### v1

# Yoga for Fourier–Mukai partnership

Abstract.
We study the behavior of integral transforms under various forms of base change. In particular, we establish a yoga of local algebra and fibers to test for derived equivalences via integral transforms. This partially generalizes a result of Orlov to singular varieties and strengthens many results in the literature to arbitrary fields. Additionally, it provides new insight into fibrations and their singularities arising in more arithmetically flavored contexts (e.g. projective and flat schemes over a DVR).
Key words and phrases:
Fourier–Mukai partners, integral transforms, singular varieties, derived categories
2020 Mathematics Subject Classification:
14A30 (primary), 14F08, 18G80, 14B05
1. Introduction
1.1. Setting the stage
The context of our work starts with Mukai [Muk81]. Loc. cit. introduced the notion of integral transforms to study triangulated equivalences between derived categories associated with abelian varieties. Over time, integral transforms have played a central role in the interplay between derived categories and algebraic geometry (e.g. [Bri02, Kaw02]), with further applications in string theory [AHR05, ASS05].
Much of this progress has focused on smooth varieties, where geometric results rely on categorical methods. However, extending these methods to singular varieties remains challenging, partly because bounded coherent complexes are typically not perfect. Thus, an important task is to develop new results or techniques that improve our understanding of integral transforms in such singular settings.
Recent work has begun to address these difficulties. Criteria determining when an integral transform is fully faithful or an equivalence have been established for varieties over algebraically closed fields, generalizing results from the smooth case [HLS07, HLS09] and building upon foundational work of Bondal–Orlov [AB95]. Likewise, Bondal–Orlov’s reconstruction theorem for smooth projective varieties has been extended to projective curves [Spe23] and (anti-)Fano Gorenstein varieties [Bal11].
There is also growing interest in studying integral transforms in the relative setting. Specifically, let $Y_{1}$ and $Y_{2}$ be proper $S$-schemes where $S$ is Noetherian. An integral transform is an exact functor $\Phi_{K}:=\mathbf{R}(p_{2})_{\ast}(\mathbf{L}p_{1}^{\ast}(-)\otimes^{\mathbf{L}}K)\colon D_{\operatorname{qc}}(Y_{1})\to D_{\operatorname{qc}}(Y_{2})$ where $K\in D_{\operatorname{qc}}(Y_{1}\times_{S}Y_{2})$ and $p_{i}\colon Y_{1}\times_{S}Y_{2}\to Y_{i}$ are the projection morphisms. If $\Phi_{K}$ is an equivalence, then we say $Y_{1}$ and $Y_{2}$ are Fourier–Mukai partners (over $S$).
Such transforms naturally arise in the study of fibrations; that is, in situations involving proper and flat morphisms. See e.g. [HLS07, HLS09], which dealt with families whose special fibers have mild singularities. In particular, loc. cit. investigated when an integral transform between locally projective fibrations $Y_{i}\to S$ is an equivalence for $S$ of finite type over an algebraically closed field. It was shown that the induced integral transform on special fibers (i.e. closed points) over $S$ controls the behavior of equivalences and full faithfulness.
In our work, we reveal additional depth to this perspective of exploiting special fibers. Specifically, we study the behavior of integral transforms under changes of the base scheme in more general settings that includes local algebra. Moreover, our approach naturally extends several results in the literature to arbitrary fields, and at the same time yields new insight for singularities of fibrations.
1.2. Preservation of adjointness
Neeman has developed a new framework called approximable triangulated categories (see e.g. [CNS25, Nee23]), which has resolved open conjectures in algebraic geometry [Nee21a, Nee24]. As a consequence of this machinery, any integral transform $\Phi_{K}$ inducing an equivalence on $D_{\operatorname{qc}}$ restricts to equivalences on the subcategories of bounded complexes with coherent cohomology $D^{b}_{\operatorname{coh}}$ and perfect complexes $\operatorname{Perf}$. Moreover, an equivalence on these ‘smaller’ categories implies an equivalence on the ‘larger’ category. See Remark˜2.7 for details.
For those concerned with equivalences, this allows the study of integral transforms to be reduced to their behavior on these subcategories. Whether an integral transform restricts to these smaller categories is now well-understood [Riz17, Bal09, DLM25]. These results shed light to the existence of left or right adjoints to integral transforms and the conditions under which such adjoints also restrict to these subcategories.
However, the behavior of adjointness under base change is not yet fully understood. Our focus is on base changes induced by affine morphisms of finite tor-dimension. In particular, this leads us to the following, which summarizes Section˜3.
Proposition 1.1.
Suppose $t\colon T\to S$ is a morphism of Noetherian schemes. Let $Y_{1}$ and $Y_{2}$ be proper and flat $S$-schemes. Denote by $\pi\colon Y_{1}\times_{S}Y_{2}\times_{S}T\to Y_{1}\times_{S}Y_{2}$ the natural morphism. Consider $K\in D^{-}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$.
  1. (1)
    If $t$ is affine and $\Phi_{K}$ restricts to $D^{b}_{\operatorname{coh}}$ (resp. $\operatorname{Perf}$), then so does $\Phi_{\mathbf{L}\pi^{\ast}K}$ (see Lemma˜3.3). Additionally, the converse holds if $t$ is faithfully flat (see Corollary˜3.6).
  2. (2)
    Assume $t$ is affine and locally of finite tor-dimension.
    1. (a)
      If $\Phi_{K}$ restricts to $\operatorname{Perf}$, then $\mathbf{L}\pi^{\ast}$ preserves the adjunction of $\Phi_{K}$ with its right adjoint (see Proposition˜3.8).
    2. (b)
      If $\Phi_{K}$ restricts to $D^{b}_{\operatorname{coh}}$, then $\mathbf{L}\pi^{\ast}$ preserves the adjunction of $\Phi_{K}$ with its left adjoint (see Proposition˜3.12).
In short, Proposition˜1.1 describes the behavior of integral transforms under base change when restricted to the small categories, specifically addressing whether the preservation of adjoints occurs. Key instances where it applies are morphisms $t$ arising from field extensions or finite surjective morphisms to a Dedekind scheme (i.e. smooth projective curves).
1.3. Fully faithfulness & equivalences
Now, we have a clearer understanding of how integral transforms behave under base change. Particularly, regarding the preservation of adjointness and the restriction to smaller categories. Next, we shift gears to studying fully faithfulness and equivalences under base change; specifically, the ascent and descent of these properties. Before doing so, we review the relevant literature a bit more.
As mentioned earlier, the work of [HLS07, HLS09] studied these questions for fibrations whose base is of finite type over an algebraically closed field. A key ingredient in loc. cit. was to look at the behavior on special fibers. More recently, such questions have also been studied for smooth proper morphisms over a Noetherian base scheme (see [Kur24, Proposition/Definition 2.3]). However, smoothness constraints can restrict the examples which one might be interested in, e.g. fibrations admitting singular fibers.
A result of Orlov addresses similar questions in the case of smooth varieties over a field. Specifically, for smooth projective varieties over a field, an integral transform is fully faithful or an equivalence if, and only if, the induced integral transform satisfies the same property after any base field extension [Orl02, Lemma 2.12]. The strategy of loc. cit. leverages ‘canonical morphisms’ between the kernels of integral transforms that describe the (co)units of adjunction. However, it remains unclear whether such canonical morphisms exist in the singular setting. This uncertainty is another source of motivation for our work, as we aimed to bypass the lack of these tools in singular cases using different techniques.
The brings attention to the following, which summarizes Section˜4.
Theorem 1.2 (see Corollary˜4.7).
Consider affine morphism $t\colon T\to S$ locally of finite tor-dimension of Noetherian schemes whose target is semiseparated. Suppose $Y_{1}$ and $Y_{2}$ are proper and flat $S$-schemes. Denote by $\pi\colon Y_{1}\times_{S}Y_{2}\times_{S}T\to Y_{1}\times_{S}Y_{2}$ the natural morphism. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over each $Y_{i}$. If $\Phi_{K}$ restricts to a fully faithful functor (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$, then so does $\Phi_{\mathbf{L}\pi^{\ast}K}$.
In practice, many schemes we may consider are separated, and hence, semiseparated. As a pleasant consequence, this allows us to partially generalize Orlov’s result to the singular setting.
Corollary 1.3 (cf. [Orl02, Lemma 2.12]).
Consider proper schemes $Y_{1}$ and $Y_{2}$ over a field $k$. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{k}Y_{2})$ be relatively perfect over each $Y_{i}$. If $\Phi_{K}$ restricts to a fully faithful functor (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$, then so does $\Phi_{K_{L}}$ for any field extension $L/k$.
Our proofs require careful bookkeeping of how the (co)units of adjunctions for integral transforms behave under base change. However, additional subtleties also arise in this context. In particular, we need to work in the larger category $D_{\operatorname{qc}}$ in order to have a firm grasp on ascending fully faithfulness; see e.g. Proposition˜4.4 and Theorem˜4.6. Notably, the techniques we use are independent of those appearing in [Orl02, Lemma 2.12].
1.4. Affine variations
Naturally, we sought analogs of Theorem˜1.2 in which $t$ need not be locally of finite tor-dimension. Ideally, such a framework would include morphisms from special fibers as in [HLS07, HLS09]. Accordingly, we work in the setting of affine morphisms to the base, while imposing a mild restriction on one of the $Y_{i}\to S$.
The following result is a slight generalization of [HLS09, Proposition 2.15]. In particular, we perform base change along affine morphisms rather than closed immersions of closed points. Note that a key difference between what follows and Sections˜3 and 4 is that we impose one of the $Y_{i}\to S$ be locally projective (as opposed to proper like our earlier results).
Theorem 1.4 (see Corollary˜5.5).
Consider affine morphism $t\colon T\to S$ to a semiseparated Noetherian scheme. Suppose $Y_{1}\to S$ is flat, locally projective and $Y_{2}\to S$ is flat, proper. Denote by $t^{\prime}\colon Y_{1}\times_{S}Y_{2}\times_{S}T\to Y_{1}\times_{S}Y_{2}$ the natural morphism. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over each $Y_{i}$. If $\Phi_{K}$ restricts to a fully faithful functor (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$, then so does $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$.
It is worthwhile to note that Theorem˜1.4 yields an independent proof of the ascending full faithfulness or equivalences in [HLS09, Proposition 2.15]. To do so, we show in suitable cases that the unit of an adjoint pair of integral transforms is preserved under base change when the left adjoint is fully faithful (see Proposition˜5.3 and Proposition˜4.2). However, it is not clear to us whether this remains true when the left adjoint is not fully faithful a priori. This condition seems necessary (at least to us) to establish the ‘descent’ part of [HLS09, Proposition 2.15] in the case of special fibers.
1.5. Applications
To conclude our work, we discuss the ramifications of Theorems˜1.2 and 1.4. Our results extend the literature on schemes of finite type over algebraically closed fields to the more general setting of arbitrary fields. They reveal an interplay between the intuition of [HLS07, HLS09] concerning special fibers and our work on faithfully flat descent along affine morphisms. In particular, certain singularities are shown to be derived invariants for fibrations over Noetherian schemes, broadening the scope to include arithmetically flavored situations of potential interest for future research.
For a pair of geometrically integral projective curves over a field $k$ of characteristic zero, Fourier–Mukai partnership implies they are $k$-forms (see Proposition˜6.1). Here, a $k$-form means there exists a finite extension $L/k$ such that their base changes along $L/k$ become isomorphic. The result holds for singular curves and enriches algebraically closed case of [Spe23]. In Example˜6.2, we exhibit a pair of non-isomorphic elliptic curves over $\mathbb{Q}$ that are nevertheless $\mathbb{Q}$-forms. For fibrations of relative dimension one over a Noetherian $\mathbb{Q}$-scheme, Fourier–Mukai partnership forces the genera of special fibers to coincide (see Corollary˜6.3).
Smoothness for proper schemes over an arbitrary field is likewise invariant under Fourier–Mukai partnership (see Proposition˜6.4). This subtlety arises especially over imperfect fields. Since regularity is characterized by $D^{b}_{\operatorname{coh}}=\operatorname{Perf}$, one might expect this condition to persist after a change of field extensions, but there exist regular projective curves that become singular after such a base change (see e.g. [Kol11, Remark 16]).
In the setting of projective fibrations, smoothness of the total space is also a derived invariant (see Corollary˜6.5). Finally, generalizing [HLS09], we show that the Cohen–Macaulay and Gorenstein properties of a morphism remain invariant under Fourier–Mukai partnership for Noetherian base schemes (see Proposition˜6.6).
Notation
Let $X$ be a Noetherian scheme. We work with the following triangulated categories:
  1. (1)
    $D(X):=D(\operatorname{Mod}(X))$ is the derived category of $\mathcal{O}_{X}$-modules.
  2. (2)
    $D_{\operatorname{qc}}(X)$ is the (strictly full) subcategory of $D(X)$ consisting of complexes with quasi-coherent cohomology.
  3. (3)
    $D_{\operatorname{coh}}^{b}(X)$ (resp. $D_{\operatorname{qc}}^{b}(X)$) is the (strictly full) subcategory of $D(X)$ consisting of complexes having bounded and coherent (resp. quasi-coherent) cohomology.
  4. (4)
    $\operatorname{Perf}(X)$ is the (strictly full) subcategory of $D_{\operatorname{qc}}(X)$ consisting of the perfect complexes on $X$.
At times if $X$ is affine, we abuse notation and write $D(R):=D_{\operatorname{qc}}(X)$ where $R:=H^{0}(X,\mathcal{O}_{X})$ are the global sections; similar conventions will occur for the other categories.
Acknowledgments.
Lank was supported under the ERC Advanced Grant 101095900-TriCatApp. Part of this work was completed while Lank visited the Basque Center for Applied Mathematics (BCAM) and thanks them for their hospitality. The authors thank Dylan Spence, Ana Cristina López-Martín, and Elías Guisado Villalgordo for discussions in an earlier version of the work.
2. Preliminaries
Let $X$ be a Noetherian scheme.
2.1. Perfect & pseudocoherent complexes
We say $P\in D_{\operatorname{qc}}(X)$ is perfect if it is affine locally isomorphic to a bounded complex of locally free sheaves of finite rank. Denote by $\operatorname{Perf}(X)$ for the full subcategory of perfect complexes. Also, given an integer $N$, we say $E\in D_{\operatorname{qc}}(X)$ is $N$-pseudocoherent if locally there is a morphism $P\to E$ from some $P\in\operatorname{Perf}(X)$ such that the induced morphism on cohomology sheaves $\mathcal{H}^{n}(P)\to\mathcal{H}^{n}(E)$ is an isomorphism for $n>N$ and is surjective for $n=N$. We say $E\in D_{\operatorname{qc}}(X)$ is pseudocoherent if it is $N$-pseudocoherent for every $N\in\mathbb{Z}$. In general, since $X$ is a scheme, $E\in D_{\operatorname{qc}}(X)$ is pseudocoherent if and only if locally on $X$ there is a quasi-isomorphism $P\to E$, with $P$ a bounded above complex of finite-rank free modules [GW23, Corollary 22.47]. As $X$ is Noetherian, $E\in D_{\operatorname{qc}}(X)$ is pseudocoherent if, and only if, $\mathcal{H}^{n}(E)$ is coherent for all $n$ and vanishes for $n\gg 0$ (see e.g. [Sta25, Tag 08E8]). The subcategory of pseudocoherent complexes of $D_{\operatorname{qc}}(X)$ coincides with $D^{-}_{\operatorname{coh}}(X)$ (i.e. the full subcategory of complexes with bounded above and coherent cohomology). Consider a morphism $f\colon Y\to X$ of Noetherian schemes. Then one has $\mathbf{L}f^{\ast}D^{-}_{\operatorname{coh}}(X)\subseteq D^{-}_{\operatorname{coh}}(Y)$ (see e.g. [Sta25, Tag 08H4]). We will use this fact liberally.
While straightforward, the following is useful.
Lemma 2.1.
Let $f\colon Y\to X$ be an affine faithfully flat morphism of Noetherian schemes. An object $E\in D_{\operatorname{qc}}(X)$ belongs to $D^{-}_{\operatorname{coh}}(X)$ (resp. $\operatorname{Perf}(X)$) if, and only if, $\mathbf{L}f^{\ast}E\in D^{-}_{\operatorname{coh}}(Y)$ (resp. $\mathbf{L}f^{\ast}E\in\operatorname{Perf}(Y)$).
Proof.
This is a special case of [GW23, Proposition 22.52], but we spell it out for convenience. The problem is affine local. So, we can reduce to the affine case. Then the claim follows from [Sta25, Tag 068R] (resp. [GW23, Lemma 21.172]). ∎
2.2. Integral transforms
Suppose $f_{1}\colon Y_{1}\to S$ and $f_{2}\colon Y_{2}\to S$ are morphisms of finite type to a Noetherian scheme. Consider the fibered square
[…]
For any $K\in D_{\operatorname{qc}}(Y_{1}\times_{S}Y_{2})$, its associated integral $S$-transform is the exact functor $\Phi_{K}\colon D_{\operatorname{qc}}(Y_{1})\to D_{\operatorname{qc}}(Y_{2})$ given by $E\mapsto\mathbf{R}p_{2,\ast}(\mathbf{L}p_{1}^{\ast}E\otimes^{\mathbf{L}}K)$. We say $Y_{1}$ and $Y_{2}$ are Fourier–Mukai $S$-partners if there is such a $K$ for which $\Phi_{K}$ yields an equivalence $D^{b}_{\operatorname{coh}}(Y_{1})\to D^{b}_{\operatorname{coh}}(Y_{2})$. When clear from context, the hyphen ‘$S$-’ is dropped.
Now, we impose each $f_{i}$ be proper and flat. An object $K\in D^{-}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ is called relatively perfect over $Y_{1}$ if
$\mathbf{R}(p_{1})_{\ast}(K\otimes^{\mathbf{L}}\operatorname{Perf}(Y_{1}\times_{S}Y_{2}))\subseteq\operatorname{Perf}(Y_{1}).$
By symmetry, the notion makes sense over $Y_{2}$. The following two facts are [DLM25, Proposition’s 3.6 & 3.2]: $K$ being relatively perfect over $Y_{1}$ is equivalent to $\Phi_{K}(D^{b}_{\operatorname{coh}}(Y_{1}))\subseteq D^{b}_{\operatorname{coh}}(Y_{2})$; $K$ being relatively perfect over $Y_{2}$ is the same as $\Phi_{K}(\operatorname{Perf}(Y_{1}))\subseteq\operatorname{Perf}(Y_{2})$.
More generally, the definition of relatively perfect objects make sense for any proper morphism of Noetherian schemes $f\colon Y\to X$. Indeed, an object $K\in D^{-}_{\operatorname{coh}}(Y)$ is called $f$-perfect if $\mathbf{R}f_{\ast}(K\otimes^{\mathbf{L}}\operatorname{Perf}(Y))\subseteq\operatorname{Perf}(X)$. So, in our setting, $K\in D^{-}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ is relatively perfect over $Y_{1}$ precisely when it is $p_{1}$-perfect. This notion has appeared in the literature, but we record a lemma for convenience to avoid any potential confusion.
Lemma 2.2.
Let $f\colon Y\to X$ be a proper morphism of Noetherian schemes. Then the following are equivalent for any $E\in D^{-}_{\operatorname{coh}}(Y)$:
  1. (1)
    $E$ is $f$-perfect.
  2. (2)
    $E\otimes^{\mathbf{L}}\mathbf{L}f^{\ast}D^{b}_{\operatorname{coh}}(X)\subseteq D^{b}_{\operatorname{coh}}(Y)$.
  3. (3)
    $E\otimes^{\mathbf{L}}\mathbf{L}f^{\ast}D^{b}_{\operatorname{qc}}(X)\subseteq D^{b}_{\operatorname{qc}}(Y)$.
Proof.
[…]
Remark 2.3.
A useful observation from Lemma˜2.2 is that any object $E\in D^{-}_{\operatorname{coh}}(Y)$ being $f$-perfect implies it belongs to $D^{b}_{\operatorname{coh}}(Y)$.
2.3. Generation for triangulated categories
Let $\mathcal{T}$ be a triangulated category with shift functor $1]\colon\mathcal{T}\to\mathcal{T}$. Fix $\mathcal{S}\subseteq\mathcal{T}$. Denote by $\operatorname{add}(\mathcal{S})$ the strictly full (that is, closed under isomorphisms) subcategory of $\mathcal{T}$ consisting of all direct summands of objects of the form $\oplus_{n\in\mathbb{Z}}S_{n}^{\oplus r_{n}}[n]$ where $S_{n}\in\mathcal{S}$ for all $n$ and finitely many $r_{n}$ are nonzero. Inductively, let
$\langle\mathcal{S}\rangle_{n}:=\begin{cases}\operatorname{add}(\varnothing)&\text{if }n=0,\\ \operatorname{add}(\mathcal{S})&\text{if }n=1,\\ \operatorname{add}(\{\operatorname{cone}\phi\mid\phi\in\operatorname{Hom}(\langle\mathcal{S}\rangle_{n-1},\langle\mathcal{S}\rangle_{1})\})&\text{if }n>1.\end{cases}$
Additionally, let $\langle\mathcal{S}\rangle:=\langle\mathcal{S}\rangle_{\infty}:=\cup_{n\geq 0}\langle\mathcal{S}\rangle_{n}$. It may be checked that $\langle\mathcal{S}\rangle$ is the smallest thick (that is, triangulated subcategory closed under direct summands) subcategory containing $\mathcal{S}$. In the case $\mathcal{S}$ consists of a single object $G$, we write the above as $\langle G\rangle_{n}$ and $\langle G\rangle$. An object $G\in\mathcal{T}$ is called a classical generator for $\mathcal{T}$ if $\langle G\rangle=\mathcal{T}$.
Now, we impose $\mathcal{T}$ be a compactly generated triangulated category. To clarify, we require that $\mathcal{T}$ admitting all small coproducts. Denote the collection of compact objects of $\mathcal{T}$ by $\mathcal{T}^{c}$. We say $G\in\mathcal{T}^{c}$ is a compact generator for $\mathcal{T}$ if for any $E\in\mathcal{T}$, one has $\operatorname{Hom}(G,E[n])=0$ for all $n\in\mathbb{Z}$ if, and only if, $E=0$. A well-known fact is that classical generators for $\mathcal{T}^{c}$ coincide with being compact generators for $\mathcal{T}$ (see e.g. [Sta25, Tag 09SR]).
The following is an important case and is freely used without mention.
Example 2.4 ([BvdB03, Theorem 3.1.1]).
Let $X$ be a quasi-compact quasi-separated scheme. The collection of compact objects in $D_{\operatorname{qc}}(X)$ coincides with $\operatorname{Perf}(X)$. Moreover, $D_{\operatorname{qc}}(X)$ is compactly generated by a single object from $\operatorname{Perf}(X)$.
The following is known for some circles, but we add it for self-containedness (see e.g. [KL15, Lemma’s 2.10 & 2.12]).
Lemma 2.5.
Let $F\colon\mathcal{T}\leftrightarrows\mathcal{S}\colon G$ be a pair of exact adjoint functors between compactly generated triangulated categories. Assume $G$ preserves small coproducts. Then $F$ is an equivalence if, and only if, $F$ restricts to an equivalence $\mathcal{T}^{c}\to\mathcal{S}^{c}$. A similar statement holds for $F$ being fully faithful. Also, if $F$ restricts to a fully faithful functor $\mathcal{T}^{c}\to\mathcal{S}^{c}$ and $F(\mathcal{T}^{c})$ compactly generates $\mathcal{S}$, then $F\colon\mathcal{S}\to\mathcal{T}$ is an equivalence.
Proof.
It is straightforward to check that $F$ being an equivalence induces an equivalence $\mathcal{T}^{c}\to\mathcal{S}^{c}$. So, we check the converse. To do so, we show that $F$ and $G$ are both fully faithful. Denote the counit and unit of the adjunction respectively by $\epsilon$ and $\eta$. Set $\mathcal{T}^{\prime}$ and $\mathcal{S}^{\prime}$ the strictly full subcategories respectively of $\mathcal{T}$ and $\mathcal{S}$ consisting of objects where $\epsilon$ and $\eta$ are isomorphisms. One may verify that $\mathcal{T}^{\prime}$ and $\mathcal{S}^{\prime}$ are closed under shifts, iterated cones, and small coproducts (hence, homotopy colimits). However, [Sta25, Tag 09SN] tells us each object in $\mathcal{T}$ and $\mathcal{S}$ can be obtained respectively by $\mathcal{T}^{c}$ and $\mathcal{S}^{c}$ using these operations. Hence, $\mathcal{T}\subseteq\mathcal{T}^{\prime}$ and $\mathcal{S}\subseteq\mathcal{S}^{\prime}$, which completes the proof of the first claim. The remaining claims follow similarly. ∎
Terminology 2.6.
Let $\Phi\colon D_{\operatorname{qc}}(Y_{1})\to D_{\operatorname{qc}}(Y_{2})$ be an exact functor where each $Y_{i}$ is Noetherian. Suppose $\mathcal{S}\in\{\operatorname{Perf},D^{b}_{\operatorname{coh}},D^{-}_{\operatorname{coh}}\}$. If $\Phi_{K}(\mathcal{S}(Y_{1}))\subseteq\mathcal{S}(Y_{2})$, then we say $\Phi_{K}$ restricts to an exact functor on $\mathcal{S}$. The same goes for properties such as being fully faithful or an equivalence.
Remark 2.7.
Let $\Phi\colon D_{\operatorname{qc}}(Y_{1})\to D_{\operatorname{qc}}(Y_{2})$ be an exact functor where each $Y_{i}$ is Noetherian. Then the following are equivalent:
  1. (1)
    $\Phi$ is an equivalence
  2. (2)
    $\Phi$ restricts to an equivalence on $D^{-}_{\operatorname{coh}}$
  3. (3)
    $\Phi$ restricts to an equivalence on $D^{b}_{\operatorname{coh}}$
  4. (4)
    $\Phi$ restricts to an equivalence on $\operatorname{Perf}$.
[…]
3. Preservation of adjointness
This section studies the behavior of adjoints for integral transforms under a change of base schemes. To avoid repeatedly stating similar constraints, the following is used as a placeholder throughout our work.
Setup 3.1.
Let $t\colon T\to S$ be a morphism of Noetherian schemes. Suppose $f_{i}\colon Y_{i}\to S$ are proper flat morphisms with $i\in\{1,2\}$. Consider the commutative diagram
[…]
obtained by base change along $t$.
Lemma 3.2.
Consider ˜3.1 with $t$ a morphism locally of finite tor-dimension. Then $t_{1},t_{2},t^{\prime}$ are locally of finite tor-dimension. In particular, each face of Equation˜3.1 is a tor-independent square (in the sense of [Lip09, Definition 3.10.2]).
Proof.
This first claim is affine local on $S$. Indeed, if needed, we can find an affine open cover of $S$ such that the base changes of $t$ along the associated open immersion are of finite tor-dimension. Then, for each such open immersion of a component in the affine cover, we have a similar diagram to Equation˜3.1 (i.e. a ‘base change cube’). If we can show the desired properties on the morphisms arising in each diagram associated to the components of the affine open cover of $S$, then the desired claim follows. So, we may impose that $t$ be of finite tor-dimension.
By base change, we know that $g_{1},g_{2},f^{\prime}_{2},f^{\prime}_{1},g_{2}^{\prime},g_{1}^{\prime}$ are proper flat morphisms. Consider the fibered square in Equation˜3.1 induced from the base change of $f_{1}$ along $t$. It follows by [GW23, Lemma 23.27] that $t_{1}$ is of finite tor-dimension. A similar argument shows $t_{2}$ is of finite tor-dimension as well. In fact, if we work with the fibered square in Equation˜3.1 obtained by base change of $f_{2}^{\prime}$ along $t_{1}$, we see from similar reasoning that $t^{\prime}$ is locally of finite tor-dimension. This completes the proof for our first claim.
To check the second claim, it can be checked by one face at a time. Note that the same argument in what follows works whether $t$ has finite tor-dimension locally or absolutely because the conditions checked are at the level of stalks. The fibered square obtained by base change of $f_{1}$ along $t$ is tor-independent. Indeed, use [GW23, Exercise 21.32(iii)] with the constraint that $f_{1}$ is flat. Then [GW23, Exercise 21.33(v)], coupled with [Lip09, Definition 3.10.2(i) & Theorem 3.10.3], tells us the particular face is tor-independent. Now, we can proceed in a similar fashion with the remaining faces to finish the proof. ∎
Lemma 3.3.
Consider ˜3.1 with $t$ an affine morphism. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{1}$. Then $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ restricts to an exact functor on $D^{b}_{\operatorname{coh}}$. Additionally, a similar statement holds for $\operatorname{Perf}$ if $K$ is relatively perfect over $Y_{2}$.
Proof.
We prove the case of $\Phi_{\mathbf{L}(t^{\prime})K}$ restricting to an exact functor on $\operatorname{Perf}$ because a similar argument applies for $D^{b}_{\operatorname{coh}}$. As $\Phi_{K}(\operatorname{Perf}(Y_{1}))\subseteq\operatorname{Perf}(Y_{2})$, [DLM25, Proposition 3.2] implies
$\mathbf{R}(f_{1}^{\prime})_{\ast}(K\otimes^{\mathbf{L}}\operatorname{Perf}(Y_{1}\times_{S}Y_{2}))\subseteq\operatorname{Perf}(Y_{2}).$
By base change, we know that $t_{2},t_{2},t^{\prime}$ are affine morphisms. Choose $G\in\operatorname{Perf}(Y_{1}\times_{S}Y_{2})$ such that $\operatorname{Perf}(Y_{1}\times_{S}Y_{2})=\langle G\rangle$. Now, $\mathbf{L}(t^{\prime})^{\ast}G$ satisfies
$\operatorname{Perf}(Y_{1}\times_{S}Y_{2}\times_{S}T)=\langle\mathbf{L}(t^{\prime})^{\ast}G\rangle,$
see e.g. [Sta25, Tag 0BQT]. Also, via base change, we know that $g_{1},g_{2},f^{\prime}_{2},f^{\prime}_{1},g_{2}^{\prime},g_{1}^{\prime}$ are proper flat morphisms. Using flat base change (see e.g. [GW23, Remark 22.94 & Theorem 22.99]), we can tie things together and see that
$\mathbf{R}(g^{\prime}_{1})_{\ast}$ $(\mathbf{L}(t^{\prime})^{\ast}K\otimes^{\mathbf{L}}\operatorname{Perf}(Y_{1}\times_{S}Y_{2}\times_{S}T))$
$\subseteq\mathbf{R}(g^{\prime}_{1})_{\ast}\langle\mathbf{L}(t^{\prime})^{\ast}(K\otimes^{\mathbf{L}}G)\rangle$
$\subseteq\langle\mathbf{R}(g^{\prime}_{1})_{\ast}\mathbf{L}(t^{\prime})^{\ast}(K\otimes^{\mathbf{L}}G)\rangle$
$\subseteq\langle\mathbf{L}t_{2}^{\ast}\mathbf{R}(f_{1}^{\prime})_{\ast}(K\otimes^{\mathbf{L}}G)\rangle$
$\subseteq\operatorname{Perf}(Y_{2}\times_{S}T).$
Therefore, by [DLM25, Proposition 3.2], $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ restricts to an exact functor on $\operatorname{Perf}$. ∎
Lemma 3.4.
Consider ˜3.1. Let $K\in D_{\operatorname{qc}}(Y_{1}\times_{S}Y_{2})$. Then there is a natural isomorphism $\beta^{K}\colon\mathbf{L}t_{2}^{\ast}\circ\Phi_{K}\to\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\circ\mathbf{L}t_{1}^{\ast}$.
Proof.
This follows from the string of isomorphisms for each $E\in D_{\operatorname{qc}}(Y_{1}\times_{S}T)$:
$\mathbf{L}t_{2}^{\ast}\Phi_{K}(E)$ $\cong\mathbf{L}t_{2}^{\ast}\mathbf{R}(f^{\prime}_{1})_{\ast}(\mathbf{L}(f^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}K)$
$\cong\mathbf{R}(g^{\prime}_{1})_{\ast}\mathbf{L}(t^{\prime})^{\ast}(\mathbf{L}(f^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}K)$
$\cong\mathbf{R}(g^{\prime}_{1})_{\ast}(\mathbf{L}(t^{\prime})^{\ast}\mathbf{L}(f^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K)$
$\cong\mathbf{R}(g^{\prime}_{1})_{\ast}(\mathbf{L}(g^{\prime}_{2})^{\ast}\mathbf{L}(t_{1}^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K))$
$=\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}(\mathbf{L}t_{1}^{\ast}E).$
Proposition 3.5.
Consider ˜3.1 where $t$ is an affine faithfully flat morphism. Let $K\in D^{-}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$. If $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}T$ (resp. $Y_{2}\times_{S}T$), then $K$ is relatively perfect over $Y_{1}$ (resp. $Y_{2}$).
Proof.
First, we show the case for $\mathbf{L}(t^{\prime})^{\ast}K$ being relatively perfect over $Y_{1}\times_{S}T$ implying $K$ is relatively perfect over $Y_{1}$. Choose $P_{i}\in\operatorname{Perf}(Y_{i})$ such that $\langle P_{i}\rangle=\operatorname{Perf}(Y_{i})$ for each $i$. From [BvdB03, Lemma 3.4.1], we know that $\langle\mathbf{L}f_{2}^{\ast}P_{1}\otimes^{\mathbf{L}}\mathbf{L}f_{1}^{\ast}P_{2}\rangle=\operatorname{Perf}(Y_{1}\times_{S}Y_{2})$. Then, by [Sta25, Tag 0BQT], it follows that
$\langle\mathbf{L}(t^{\prime})^{\ast}(\mathbf{L}f_{2}^{\ast}P_{1}\otimes^{\mathbf{L}}\mathbf{L}f_{1}^{\ast}P_{2})\rangle=\operatorname{Perf}(Y_{1}\times_{S}Y_{2}\times_{S}T).$
However, we have an isomorphism in Equation˜3.1,
$\mathbf{L}(t^{\prime})^{\ast}(\mathbf{L}f_{2}^{\ast}P_{1}\otimes^{\mathbf{L}}\mathbf{L}f_{1}^{\ast}P_{2})$ $\cong\mathbf{L}(t^{\prime})^{\ast}\mathbf{L}f_{2}^{\ast}P_{1}\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}\mathbf{L}f_{1}^{\ast}P_{2}$
$\cong\mathbf{L}(g_{2}^{\prime})^{\ast}\mathbf{L}t_{1}^{\ast}P_{1}\otimes^{\mathbf{L}}\mathbf{L}(g_{1}^{\prime})^{\ast}\mathbf{L}t_{2}^{\ast}P_{2}.$
Now, we finish the proof for the claim. Let $E\in D^{b}_{\operatorname{coh}}(Y_{1})$. As $t^{\prime}$ is faithfully flat, it follows that $\mathbf{L}t_{1}^{\ast}E\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}T)$. By Lemma˜3.4, we know that $\mathbf{L}t_{2}^{\ast}\Phi_{K}(E)\cong\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}(\mathbf{L}t^{\ast}_{1}E)$. From Lemma˜2.1, the hypothesis implies $\Phi_{K}(E)\in D_{\operatorname{coh}}^{-}(Y_{2})$. So, it suffices to show that $\mathcal{H}^{j}(\Phi_{K}(E))=0$ for all but finitely many $j\in\mathbb{Z}$.
Our hypothesis ensures that $\Phi_{\mathbf{L}(t^{\prime})K}(\mathbf{L}t_{1}^{\ast}E)\in D^{b}_{\operatorname{coh}}(Y_{2}\times_{S}T)$. Then, as $t_{2}$ is flat, for each $j\in\mathbb{Z}$,
$\operatorname{supp}(\mathcal{H}^{j}(\Phi_{K}(E)))$ $=t_{2}(\operatorname{supp}(t_{2}^{\ast}\mathcal{H}^{j}(\Phi_{K}(E))))$ ([Sta25, Tag 056J])
$=t_{2}(\operatorname{supp}(\mathbf{L}t_{2}^{\ast}\mathcal{H}^{j}(\Phi_{K}(E))))$ (Flatness of $t_{2}$)
$=t_{2}(\operatorname{supp}(\mathcal{H}^{j}(\mathbf{L}t_{2}^{\ast}\Phi_{K}(E)))).$
Hence, the desired claim regarding bounded cohomology follows. Consequently, we have $\Phi_{K}(D^{b}_{\operatorname{coh}}(Y_{1}))\subseteq D^{b}_{\operatorname{coh}}(Y_{2})$, which implies $K$ is relatively perfect over $Y_{1}$. To show the case of perfect complexes being preserved by integral transforms, we can argue as above and use Lemma˜2.1. ∎
Corollary 3.6.
Consider ˜3.1 where $t$ is an affine faithfully flat morphism. Let $K\in D^{-}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$. Then $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}T$ (resp. $Y_{2}\times_{S}T$) if, and only if, $K$ is relatively perfect over $Y_{1}$ (resp. $Y_{2}$).
Proof.
This follows from Lemma˜3.3 and Proposition˜3.5
Lemma 3.7 (Base change for relative dualizing complex).
Let $S$ be a quasi-compact quasi-separated scheme. Consider a fibered square
[…]
where $f$ is proper, flat, and of finite presentation. Then $\mathbf{L}(g^{\prime})^{\ast}f^{!}\mathcal{O}_{S}$ is a relative dualizing complex for $f^{\prime}$. In fact, there is an isomorphism $\mathbf{L}(g^{\prime})^{\ast}f^{!}\mathcal{O}_{S}\to(f^{\prime})^{!}\mathcal{O}_{S^{\prime}}$.
Proof.
This is well-known to some, but we include it for convenience. By [Sta25, Tag 0B6S & Tag 0E2Z], we know that $(f^{\prime})^{!}\mathcal{O}_{S^{\prime}}$ and $f^{!}\mathcal{O}_{S}$ are respectively relative dualizing complexes for $f^{\prime}$ and $f$. Moreover, [Sta25, Tag 0E2Y] tells us that $\mathbf{L}(g^{\prime})^{\prime}f^{!}\mathcal{O}_{S}$ is a relative dualizing complex for $f^{\prime}$. However, [Sta25, Tag 0E2W] implies relative dualizing complexes are unique up to isomorphism, so $\mathbf{L}(g^{\prime})^{\prime}f^{!}\mathcal{O}_{S}\cong(f^{\prime})^{!}\mathcal{O}_{S^{\prime}}$. ∎
Proposition 3.8.
Consider ˜3.1 where $t$ is an affine morphism locally of finite tor-dimension (e.g. $t=1_{S}$). Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{2}$. Then $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ is right adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ on $D_{\operatorname{qc}}$ where $K^{\prime}:=\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$. In particular, $K^{\prime}\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$.
Proof.
That the last claim holds follows from [Riz17, Definition 2.11 & Lemma 2.12]. So, we check the first claim. For any $E\in D_{\operatorname{qc}}(Y_{1}\times_{S}T)$ and $A\in D_{\operatorname{qc}}(Y_{2}\times_{S}T)$, there is a string of natural isomorphisms obtained by adjunctions:
$\operatorname{Hom}$ $(\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}(E),A)$
$=\operatorname{Hom}(\mathbf{R}(g^{\prime}_{1})_{\ast}(\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K),A)$ (Definition)
$\cong\operatorname{Hom}(\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K,(g^{\prime}_{1})^{!}A)$ (see e.g. [Sta25, Tag 0A9E])
$\cong\operatorname{Hom}(\mathbf{L}(g^{\prime}_{2})^{\ast}E,\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,(g^{\prime}_{1})^{!}A))$ (see e.g. [Sta25, Tag 08DH])
$\cong\operatorname{Hom}(E,\mathbf{R}(g^{\prime}_{2})_{\ast}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,(g^{\prime}_{1})^{!}A)).$
The desired claim follows if can find an isomorphism
$\mathbf{R}$ $(g^{\prime}_{2})_{\ast}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,(g^{\prime}_{1})^{!}A)$
$\cong\mathbf{R}(g^{\prime}_{2})_{\ast}\bigg(\mathbf{L}(t^{\prime})^{\ast}\big(\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})\big)\otimes^{\mathbf{L}}\mathbf{L}(g^{\prime}_{1})^{\ast}A\bigg)=:\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A).$
Note that Lemma˜3.2, as well as affineness being stable under base change, implies each vertical morphism in Equation˜3.1 are affine and locally of finite tor-dimension. Moreover, $(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}}\in D^{+}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ as upper shriek functors for proper morphisms preserve complexes with bounded below and coherent cohomology (see e.g. [Sta25, Tag 0AU1]). So, we have a string of isomorphisms
$\mathbf{L}(t^{\prime})^{\ast}$ $\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$
$\cong\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,\mathbf{L}(t^{\prime})^{\ast}(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$ ([GW23, Proposition 22.70])
[…]
By Lemma˜3.3, we see that $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ restricts to an exact functor $\operatorname{Perf}(Y_{1}\times_{S}T)\to\operatorname{Perf}(Y_{2}\times_{S}T)$. Then [DLM25, Lemma 3.2] implies
$\mathbf{R}(g_{1}^{\prime})_{\ast}(\mathbf{L}(t^{\prime})^{\ast}K\otimes^{\mathbf{L}}\operatorname{Perf}(Y_{1}\times_{S}Y_{2}\times_{S}T))\subseteq\operatorname{Perf}(Y_{2}\times_{S}T).$
There is another string of isomorphisms:
$\mathbf{R}$ $(g^{\prime}_{2})_{\ast}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,(g^{\prime}_{1})^{!}A)$
$\cong\mathbf{R}(g^{\prime}_{2})_{\ast}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}\bigg(\mathbf{L}(t^{\prime})^{\ast}K,\mathbf{L}(g^{\prime}_{1})^{\ast}A\otimes^{\mathbf{L}}(g^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}\times_{S}T}\bigg)$ ([Sta25, Tag 0B6S])
$\cong\mathbf{R}(g^{\prime}_{2})_{\ast}\bigg(\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}\big(\mathbf{L}(t^{\prime})^{\ast}K,(g^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}\times_{S}T}\big)\otimes^{\mathbf{L}}\mathbf{L}(g^{\prime}_{1})^{\ast}A\bigg)$
$\cong\mathbf{R}(g^{\prime}_{2})_{\ast}\bigg(\mathbf{L}(t^{\prime})^{\ast}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})\otimes^{\mathbf{L}}\mathbf{L}(g^{\prime}_{1})^{\ast}A\bigg).$
The last isomorphism comes from the work above, whereas the second isomorphism is [Riz17, Lemma 2.13], which completes the proof. ∎
Corollary 3.9.
Consider ˜3.1 where $t=1_{S}$ and $S$ is semiseparated (i.e. that intersection of two affine open subschemes is affine, see e.g. [AJPV08, §2]). Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{2}$. Denote by $K^{\prime}$ for the kernel of the integral transform obtained in Proposition˜3.8 which is right adjoint to $\Phi_{K}$ on $D_{\operatorname{qc}}$. Then $\Phi_{K^{\prime}}$ restricts to $D^{b}_{\operatorname{coh}}$.
Proof.
Let $U_{i}$ be an affine open cover of $S$. After base changing along each open immersion $U_{i}\to S$, we can form open covers of each $Y_{i}$. To check $\Phi_{K^{\prime}}(E)\in D^{b}_{\operatorname{coh}}(Y_{2})$, it suffices to show the restriction of $\Phi_{K^{\prime}}(E)$ to each component of the induced open cover on $Y_{2}$ has bounded and coherent cohomology. Hence, using Proposition˜3.8 and Lemma˜3.3, we can reduce to the case $S$ is affine. In this setting, the desired claim follows from [DLM25, Lemma 3.1]. Specifically, $\Phi_{K^{\prime}}\colon D^{b}_{\operatorname{coh}}(Y_{2})\to D_{\operatorname{qc}}(Y_{1})$ agrees with the unique functor $\Phi\colon D^{b}_{\operatorname{coh}}(Y_{2})\to D^{b}_{\operatorname{coh}}(Y_{1})$ of loc. cit. (indeed, [Nee21b, Example 0.7] gives a triangulated equivalence between the category of finite cohomological functors on perfect complexes and the bounded derived category of bounded pseudocoherent complexes). ∎
Remark 3.10.
In Corollary˜3.9, we require $S$ to be semiseparated solely so that the open immersion from an affine scheme to $S$ is an affine morphism. This allows us to leverage Proposition˜3.8. Note that semiseparated is a mild constraint because any separated scheme is such.
Lemma 3.11.
Let $f_{i}\colon Y_{i}\to S$ be proper flat morphisms to a Noetherian scheme where $i\in\{1,2\}$. Denote by $g_{i}\colon Y_{1}\times_{S}Y_{2}\to Y_{i}$ the natural morphisms. Suppose $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ is relatively perfect over $Y_{1}$. Then $\Phi_{K}$ admits a left adjoint on $D_{\operatorname{qc}}$. In particular, the left adjoint is of the form $\Phi_{K^{\prime}}$ where $K^{\prime}:=\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,g^{!}_{1}\mathcal{O}_{Y_{1}})$.
Proof.
This is baked into the proof of [DLM25, Proposition 4.3], but we spell it out for convenience. In short, the desired claim follows from the string of natural isomorphisms for all $E\in D_{\operatorname{qc}}(Y_{2})$ and $G\in D_{\operatorname{qc}}(Y_{1})$:
$\operatorname{Hom}$ $(E,\Phi_{K}(G))\cong\operatorname{Hom}(E,\mathbf{R}(g_{2})_{\ast}(K\otimes^{\mathbf{L}}\mathbf{L}g_{1}^{\ast}G))$ (Definition)
$\cong\operatorname{Hom}(\mathbf{L}g_{2}^{\ast}E,K\otimes^{\mathbf{L}}\mathbf{L}g_{1}^{\ast}G)$ (Adjunction)
[…]
Proposition 3.12.
Consider ˜3.1 where $t$ is an affine morphism locally of finite tor-dimension. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{1}$. Then $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ is left adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ on $D_{\operatorname{qc}}$ where $K^{\prime}:=\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}})$. In particular, $K^{\prime}\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$.
Proof.
That the last claim holds follows from [Riz17, Definition 2.11 & Lemma 2.12]. So, we check the first claim. By Lemma˜3.3, $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}T$. Moreover, from Lemma˜3.11, the kernel of the integral transform which is left adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is given $\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,(g^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}\times_{S}T})$. So, it suffices to show there is an isomorphism
$\mathbf{R}(g_{2}^{\prime})_{\ast}$ $(\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,(g^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}\times_{S}T})\otimes^{\mathbf{L}}\mathbf{L}(g_{1}^{\prime})^{\ast}E)$
$\to\mathbf{R}(g_{2}^{\prime})_{\ast}(\mathbf{L}(t^{\prime})^{\ast}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}})\otimes^{\mathbf{L}}\mathbf{L}(g_{1}^{\prime})^{\ast}E).$
This follows if we can find an isomorphism
$\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,(g^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}\times_{S}T})$
$\to\mathbf{L}(t^{\prime})^{\ast}(\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}})).$
As Lemma˜3.2 and affineness is stable under base change, it follows that each vertical morphism in Equation˜3.1 are affine and locally of finite tor-dimension. Also, $(g^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}}\in D^{+}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ as upper shriek functors for proper morphisms preserve complexes with bounded below and coherent cohomology (see e.g. [Sta25, Tag 0AU1]). Hence, we have a string of isomorphisms
$\mathbf{L}(t^{\prime})^{\ast}$ $\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}})$
$\cong\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,\mathbf{L}(t^{\prime})^{\ast}(f^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}})$ ([GW23, Proposition 22.70])
[…]
This completes the proof. ∎
4. Fully faithfulness & equivalences
This section is concerned with the behavior of fully faithfulness or an equivalence for an integral transform under a change of base scheme.
Lemma 4.1.
Consider ˜3.1. Let $K\in D_{\operatorname{qc}}(Y_{1}\times_{S}Y_{2})$. Then there is a natural isomorphism $\alpha^{K}\colon(t_{2})_{\ast}\circ\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\to\Phi_{K}\circ(t_{1})_{\ast}$.
Proof.
This follows from the string of natural isomorphisms for each $E\in D_{\operatorname{qc}}(Y_{1}\times_{S}T)$:
$\mathbf{R}(t_{2})_{\ast}\circ\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}(E)$ $=\mathbf{R}(t_{2})_{\ast}\mathbf{R}(g^{\prime}_{1})_{\ast}\big(\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K\big)$
$\cong\mathbf{R}(t_{2}\circ g^{\prime}_{1})_{\ast}\big(\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K\big)$
$\cong\mathbf{R}(f_{1}^{\prime}\circ t^{\prime})_{\ast}\big(\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K\big)$
$\cong\mathbf{R}(f_{1}^{\prime})_{\ast}\bigg(\mathbf{R}(t^{\prime})_{\ast}\big(\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K\big)\bigg)$
$\cong\mathbf{R}(f_{1}^{\prime})_{\ast}\bigg(\mathbf{R}(t^{\prime})_{\ast}\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}K\bigg)$
$\cong\mathbf{R}(f_{1}^{\prime})_{\ast}\bigg(\mathbf{L}(f^{\prime}_{2})^{\ast}\mathbf{R}(t_{1})_{\ast}E\otimes^{\mathbf{L}}K\bigg)$
$=\Phi_{K}\circ\mathbf{R}(t_{1})_{\ast}(E).$
Proposition 4.2.
Consider ˜3.1 where $t$ is an affine morphism locally of finite tor-dimension. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{2}$. Denote by $K^{\prime}$ for the kernel of the integral transform obtained in Proposition˜3.8 which is right adjoint to $\Phi_{K}$ on $D_{\operatorname{qc}}$. If $\Phi_{K}$ is fully faithful on $D_{\operatorname{qc}}$, then $\mathbf{L}t^{\ast}_{1}$ preserves the unit $\eta$ of the adjoint pair $\Phi_{K}$ and $\Phi_{K^{\prime}}$ on $D_{\operatorname{qc}}$. That is, $\mathbf{L}t_{1}^{\ast}\eta_{E}$ is the unit of $\mathbf{L}t_{1}^{\ast}E$ for the adjoint pair $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ and $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$.
Proof.
By Proposition˜3.8, $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ and $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ are adjoint functors on $D_{\operatorname{qc}}$. Choose $E\in D_{\operatorname{qc}}(Y_{1})$. We want to show that $\mathbf{L}t_{1}^{\ast}(\eta_{E})$ is the unit of $\mathbf{L}t_{1}^{\ast}E$ for this adjunction.
Let $\beta\colon\mathbf{L}t^{\ast}_{1}E\to\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$ be a morphism in $D_{\operatorname{qc}}(Y_{1}\times_{S}T)$ for any $A\in D_{\operatorname{qc}}(Y_{2}\times_{S}T)$. Using that $\mathbf{L}t_{1}^{\ast}$ is left adjoint to $\mathbf{R}(t_{1})_{\ast}$ on $D_{\operatorname{qc}}$, adjunction tells us there exists a unique morphism $\beta^{\prime}$ making the following diagram commute,
$\mathbf{L}t^{\ast}_{1}E$$\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$$\mathbf{L}t_{1}^{\ast}\mathbf{R}(t_{1})_{\ast}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$$\scriptstyle{\beta}$$\scriptstyle{\beta^{\prime}}$$\scriptstyle{\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}}$
where $\xi^{t_{1}}$ is the counit of derived pullback and pushforward along $t_{1}$. Now, $\beta^{\prime}$ is of the form $\mathbf{L}t^{\ast}_{1}(\beta^{\prime\prime})$ where $\beta^{\prime\prime}\colon E\to\mathbf{R}(t_{1})_{\ast}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$. Furthermore, by the universal property of the counit, the morphism $\beta^{\prime\prime}$ is the unique morphism such that $\beta=\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\mathbf{L}t^{\ast}_{1}(\beta^{\prime\prime})$. Similarly, using that $\Phi_{K}$ is left adjoint to $\Phi_{K^{\prime}}$ and the natural isomorphism $\alpha$ of Lemma˜4.1, there is a $\beta^{\prime\prime\prime}$ giving a diagram
[…]
which commutes. In other words,
$\beta^{\prime}=\mathbf{L}t_{1}^{\ast}((\alpha^{A})^{-1}\circ\beta^{\prime\prime\prime})\circ\mathbf{L}t_{1}^{\ast}(\eta_{E}),$
which tells us $\beta$ factors through $\mathbf{L}t_{1}^{\ast}(\eta_{E})$.
It remains to show that this factorization is unique. So, suppose there is a morphism $h\colon\mathbf{L}t_{1}^{\ast}\Phi_{K^{\prime}}\Phi_{K}(E)\to\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$ such that $\beta=h\circ\mathbf{L}t_{1}^{\ast}(\eta_{E})$. Once again, using that $\mathbf{L}t_{1}^{\ast}$ is left adjoint to $\mathbf{R}(t_{1})_{\ast}$, we have a morphism
$h^{\prime}\colon\mathbf{L}t_{1}^{\ast}\Phi_{K^{\prime}}\Phi_{K}(E)\to\mathbf{L}t_{1}^{\ast}\mathbf{R}(t_{1})_{\ast}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$
giving a commutative diagram
$\mathbf{L}t_{1}^{\ast}\Phi_{K^{\prime}}\Phi_{K}(E)$$\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$$\mathbf{L}t_{1}^{\ast}\mathbf{R}(t_{1})_{\ast}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$$\scriptstyle{h}$$\scriptstyle{h^{\prime}}$$\scriptstyle{\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}}$
with $\xi^{t_{1}}$ as above and $h^{\prime}=\mathbf{L}t^{\ast}_{1}h^{\prime\prime}$ for a morphism $h^{\prime\prime}\colon\Phi_{K^{\prime}}\Phi_{K}(E)\to\mathbf{R}(t_{1})_{\ast}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$. Here, $h^{\prime\prime}$ is the unique morphism such that $h=\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\mathbf{L}t^{\ast}_{1}h^{\prime\prime}$.
Tying things together, we see that
$\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\mathbf{L}t^{\ast}_{1}(h^{\prime\prime}\circ\eta_{E})=(\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\mathbf{L}t^{\ast}_{1}(h^{\prime\prime}))\circ\mathbf{L}t^{\ast}_{1}(\eta_{E})=h\circ\mathbf{L}t^{\ast}_{1}(\eta_{E})=\beta.$
However, by the previous work, we know that
$\beta=\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\beta^{\prime}=\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\mathbf{L}t_{1}^{\ast}((\alpha^{A})^{-1}\circ\beta^{\prime\prime\prime})\circ\mathbf{L}t_{1}^{\ast}(\eta_{E}),$
So, from uniqueness, it follows that
$h^{\prime\prime}\circ\eta_{E}=(\alpha^{A})^{-1}\circ\beta^{\prime\prime\prime}\circ\eta_{E}.$
As $\Phi_{K}$ is fully faithful on $D^{b}_{\operatorname{coh}}$, $\eta_{E}$ must be an isomorphism, which implies $h^{\prime\prime}=(\alpha^{A})^{-1}\circ\beta^{\prime\prime\prime}$. Consequently, we have
$h=\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\mathbf{L}t^{\ast}_{1}h^{\prime\prime}=\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\mathbf{L}t_{1}^{\ast}((\alpha^{A})^{-1}\circ\beta^{\prime\prime\prime}),$
and hence, uniqueness as desired to complete the proof. ∎
Lemma 4.3.
Let $F\colon\mathcal{S}_{1}\leftrightarrows\mathcal{S}_{2}\colon G$ be an adjoint pair of exact functors between compactly generated triangulated categories. Assume $G$ preserves small coproducts. Denote by $\eta$ for the unit of the adjunction. Then the collection $\mathcal{T}$ in $\mathcal{S}_{1}$ consisting of objects $E$ for which $\eta_{E}$ is an isomorphism is a triangulated subcategory of $\mathcal{S}_{1}$ which is closed under small coproducts. Moreover, $F$ is fully faithful if, and only if, $\eta_{E}$ is an isomorphism for every $E\in\mathcal{S}_{1}^{c}$.
Proof.
It is straightforward to see that $\mathcal{T}$ is triangulated. So, we start by showing $\mathcal{T}$ is closed under small coproducts. Let $E_{i}\in\mathcal{T}$ with $i\in I$ for $I$ a set. As $F$ and $G$ both commute with small coproducts, the natural morphism $\alpha\colon\oplus_{i\in I}GF(E_{i})\to GF(\oplus_{i\in I}E_{i})$ is an isomorphism. Also, we have a commutative diagram for each $i\in I$,
$E_{i}$$GF(E_{i})$$\bigoplus_{i\in I}E_{i}$$GF(\bigoplus_{i}E_{i})$$\scriptstyle{\eta_{E_{i}}}$$\scriptstyle{\pi_{i}}$$\scriptstyle{GF(\pi_{i})}$$\scriptstyle{\eta_{\oplus_{i\in I}E_{i}}}$
where $\pi_{i}$ is the natural injection. If $\pi^{\prime}_{i}\colon GF(E_{i})\to\oplus_{i\in I}GF(E_{i})$ is the natural injection, then we know that $GF(\pi_{i})=\alpha\circ\pi_{i}^{\prime}$.
It suffices to show $\eta_{\oplus_{i\in I}E_{i}}$ is an isomorphism. Note that $\eta_{\oplus_{i\in I}E_{i}}$ is the unique morphism arising from the universal property of $\oplus_{i\in I}E_{i}$ being a coproduct of the $E_{i}$. Hence, if $GF(\oplus_{i}E_{i})$ with morphisms $GF(\pi_{i})\circ\eta_{E_{i}}$ satisfies the necessary universal property, then $\eta_{\oplus_{i\in I}E_{i}}$ is an isomorphism. Consider an $A\in\mathcal{T}$ and morphisms $f_{i}\colon E_{i}\to A$. This gives us morphisms $f_{i}\colon\eta^{-1}_{E_{i}}\colon GF(E_{i})\to A$, and hence, a unique morphism $f\colon\oplus_{i\in I}GF(E_{i})\to A$ such that $f_{i}\circ\eta^{-1}_{E_{i}}=f\circ\pi^{\prime}_{i}$ for all $i\in I$. It can be checked that $f_{i}\circ\eta^{-1}_{E_{i}}=f\circ\alpha^{-1}\circ GF(\pi_{i})$, and by uniqueness of $f$, we can see that $f\circ\alpha^{-1}$ is the only morphism from $GF(\oplus_{i\in I}E_{i})$ which can satisfy this property. Consequently, we have $GF(\oplus_{i\in I}E_{i})$ satisfying the necessary universal property to show that $\eta_{\oplus_{i\in I}E_{i}}$ is an isomorphism. So, $\mathcal{T}$ is closed under small coproducts.
Now, we check the second claim. If $F$ is fully faithful, then $\eta_{E}$ is an isomorphism for all $E\in\mathcal{S}_{1}$. So, we need to check that $\eta_{E}$ being an isomorphism for every $E\in\mathcal{S}_{1}^{c}$ characterizes fully faithfulness of $F$ on $\mathcal{S}_{1}$. In such a case, we know that $\mathcal{S}_{1}^{c}\subseteq\mathcal{T}$. However, as $\mathcal{T}$ is a triangulated subcategory of $\mathcal{S}_{1}$ that is closed under small coproducts and contains $\mathcal{S}_{1}^{c}$, it follows that $\mathcal{S}_{1}\subseteq\mathcal{T}$. This completes the proof. ∎
Proposition 4.4.
Consider ˜3.1 where $t=1_{S}$ and $S$ is semiseparated. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{2}$. Denote by $K^{\prime}$ for the kernel of the right adjoint for $\Phi_{K}$ obtained in Proposition˜3.8. Set $\eta$ to be the unit of the adjunction $\Phi_{K}$ and $\Phi_{K^{\prime}}$ on $D_{\operatorname{qc}}$. Then the following are equivalent:
  1. (1)
    $\Phi_{K}$ is fully faithful on $D_{\operatorname{qc}}$
  2. (2)
    $\eta_{E}$ is an isomorphism for all $E\in D^{-}_{\operatorname{coh}}(Y_{1})$
  3. (3)
    $\eta_{E}$ is an isomorphism for all $E\in D^{b}_{\operatorname{coh}}(Y_{1})$
  4. (4)
    $\eta_{E}$ is an isomorphism for all $E\in\operatorname{Perf}(Y_{1})$
Additionally, if $K$ is relatively perfect over $Y_{1}$, then these are equivalent to $\Phi_{K}$ is fully faithful on $D^{b}_{\operatorname{coh}}$.
Proof.
[…]
Remark 4.5.
In Proposition˜4.4, a reasonable generalization of the last claim could be to replace $D^{b}_{\operatorname{coh}}$ with $\operatorname{Perf}$ or $D^{-}_{\operatorname{coh}}$.
Theorem 4.6.
Consider ˜3.1 where $t$ is an affine morphism locally of finite tor-dimension and $S$ is semiseparated. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{2}$. If $\Phi_{K}$ is fully faithful (resp. an equivalence) on $D_{\operatorname{qc}}$, then so is $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$.
Proof.
First, we check the case of fully faithfulness. Denote by $K^{\prime}$ for the kernel of the right adjoint for $\Phi_{K}$ on $D_{\operatorname{qc}}$ that is obtained in Proposition˜3.8. Recall that Proposition˜3.8 ensures $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is left adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ on $D_{\operatorname{qc}}$. Moreover, $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ preserves small coproducts as all integral transforms do. Let $\mathcal{T}$ be the subcategory of $D_{\operatorname{qc}}(Y_{1}\times_{S}T)$ for which the unit of $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ and $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ is an isomorphism. By Lemma˜4.3, $\mathcal{T}$ is a triangulated subcategory which is closed under small coproducts. Using Proposition˜4.2, we see that $\mathbf{L}t_{1}^{\ast}\operatorname{Perf}(Y_{1})\subseteq\mathcal{T}$. However, $t_{1}$ is affine, so $\mathbf{L}t_{1}^{\ast}P$ is a compact generator of $D_{\operatorname{qc}}(Y_{1}\times_{S}T)$ if $P$ is such for $D_{\operatorname{qc}}(Y_{1})$ (see e.g. [Sta25, Tag 0BQT]). Hence, Proposition˜4.4 ensures that $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is fully faithful on $D_{\operatorname{qc}}$ as desired. Indeed, $\mathcal{T}$ containing $\mathbf{L}t_{1}^{\ast}P$ ensures it contains all the compact objects because
$\operatorname{Perf}(Y_{1}\times_{S}T)=\langle\mathbf{L}t_{1}^{\ast}P\rangle\subseteq\mathcal{T}.$
To see the the inclusion, use that any triangulated subcategory closed under small coproducts is closed under direct summands (e.g. via the ‘Eilenberg swindle’).
Lastly, we check the case of $\Phi_{K}$ being an equivalence. From the first claim, it suffices to show $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is essentially surjective. If $\Phi_{K}$ is an equivalence and is left adjoint to $\Phi_{K^{\prime}}$ on $D_{\operatorname{qc}}$, then we can arrange this data so that $\Phi_{K^{\prime}}$ is left adjoint to $\Phi_{K}$ on $D_{\operatorname{qc}}$. By Proposition˜3.8, $K^{\prime}\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$. Moreover, as $\Phi_{K^{\prime}}$ is an equivalence on $D_{\operatorname{qc}}$, it must also restrict to an equivalence on $\operatorname{Perf}$. Hence, $K^{\prime}$ must be relatively perfect over $Y_{2}$. Now, we may argue as above with the adjoint pair $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ and $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ to show the unit of this adjunction is an isomorphism. However, this implies $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is essentially surjective, which completes the proof. ∎
Corollary 4.7.
Consider ˜3.1 where $t$ is an affine morphism locally of finite tor-dimension and $S$ be semiseparated. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over each $Y_{i}$. If $\Phi_{K}$ restricts to a fully faithful functor (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$, then so does $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$.
Proof.
First, we show the case of being fully faithful. By Proposition˜4.4, $\Phi_{K}$ restricting to a fully faithful functor on $D^{b}_{\operatorname{coh}}$ is equivalent to $\Phi_{K}$ being fully faithful on $D_{\operatorname{qc}}$ because $K$ is relatively perfect over each $Y_{i}$. So, the desired claim follows from Theorem˜4.6 and a further application of Proposition˜4.4.
Next, we check the case of $\Phi_{K}$ being an equivalence on $D^{b}_{\operatorname{coh}}$. By Remark˜2.7, $\Phi_{K}$ restricts to an equivalence on $D^{b}_{\operatorname{coh}}$ implies it restricts to an equivalence on $D_{\operatorname{qc}}$. From Theorem˜4.6, $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is an equivalence on $D_{\operatorname{qc}}$. However, once again, Remark˜2.7 tells us $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ restricts to an equivalence on $D^{b}_{\operatorname{coh}}$ as desired. ∎
5. Affine variations
This section is concerned with analogs of Sections˜3 and 4 where $t$ in ˜3.1 need not be locally of finite tor-dimension. Instead, we consider the special case that $t$ is an affine morphism to the base with a mild restriction to one of the $Y_{i}\to S$.
Lemma 5.1.
Consider ˜3.1 where $t$ is an affine morphism and $f_{1}$ is locally projective. Suppose $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ is relatively perfect over $Y_{2}$. Then $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ is right adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ on $D_{\operatorname{qc}}$ where $K^{\prime}:=\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$. In particular, $K^{\prime}\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$.
Proof.
That the last claim holds follows from [Riz17, Definition 2.11 & Lemma 2.12]. So, we check the first claim. To start, as $K$ is relatively perfect over $Y_{2}$, Lemma˜3.3 tells us that $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{2}\times_{S}T$. Hence, via Lemma˜3.11, we know that $\Phi_{\widetilde{K^{\prime}}}$ is left adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ where
$\widetilde{K^{\prime}}=\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,(g^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}\times_{S}T}).$
So, it is enough to show that
$\mathbf{L}(t^{\prime})^{\ast}K^{\prime}=\mathbf{L}(t^{\prime})^{\ast}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})\cong\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,(g^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}\times_{S}T}).$
Observe, by [Sta25, Tag 08I3] and Lemma˜3.7, there exists a natural morphism
$\mathbf{L}(t^{\prime})^{\ast}$ $\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})\to\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,\mathbf{L}(t^{\prime})^{\ast}(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$
$\xrightarrow{\cong}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,(g^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}\times_{S}T}).$
However, $t$, and hence $t^{\prime}$, is an affine morphism, which ensures their derived pushforwards reflect the zero objects (see e.g. [Sta25, Tag 0BQT]). Hence, it suffices to check that the induced map
$\mathbf{R}t^{\prime}_{\ast}\mathbf{L}(t^{\prime})^{\ast}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})\to\mathbf{R}t^{\prime}_{\ast}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,\mathbf{L}(t^{\prime})^{\ast}(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$
is an isomorphism after applying $\mathbf{R}t^{\prime}_{\ast}$. Now, we have a string of isomorphisms,
$\mathbf{R}t^{\prime}_{\ast}$ $\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,\mathbf{L}(t^{\prime})^{\ast}(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$
$\cong\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,\mathbf{R}t^{\prime}_{\ast}(\mathbf{L}(t^{\prime})^{\ast}(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}}))$ ([GW23, Theorem 21.128])
$\cong\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,\mathbf{R}t^{\prime}_{\ast}\mathcal{O}_{Y_{1}\times_{S}Y_{2}\times_{S}T}\otimes^{\mathbf{L}}(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$ (Projection formula)
$\cong\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}}\otimes\mathbf{L}(f_{1}^{\prime})^{\ast}\mathbf{R}(t_{2})_{\ast}\mathcal{O}_{Y_{2}\times T})$ $\textrm{(Flat base change)}.$
Then we have a commutative square
$\mathbf{R}t^{\prime}_{\ast}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,\mathbf{L}(t^{\prime})^{\ast}(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$$\mathbf{R}t^{\prime}_{\ast}\mathbf{L}(t^{\prime})^{\ast}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$$\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})\otimes\mathbf{L}(f_{1}^{\prime})^{\ast}\mathbf{R}(t_{2})_{\ast}\mathcal{O}_{Y_{2}\times T}$$\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}}\otimes\mathbf{L}(f_{1}^{\prime})^{\ast}\mathbf{R}(t_{2})_{\ast}\mathcal{O}_{Y_{2}\times T})$$\scriptstyle{\cong}\scriptstyle{\cong}$
where the bottom vertical morphism is an isomorphism (see [Riz17, Lemma 2.15]). This gives us the required isomorphism and completes the proof. ∎
Lemma 5.2.
Consider ˜3.1 where $t$ is affine, $f_{1}$ is locally projective, and $S$ is semiseparated. Suppose $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ is relatively perfect over $Y_{2}$. Then the right adjoint of $\Phi_{K}$ restricts to an exact functor $D^{b}_{\operatorname{coh}}(Y_{2})\to D^{b}_{\operatorname{coh}}(Y_{1})$. A similar statement holds for the right adjoint of $\Phi_{\mathbf{L}(t^{\prime})K}$.
Proof.
The first claim is exactly Corollary˜3.9. By Lemma˜3.3, our hypothesis implies $\mathbf{L}(t^{\prime})K$ is relatively perfect over each $Y_{i}\times_{S}T$. So, once again, the second claim follows from Corollary˜3.9. ∎
Proposition 5.3.
Consider ˜3.1 where $t$ is affine, $f_{1}$ is locally projective, and $S$ is semiseparated. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{2}$. Denote by $K^{\prime}$ the kernel of the integral transform associated to the right adjoint of $\Phi_{K}\colon D_{\operatorname{qc}}(Y_{1})\to D_{\operatorname{qc}}(Y_{2})$ obtained in Proposition˜3.8. If $\Phi_{K}$ is fully faithful on $D_{\operatorname{qc}}$, then $\mathbf{L}t^{\ast}_{1}$ preserves the unit $\eta$ of the adjoint pair $\Phi_{K}$ and $\Phi_{K^{\prime}}$ on $D_{\operatorname{qc}}$. That is, $\mathbf{L}t_{1}^{\ast}\eta_{E}$ is the unit of $\mathbf{L}t_{1}^{\ast}E$ for the adjoint pair $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ and $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$.
Proof.
The claims follow essentially verbatim to the proof of Proposition˜4.2 if coupled with Lemmas˜5.1 and 5.2. ∎
Theorem 5.4.
Consider ˜3.1 where $t$ is affine, $f_{1}$ is locally projective, and $S$ is semiseparated. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ is relatively perfect over $Y_{2}$. If $\Phi_{K}$ is fully faithful (resp. an equivalence) on $D_{\operatorname{qc}}$, then so is $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$.
Proof.
We can argue as in Theorem˜4.6 by replacing Proposition˜3.8 with Lemma˜5.1 and Proposition˜4.2 with Proposition˜5.3. ∎
Corollary 5.5.
Consider ˜3.1 where $t$ is affine, $f_{1}$ is locally projective, and $S$ is semiseparated. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over each $Y_{i}$. If $\Phi_{K}$ restricts to a fully faithful functor (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$, then so does $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$.
Proof.
This is argued similar to Corollary˜4.7 if coupled with Theorem˜5.4. ∎
6. Applications
In this section, we study the consequences of utilizing Theorems˜4.6 and 5.4 together.
Proposition 6.1.
Let $k$ be a field of characteristic zero. Suppose $C$ and $C^{\prime}$ are geometrically integral projective curves over $k$. If $C$ and $C^{\prime}$ are Fourier–Mukai partners over $k$, then $\operatorname{Spec}(L)\times_{k}C\cong\operatorname{Spec}(L)\times_{k}C^{\prime}$ for some finite field extension $L/k$ (i.e. are $k$-forms).
Proof.
Let $\overline{k}$ be the algebraic closure of $k$. Suppose $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{k}Y_{2})$ has an associated integral transform $\Phi_{K}$ restricting to an equivalence $D^{b}_{\operatorname{coh}}(C)\to D^{b}_{\operatorname{coh}}(C^{\prime})$. Denote by $\pi\colon(Y_{1}\times_{k}Y_{2})\times_{k}\operatorname{Spec}(\overline{k})\to Y_{1}\times_{k}Y_{2}$ the natural morphism. From Theorem˜4.6, we see that $\Phi_{\mathbf{L}\pi^{\ast}K}$ restricts to an equivalence $D^{b}_{\operatorname{coh}}(C\times_{k}\operatorname{Spec}(\overline{k}))\to D^{b}_{\operatorname{coh}}(C^{\prime}\times_{k}\operatorname{Spec}(\overline{k}))$. Then $C\times_{k}\operatorname{Spec}(\overline{k})$ and $C^{\prime}\times_{k}\operatorname{Spec}(\overline{k})$ are Fourier–Mukai partners over an algebraically closed field of characteristic zero. By [CNS25, Corollary 5.4], we have that $\Phi_{\mathbf{L}\pi^{\ast}K}$ restricts to an equivalence $\operatorname{Perf}(C\times_{k}\operatorname{Spec}(\overline{k}))\to\operatorname{Perf}(C^{\prime}\times_{k}\operatorname{Spec}(\overline{k}))$. From the following facts, we know that $C\times_{k}\operatorname{Spec}(\overline{k})\cong C^{\prime}\times_{k}\operatorname{Spec}(\overline{k})$:
  • $C\times_{k}\operatorname{Spec}(\overline{k})$ is Gorenstein with ample or anti-ample canonical bundle [Bal11, Theorem 2]
  • $C\times_{k}\operatorname{Spec}(\overline{k})$ is not Gorenstein [Spe23, Theorem A]
  • $C\times_{k}\operatorname{Spec}(\overline{k})$ Gorenstein of arithmetic genus one and trivial canonical line bundle [LM14, Theorem 1.1] ($C\times_{k}\operatorname{Spec}(\overline{k})$ is integral with arithmetic genus one such that $\operatorname{deg}(\omega_{X})=0$ implies $\omega_{X}\cong\mathcal{O}_{X}$ by [Cat82, Proposition 1.11] where $\omega_{X}$ is the canonical sheaf).
Observe these are the only possible cases for projective curves over an algebraically closed field in characteristic zero. Hence, there is a finite field extension $L/k$ such that $C\times_{k}\operatorname{Spec}(L)\cong C^{\prime}\times_{k}\operatorname{Spec}(L)$ (see e.g. [Poo17, Proposition 3.2.5]), which completes the proof. ∎
Example 6.2.
There is a pair of non-isomorphic elliptic curves $E_{1}$ and $E_{2}$ over $\mathbb{Q}$ that are Fourier–Mukai partners (see [War14, Example 2.3.11]). From [Orl02, Theorem 2.2], it follows that any triangulated equivalence $D^{b}_{\operatorname{coh}}(E_{1})\to D^{b}_{\operatorname{coh}}(E_{2})$ is given by an integral transform. Note that an elliptic curve is geometrically integral and projective. By Proposition˜6.1, there is a finite field extension $L/\mathbb{Q}$ such that $E_{1}\times_{\mathbb{Q}}\operatorname{Spec}(L)\cong E_{2}\times_{\mathbb{Q}}\operatorname{Spec}(L)$.
Corollary 6.3.
Let $S$ be a separated Noetherian $\mathbb{Q}$-scheme. Suppose $f_{i}\colon Y_{i}\to S$ are proper flat morphisms whose special fibers are geometrically integral curves. Assume $f_{1}$ is locally projective. If $Y_{1}$ and $Y_{2}$ are Fourier–Mukai partners over $S$, then then genii of the special fibers of $Y_{1}$ coincide with that of $Y_{2}$.
Proof.
Let $s\in S$ be a closed point. Then the natural morphism
$\operatorname{Spec}(\kappa(s))\to S\to\operatorname{Spec}(\mathbb{Q})$
implies that $\mathbb{Q}\subseteq\kappa(s)$ (i.e. $\kappa(s)$ has characteristic zero). Note that, as $S$ is separated, $\operatorname{Spec}(\kappa(s))\to S$ is affine (see e.g. [Sta25, Tag 01SG]). Denote by $Y_{i}^{s}$ for the base change of $Y_{i}$ along $\operatorname{Spec}(\kappa(s))\to S$. Our hypothesis ensures that each $Y_{i}^{s}$ is a geometrically integral proper curve over $\kappa(s)$. Then, by Theorem˜5.4, $Y^{s}_{1}$ and $Y_{2}^{s}$ are Fourier–Mukai partners over $\kappa(s)$. Furthermore, Proposition˜6.1 guarantees that there is a finite field extension $K(s)/\kappa(s)$ such that $Y^{s}_{1}\times_{\kappa(s)}\operatorname{Spec}(K(s))$ is isomorphism to $Y^{s}_{2}\times_{\kappa(s)}\operatorname{Spec}(K(s))$. Hence, if coupled with [Sta25, Tag 0BY9 & Tag 0FD2], the genii of each $Y^{s}_{i}\times_{\kappa(s)}\operatorname{Spec}(K(s))$ coincides with $Y^{s}_{j}$ for $i\not=j$. In other words, the genii of $Y^{s}_{1}$ and $Y_{2}^{s}$ coincide, which completes the proof. ∎
Proposition 6.4.
Let $Y_{1}$ and $Y_{2}$ be proper schemes over a field $k$ (which need not be perfect). Suppose $Y_{1}$ and $Y_{2}$ are Fourier–Mukai partners. If $Y_{1}$ is geometrically regular, then so is $Y_{2}$. In other words, smoothness is a derived invariance.
Proof.
By Theorem˜4.6, $Y_{1}\times_{k}\operatorname{Spec}(L)$ and $Y_{2}\times_{k}\operatorname{Spec}(L)$ are Fourier–Mukai partners for every finitely generated field extension $L/k$. So, $Y_{1}\times_{k}\operatorname{Spec}(L)$ is regular. Then [CNS25, Corollary 5.4] implies the following string of triangulated equivalences:
$D^{b}_{\operatorname{coh}}(Y_{1}\times_{k}\operatorname{Spec}(L))\cong\operatorname{Perf}(Y_{1}\times_{k}\operatorname{Spec}(L))\cong\operatorname{Perf}(Y_{2}\times_{k}\operatorname{Spec}(L)).$
Hence, $\operatorname{Perf}(Y_{2}\times_{k}\operatorname{Spec}(L))=D^{b}_{\operatorname{coh}}(Y_{2}\times_{k}\operatorname{Spec}(L))$, telling us that $Y_{2}\times_{k}\operatorname{Spec}(L)$ is regular. In other words, $Y_{2}$ must be geometrically regular. ∎
Corollary 6.5.
Let $f_{i}\colon Y_{i}\to S$ be locally projective flat morphisms to a separated Noetherian scheme. Assume that $Y_{1}$ and $Y_{2}$ are Fourier–Mukai partners over $S$. Then $f_{1}$ is smooth if, and only if, $f_{2}$ is smooth.
Proof.
We prove only one direction as the other follows nearly verbatim. Assume $f_{1}$ is smooth. It suffices, by [Sta25, Tag 01V8], to show that all the fibers $Y_{2}\times_{S}\operatorname{Spec}(\kappa(s))$ are smooth over $\kappa(s)$ where $s\in S$ is any point. Fix $s\in S$ (perhaps, not closed). The natural morphism $t_{s}\colon\operatorname{Spec}(\mathcal{O}_{S,s})\to S$ is affine (see e.g. [Sta25, Tag 01SG]) and of finite tor-dimension (due to being flat). Hence, Theorem˜4.6 tells us the base changes of $Y_{1}$ and $Y_{2}$ along $t_{s}$ are Fourier–Mukai partners. Then Theorem˜5.4 implies these base changes remain Fourier–Mukai partners if we base change further along the natural closed immersion $\operatorname{Spec}(\kappa(p))\to\operatorname{Spec}(\mathcal{O}_{S,s})$. However, smoothness is stable under base change, so each base change of $Y_{1}$ along such morphisms remains smooth. Hence, the desired claim follows from Proposition˜6.4. ∎
Proposition 6.6.
Let $f_{i}\colon Y_{i}\to S$ be projective flat morphisms to a separated Noetherian scheme with geometrically integral fibers. Assume that $Y_{1}$ and $Y_{2}$ are Fourier–Mukai partners over $S$. Then $f_{1}$ is Cohen–Macaulay (resp. Gorenstein) if, and only if, $f_{2}$ satisfies the same condition.
Proof.
It suffices to show $f_{2}$ is Cohen–Macaulay (resp. Gorenstein) if $f_{1}$ satisfies the same condition because the proof is analogous for the converse. In fact, we argue in a similar fashion to Corollary˜6.5. Moreover, by [Sta25, Tag 045O] (resp. [Sta25, Tag 0C03]), we only need to show that all the fibers $Y_{2}\times_{S}\operatorname{Spec}(\kappa(s))$ are Cohen–Macaulay (resp. Gorenstein) over an algebraic closure of each $\kappa(s)$ where $s\in S$ is any point. Fix $s\in S$ (perhaps, not closed). The natural morphism $t_{s}\colon\operatorname{Spec}(\mathcal{O}_{S,s})\to S$ is affine (see e.g. [Sta25, Tag 01SG]) and of finite tor-dimension (due to being flat). Hence, Theorem˜4.6 tells us the base changes of $Y_{1}$ and $Y_{2}$ along $t_{s}$ are Fourier–Mukai partners. Then Theorem˜5.4 implies these base changes remain Fourier–Mukai partners if we base change further along the natural closed immersion $\operatorname{Spec}(\kappa(p))\to\operatorname{Spec}(\mathcal{O}_{S,s})$. Furthermore, if we base change further along $\overline{\kappa(s)}/\kappa(s)$ with $\overline{\kappa(s)}$ an algebraic closure of $\kappa(s)$, these base changes remain Fourier–Mukai partners via Theorem˜4.6. However, it follows (as fibers are geometrically integral) from [HLS09, Theorem 4.4], that these integral projective schemes over $\overline{\kappa(s)}$ enjoy the property that Cohen–Macaulayness (resp. Gorensteinnes) is a derived invariance. This completes the proof. ∎

### v2

@@
 1. Introduction
-1.1. Setting the stage
+1.1. What is known
 The context of our work starts with Mukai [Muk81].
@@
 cit.
-introduced the notion of integral transforms to study triangulated equivalences between derived categories associated with abelian varieties.
+introduced integral transforms to study triangulated equivalences between derived categories associated with abelian varieties.
 Over time, integral transforms have played a central role in the interplay between derived categories and algebraic geometry (e.g.
@@
 Much of this progress has focused on smooth varieties, where geometric results rely on categorical methods.
-However, extending these methods to singular varieties remains challenging, partly because bounded coherent complexes are typically not perfect.
-Thus, an important task is to develop new results or techniques that improve our understanding of integral transforms in such singular settings.
+However, extending these methods to singular varieties remains challenging, largely because bounded coherent complexes need not be perfect.
+Thus, an important task is to develop techniques that improve our understanding of integral transforms in such singular settings.
 Recent work has begun to address these difficulties.
@@
 closed points) over $S$ controls the behavior of equivalences and full faithfulness.
-In our work, we reveal additional depth to this perspective of exploiting special fibers.
-Specifically, we study the behavior of integral transforms under changes of the base scheme in more general settings that includes local algebra.
+1.2. What we do
+In this work, we deepen this perspective by systematically exploiting special fibers.
+Specifically, we study the behavior of integral transforms under base change in settings that incorporate local algebra.
 Moreover, our approach naturally extends several results in the literature to arbitrary fields, and at the same time yields new insight for singularities of fibrations.
-1.2. Preservation of adjointness
+Preservation of adjointness
 Neeman has developed a new framework called approximable triangulated categories (see e.g.
@@
 See Remark˜2.7 for details.
-For those concerned with equivalences, this allows the study of integral transforms to be reduced to their behavior on these subcategories.
+Consequently, questions about equivalences may be reduced to the behavior of integral transforms on these smaller subcategories.
 Whether an integral transform restricts to these smaller categories is now well-understood [Riz17, Bal09, DLM25].
-These results shed light to the existence of left or right adjoints to integral transforms and the conditions under which such adjoints also restrict to these subcategories.
+These results shed light on the existence of left or right adjoints to integral transforms and the conditions under which such adjoints also restrict to these subcategories.
 However, the behavior of adjointness under base change is not yet fully understood.
@@
 smooth projective curves).
-1.3. Fully faithfulness & equivalences
-Now, we have a clearer understanding of how integral transforms behave under base change.
+Fully faithfulness & equivalences
+With this understanding of adjointness under base change in hand, we turn to questions of fully faithfulness and equivalence.
 Particularly, regarding the preservation of adjointness and the restriction to smaller categories.
@@
 More recently, such questions have also been studied for smooth proper morphisms over a Noetherian base scheme (see [Kur24, Proposition/Definition 2.3]).
-However, smoothness constraints can restrict the examples which one might be interested in, e.g.
+However, smoothness assumptions restrict the range of examples one can consider, e.g.
 fibrations admitting singular fibers.
@@
 However, it remains unclear whether such canonical morphisms exist in the singular setting.
-This uncertainty is another source of motivation for our work, as we aimed to bypass the lack of these tools in singular cases using different techniques.
-The brings attention to the following, which summarizes Section˜4.
+This uncertainty provides further motivation for our work: we bypass the absence of such tools in the singular setting by developing different techniques.
+This brings attention to the following, which summarizes Section˜4.
 Theorem 1.2 (see Corollary˜4.7).
@@
 Notably, the techniques we use are independent of those appearing in [Orl02, Lemma 2.12].
-1.4. Affine variations
-Naturally, we sought analogs of Theorem˜1.2 in which $t$ need not be locally of finite tor-dimension.
-Ideally, such a framework would include morphisms from special fibers as in [HLS07, HLS09].
-Accordingly, we work in the setting of affine morphisms to the base, while imposing a mild restriction on one of the $Y_{i}\to S$.
-The following result is a slight generalization of [HLS09, Proposition 2.15].
-In particular, we perform base change along affine morphisms rather than closed immersions of closed points.
-Note that a key difference between what follows and Sections˜3 and 4 is that we impose one of the $Y_{i}\to S$ be locally projective (as opposed to proper like our earlier results).
-Theorem 1.4 (see Corollary˜5.5).
-Consider affine morphism $t\colon T\to S$ to a semiseparated Noetherian scheme.
-Suppose $Y_{1}\to S$ is flat, locally projective and $Y_{2}\to S$ is flat, proper.
-Denote by $t^{\prime}\colon Y_{1}\times_{S}Y_{2}\times_{S}T\to Y_{1}\times_{S}Y_{2}$ the natural morphism.
-Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over each $Y_{i}$.
-If $\Phi_{K}$ restricts to a fully faithful functor (resp.
-an equivalence) on $D^{b}_{\operatorname{coh}}$, then so does $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$.
-It is worthwhile to note that Theorem˜1.4 yields an independent proof of the ascending full faithfulness or equivalences in [HLS09, Proposition 2.15].
-To do so, we show in suitable cases that the unit of an adjoint pair of integral transforms is preserved under base change when the left adjoint is fully faithful (see Proposition˜5.3 and Proposition˜4.2).
-However, it is not clear to us whether this remains true when the left adjoint is not fully faithful a priori.
-This condition seems necessary (at least to us) to establish the ‘descent’ part of [HLS09, Proposition 2.15] in the case of special fibers.
-1.5. Applications
-To conclude our work, we discuss the ramifications of Theorems˜1.2 and 1.4.
+1.3. Why it matters
+To conclude our work, we discuss the ramifications of the results above.
 Our results extend the literature on schemes of finite type over algebraically closed fields to the more general setting of arbitrary fields.
-They reveal an interplay between the intuition of [HLS07, HLS09] concerning special fibers and our work on faithfully flat descent along affine morphisms.
+They reveal a new interplay between the intuition of [HLS07, HLS09] concerning special fibers and faithfully flat descent along affine morphisms.
 In particular, certain singularities are shown to be derived invariants for fibrations over Noetherian schemes, broadening the scope to include arithmetically flavored situations of potential interest for future research.
-For a pair of geometrically integral projective curves over a field $k$ of characteristic zero, Fourier–Mukai partnership implies they are $k$-forms (see Proposition˜6.1).
+For a pair of geometrically integral projective curves over a field $k$ of characteristic zero, Fourier–Mukai partnership implies they are $k$-forms (see Proposition˜5.1).
 Here, a $k$-form means there exists a finite extension $L/k$ such that their base changes along $L/k$ become isomorphic.
 The result holds for singular curves and enriches algebraically closed case of [Spe23].
-In Example˜6.2, we exhibit a pair of non-isomorphic elliptic curves over $\mathbb{Q}$ that are nevertheless $\mathbb{Q}$-forms.
-For fibrations of relative dimension one over a Noetherian $\mathbb{Q}$-scheme, Fourier–Mukai partnership forces the genera of special fibers to coincide (see Corollary˜6.3).
-Smoothness for proper schemes over an arbitrary field is likewise invariant under Fourier–Mukai partnership (see Proposition˜6.4).
+For fibrations of relative dimension one over a regular Noetherian $\mathbb{Q}$-scheme, Fourier–Mukai partnership forces the genera of special fibers to coincide (see Corollary˜5.2).
+Smoothness for proper schemes over an arbitrary field is likewise invariant under Fourier–Mukai partnership (see Proposition˜5.3).
 This subtlety arises especially over imperfect fields.
@@
 [Kol11, Remark 16]).
-In the setting of projective fibrations, smoothness of the total space is also a derived invariant (see Corollary˜6.5).
-Finally, generalizing [HLS09], we show that the Cohen–Macaulay and Gorenstein properties of a morphism remain invariant under Fourier–Mukai partnership for Noetherian base schemes (see Proposition˜6.6).
+In the setting of projective fibrations, smoothness of the total space is also a derived invariant if the base is regular (see Corollary˜5.4).
+Finally, generalizing [HLS09], we show that the Cohen–Macaulay and Gorenstein properties of a morphism remain invariant under Fourier–Mukai partnership for regular Noetherian base schemes (see Proposition˜5.5).
+In a related but distinct direction, we study a variation of Proposition˜5.5 in the case of derived equivalences.
+Particularly, we prove a version for the Gorenstein condition for arbitrary derived equivalences where the required functor need not be an integral transform.
+See Proposition˜5.6.
 Notation
@@
     $\operatorname{Perf}(X)$ is the (strictly full) subcategory of $D_{\operatorname{qc}}(X)$ consisting of the perfect complexes on $X$.
-At times if $X$ is affine, we abuse notation and write $D(R):=D_{\operatorname{qc}}(X)$ where $R:=H^{0}(X,\mathcal{O}_{X})$ are the global sections; similar conventions will occur for the other categories.
+At times if $X$ is affine, we abuse notation and write $D_{\operatorname{qc}}(R):=D_{\operatorname{qc}}(X)$ where $R:=H^{0}(X,\mathcal{O}_{X})$ are the global sections; similar conventions will occur for the other categories mentioned here and later.
 Acknowledgments.
@@
 [Sta25, Tag 08H4]).
-We will use this fact liberally.
+We will use this characterization liberally.
 While straightforward, the following is useful.
@@
 Let $\mathcal{T}$ be a triangulated category with shift functor $1]\colon\mathcal{T}\to\mathcal{T}$.
-Fix $\mathcal{S}\subseteq\mathcal{T}$.
+Fix a subcategory $\mathcal{S}\subseteq\mathcal{T}$.
 Denote by $\operatorname{add}(\mathcal{S})$ the strictly full (that is, closed under isomorphisms) subcategory of $\mathcal{T}$ consisting of all direct summands of objects of the form $\oplus_{n\in\mathbb{Z}}S_{n}^{\oplus r_{n}}[n]$ where $S_{n}\in\mathcal{S}$ for all $n$ and finitely many $r_{n}$ are nonzero.
@@
 Now, we impose $\mathcal{T}$ be a compactly generated triangulated category.
-To clarify, we require that $\mathcal{T}$ admitting all small coproducts.
+To clarify, we require that $\mathcal{T}$ admits all small coproducts.
 Denote the collection of compact objects of $\mathcal{T}$ by $\mathcal{T}^{c}$.
@@
 Moreover, $D_{\operatorname{qc}}(X)$ is compactly generated by a single object from $\operatorname{Perf}(X)$.
+See [BvdB03, Theorem 3.1.1] for details.
 The following is known for some circles, but we add it for self-containedness (see e.g. [KL15, Lemma’s 2.10 & 2.12]).
@@
 Then $t_{1},t_{2},t^{\prime}$ are locally of finite tor-dimension.
-In particular, each face of Equation˜3.1 is a tor-independent square (in the sense of [Lip09, Definition 3.10.2]).
+In particular, each face of Diagram˜3.1 is a tor-independent square (in the sense of [Lip09, Definition 3.10.2]).
 Proof.
@@
 Indeed, if needed, we can find an affine open cover of $S$ such that the base changes of $t$ along the associated open immersion are of finite tor-dimension.
-Then, for each such open immersion of a component in the affine cover, we have a similar diagram to Equation˜3.1 (i.e.
+Then, for each such open immersion of a component in the affine cover, we have a similar diagram to Diagram˜3.1 (i.e.
 a ‘base change cube’).
@@
 By base change, we know that $g_{1},g_{2},f^{\prime}_{2},f^{\prime}_{1},g_{2}^{\prime},g_{1}^{\prime}$ are proper flat morphisms.
-Consider the fibered square in Equation˜3.1 induced from the base change of $f_{1}$ along $t$.
+Consider the fibered square in Diagram˜3.1 induced from the base change of $f_{1}$ along $t$.
 It follows by [GW23, Lemma 23.27] that $t_{1}$ is of finite tor-dimension.
 A similar argument shows $t_{2}$ is of finite tor-dimension as well.
-In fact, if we work with the fibered square in Equation˜3.1 obtained by base change of $f_{2}^{\prime}$ along $t_{1}$, we see from similar reasoning that $t^{\prime}$ is locally of finite tor-dimension.
+In fact, if we work with the fibered square in Diagram˜3.1 obtained by base change of $f_{2}^{\prime}$ along $t_{1}$, we see from similar reasoning that $t^{\prime}$ is locally of finite tor-dimension.
 This completes the proof for our first claim.
@@
 Then $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ restricts to an exact functor on $D^{b}_{\operatorname{coh}}$.
-Additionally, a similar statement holds for $\operatorname{Perf}$ if $K$ is relatively perfect over $Y_{2}$.
+Similarly, if $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ is relatively perfect over $Y_{2}$, then $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ restricts to an exact functor on $\operatorname{Perf}$.
 Proof.
-We prove the case of $\Phi_{\mathbf{L}(t^{\prime})K}$ restricting to an exact functor on $\operatorname{Perf}$ because a similar argument applies for $D^{b}_{\operatorname{coh}}$.
+We only prove the case of $K$ being relatively perfect over $Y_{2}$ implying $\Phi_{\mathbf{L}(t^{\prime})K}$ restricts to an exact functor on $\operatorname{Perf}$ because a similar argument applies for $D^{b}_{\operatorname{coh}}$.
 As $\Phi_{K}(\operatorname{Perf}(Y_{1}))\subseteq\operatorname{Perf}(Y_{2})$, [DLM25, Proposition 3.2] implies
@@
 Let $K\in D_{\operatorname{qc}}(Y_{1}\times_{S}Y_{2})$.
-Then there is a natural isomorphism $\beta^{K}\colon\mathbf{L}t_{2}^{\ast}\circ\Phi_{K}\to\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\circ\mathbf{L}t_{1}^{\ast}$.
+Then there is a natural isomorphism $\beta^{K}\colon\mathbf{L}t_{2}^{\ast}\circ\Phi_{K}\to\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\circ\mathbf{L}t_{1}^{\ast}$ on $D_{\operatorname{qc}}$.
 Proof.
-This follows from the string of isomorphisms for each $E\in D_{\operatorname{qc}}(Y_{1}\times_{S}T)$:
+This follows from the string of functorial isomorphisms for each $E\in D_{\operatorname{qc}}(Y_{1}\times_{S}T)$:
 $\mathbf{L}t_{2}^{\ast}\Phi_{K}(E)$ $\cong\mathbf{L}t_{2}^{\ast}\mathbf{R}(f^{\prime}_{1})_{\ast}(\mathbf{L}(f^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}K)$
@@
 $\langle\mathbf{L}(t^{\prime})^{\ast}(\mathbf{L}f_{2}^{\ast}P_{1}\otimes^{\mathbf{L}}\mathbf{L}f_{1}^{\ast}P_{2})\rangle=\operatorname{Perf}(Y_{1}\times_{S}Y_{2}\times_{S}T).$
-However, we have an isomorphism in Equation˜3.1,
+However, we have an isomorphism in Diagram˜3.1,
 $\mathbf{L}(t^{\prime})^{\ast}(\mathbf{L}f_{2}^{\ast}P_{1}\otimes^{\mathbf{L}}\mathbf{L}f_{1}^{\ast}P_{2})$ $\cong\mathbf{L}(t^{\prime})^{\ast}\mathbf{L}f_{2}^{\ast}P_{1}\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}\mathbf{L}f_{1}^{\ast}P_{2}$
@@
 $=t_{2}(\operatorname{supp}(\mathbf{L}t_{2}^{\ast}\mathcal{H}^{j}(\Phi_{K}(E))))$ (Flatness of $t_{2}$)
-$=t_{2}(\operatorname{supp}(\mathcal{H}^{j}(\mathbf{L}t_{2}^{\ast}\Phi_{K}(E)))).$
+$=t_{2}(\operatorname{supp}(\mathcal{H}^{j}(\mathbf{L}t_{2}^{\ast}\Phi_{K}(E))))$ $\textrm{(Flatness of $t_{2}$)}.$
 Hence, the desired claim regarding bounded cohomology follows.
@@
 Then $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ is right adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ on $D_{\operatorname{qc}}$ where $K^{\prime}:=\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$.
-In particular, $K^{\prime}\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$.
+In this case, $K^{\prime}\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$.
 Proof.
@@
 $\cong\operatorname{Hom}(E,\mathbf{R}(g^{\prime}_{2})_{\ast}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,(g^{\prime}_{1})^{!}A)).$
-The desired claim follows if can find an isomorphism
+The desired claim follows if we can find an isomorphism
 $\mathbf{R}$ $(g^{\prime}_{2})_{\ast}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,(g^{\prime}_{1})^{!}A)$
 $\cong\mathbf{R}(g^{\prime}_{2})_{\ast}\bigg(\mathbf{L}(t^{\prime})^{\ast}\big(\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})\big)\otimes^{\mathbf{L}}\mathbf{L}(g^{\prime}_{1})^{\ast}A\bigg)=:\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A).$
-Note that Lemma˜3.2, as well as affineness being stable under base change, implies each vertical morphism in Equation˜3.1 are affine and locally of finite tor-dimension.
+Note that Lemma˜3.2, as well as affineness being stable under base change, implies each vertical morphism in Diagram˜3.1 are affine and locally of finite tor-dimension.
 Moreover, $(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}}\in D^{+}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ as upper shriek functors for proper morphisms preserve complexes with bounded below and coherent cohomology (see e.g.
@@
 […]
+in $D^{+}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2}\times_{S}T)$.
 By Lemma˜3.3, we see that $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ restricts to an exact functor $\operatorname{Perf}(Y_{1}\times_{S}T)\to\operatorname{Perf}(Y_{2}\times_{S}T)$.
@@
 $\mathbf{R}(g_{1}^{\prime})_{\ast}(\mathbf{L}(t^{\prime})^{\ast}K\otimes^{\mathbf{L}}\operatorname{Perf}(Y_{1}\times_{S}Y_{2}\times_{S}T))\subseteq\operatorname{Perf}(Y_{2}\times_{S}T).$
-There is another string of isomorphisms:
+There is another string of isomorphisms for all $A\in D_{\operatorname{qc}}$:
 $\mathbf{R}$ $(g^{\prime}_{2})_{\ast}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,(g^{\prime}_{1})^{!}A)$
@@
 $\cong\mathbf{R}(g^{\prime}_{2})_{\ast}\bigg(\mathbf{L}(t^{\prime})^{\ast}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})\otimes^{\mathbf{L}}\mathbf{L}(g^{\prime}_{1})^{\ast}A\bigg).$
-The last isomorphism comes from the work above, whereas the second isomorphism is [Riz17, Lemma 2.13], which completes the proof. ∎
+The last isomorphism comes from the work above, whereas the second isomorphism is [Riz17, Lemma 2.13] together with the fact that $\mathbf{L}(t^{\prime})^{\ast}K$ is $g_{1}^{\prime}$-perfect (see see Corollary˜3.6 and the discussion in Section˜2.2), which completes the proof.
+∎
 Corollary 3.9.
@@
 Let $f_{i}\colon Y_{i}\to S$ be proper flat morphisms to a Noetherian scheme where $i\in\{1,2\}$.
-Denote by $g_{i}\colon Y_{1}\times_{S}Y_{2}\to Y_{i}$ the natural morphisms.
+Denote by $f^{\prime}_{i}\colon Y_{1}\times_{S}Y_{2}\to Y_{i}$ the natural morphisms.
 Suppose $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ is relatively perfect over $Y_{1}$.
 Then $\Phi_{K}$ admits a left adjoint on $D_{\operatorname{qc}}$.
-In particular, the left adjoint is of the form $\Phi_{K^{\prime}}$ where $K^{\prime}:=\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,g^{!}_{1}\mathcal{O}_{Y_{1}})$.
+In particular, the left adjoint is of the form $\Phi_{K^{\prime}}$ where $K^{\prime}:=\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{1}})$.
 Proof.
@@
 In short, the desired claim follows from the string of natural isomorphisms for all $E\in D_{\operatorname{qc}}(Y_{2})$ and $G\in D_{\operatorname{qc}}(Y_{1})$:
-$\operatorname{Hom}$ $(E,\Phi_{K}(G))\cong\operatorname{Hom}(E,\mathbf{R}(g_{2})_{\ast}(K\otimes^{\mathbf{L}}\mathbf{L}g_{1}^{\ast}G))$ (Definition)
-$\cong\operatorname{Hom}(\mathbf{L}g_{2}^{\ast}E,K\otimes^{\mathbf{L}}\mathbf{L}g_{1}^{\ast}G)$ (Adjunction)
+$\operatorname{Hom}$ $(E,\Phi_{K}(G))\cong\operatorname{Hom}(E,\mathbf{R}(f_{2}^{\prime})_{\ast}(K\otimes^{\mathbf{L}}\mathbf{L}(f^{\prime}_{1})^{\ast}G))$ (Definition)
+$\cong\operatorname{Hom}(\mathbf{L}(f_{2}^{\prime})^{\ast}E,K\otimes^{\mathbf{L}}\mathbf{L}(f_{1}^{\prime})^{\ast}G)$ (Adjunction)
 […]
@@
 Then $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ is left adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ on $D_{\operatorname{qc}}$ where $K^{\prime}:=\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}})$.
-In particular, $K^{\prime}\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$.
+In such a case, $K^{\prime}\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$.
 Proof.
@@
 This follows if we can find an isomorphism
-$\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,(g^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}\times_{S}T})$
-$\to\mathbf{L}(t^{\prime})^{\ast}(\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}})).$
-As Lemma˜3.2 and affineness is stable under base change, it follows that each vertical morphism in Equation˜3.1 are affine and locally of finite tor-dimension.
+$\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,(g^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}\times_{S}T})\to\mathbf{L}(t^{\prime})^{\ast}(\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}})).$
+As Lemma˜3.2 and affineness is stable under base change, it follows that each vertical morphism in Diagram˜3.1 are affine and locally of finite tor-dimension.
 Also, $(g^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}}\in D^{+}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ as upper shriek functors for proper morphisms preserve complexes with bounded below and coherent cohomology (see e.g.
@@
 […]
-This completes the proof. ∎
+in $D_{\operatorname{coh}}^{+}$. This completes the proof. ∎
 4. Fully faithfulness & equivalences
@@
 Let $K\in D_{\operatorname{qc}}(Y_{1}\times_{S}Y_{2})$.
-Then there is a natural isomorphism $\alpha^{K}\colon(t_{2})_{\ast}\circ\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\to\Phi_{K}\circ(t_{1})_{\ast}$.
+Then there is a natural isomorphism $\alpha^{K}\colon\mathbf{R}(t_{2})_{\ast}\circ\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\to\Phi_{K}\circ\mathbf{R}(t_{1})_{\ast}$ on $D_{\operatorname{qc}}$.
 Proof.
 This follows from the string of natural isomorphisms for each $E\in D_{\operatorname{qc}}(Y_{1}\times_{S}T)$:
-$\mathbf{R}(t_{2})_{\ast}\circ\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}(E)$ $=\mathbf{R}(t_{2})_{\ast}\mathbf{R}(g^{\prime}_{1})_{\ast}\big(\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K\big)$
+$\mathbf{R}(t_{2})_{\ast}$ $\circ\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}(E)$
+$=\mathbf{R}(t_{2})_{\ast}\mathbf{R}(g^{\prime}_{1})_{\ast}\big(\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K\big)$
 $\cong\mathbf{R}(t_{2}\circ g^{\prime}_{1})_{\ast}\big(\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K\big)$
 $\cong\mathbf{R}(f_{1}^{\prime}\circ t^{\prime})_{\ast}\big(\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K\big)$
-$\cong\mathbf{R}(f_{1}^{\prime})_{\ast}\bigg(\mathbf{R}(t^{\prime})_{\ast}\big(\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K\big)\bigg)$
+$\cong\mathbf{R}(f_{1}^{\prime})_{\ast}\bigg(\mathbf{R}(t^{\prime})_{\ast}\big(\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K\big)\bigg)$ ([GW23, Proposition 22.84])
 $\cong\mathbf{R}(f_{1}^{\prime})_{\ast}\bigg(\mathbf{R}(t^{\prime})_{\ast}\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}K\bigg)$
@@
 Let $\beta\colon\mathbf{L}t^{\ast}_{1}E\to\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$ be a morphism in $D_{\operatorname{qc}}(Y_{1}\times_{S}T)$ for any $A\in D_{\operatorname{qc}}(Y_{2}\times_{S}T)$.
-Using that $\mathbf{L}t_{1}^{\ast}$ is left adjoint to $\mathbf{R}(t_{1})_{\ast}$ on $D_{\operatorname{qc}}$, adjunction tells us there exists a unique morphism $\beta^{\prime}$ making the following diagram commute,
+Using that $\mathbf{L}t_{1}^{\ast}$ is left adjoint to $\mathbf{R}(t_{1})_{\ast}$ on $D_{\operatorname{qc}}$, adjunction tells us there exists a unique morphism $\beta^{\prime}$ of the form $\mathbf{L}t^{\ast}_{1}(\beta^{\prime\prime})$ where $\beta^{\prime\prime}\colon E\to\mathbf{R}(t_{1})_{\ast}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$, making the following diagram commute,
 $\mathbf{L}t^{\ast}_{1}E$$\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$$\mathbf{L}t_{1}^{\ast}\mathbf{R}(t_{1})_{\ast}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$$\scriptstyle{\beta}$$\scriptstyle{\beta^{\prime}}$$\scriptstyle{\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}}$
 where $\xi^{t_{1}}$ is the counit of derived pullback and pushforward along $t_{1}$.
-Now, $\beta^{\prime}$ is of the form $\mathbf{L}t^{\ast}_{1}(\beta^{\prime\prime})$ where $\beta^{\prime\prime}\colon E\to\mathbf{R}(t_{1})_{\ast}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$.
 Furthermore, by the universal property of the counit, the morphism $\beta^{\prime\prime}$ is the unique morphism such that $\beta=\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\mathbf{L}t^{\ast}_{1}(\beta^{\prime\prime})$.
@@
 […]
-which commutes. In other words,
+which commutes. Pulling back this diagram via $\mathbf{L}t_{1}^{\ast}$, we obtain
 $\beta^{\prime}=\mathbf{L}t_{1}^{\ast}((\alpha^{A})^{-1}\circ\beta^{\prime\prime\prime})\circ\mathbf{L}t_{1}^{\ast}(\eta_{E}),$
 which tells us $\beta$ factors through $\mathbf{L}t_{1}^{\ast}(\eta_{E})$.
-It remains to show that this factorization is unique.
+It remains to show that the above factorization for $\beta^{\prime}$ is unique.
 So, suppose there is a morphism $h\colon\mathbf{L}t_{1}^{\ast}\Phi_{K^{\prime}}\Phi_{K}(E)\to\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$ such that $\beta=h\circ\mathbf{L}t_{1}^{\ast}(\eta_{E})$.
@@
 $\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\mathbf{L}t^{\ast}_{1}(h^{\prime\prime}\circ\eta_{E})=(\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\mathbf{L}t^{\ast}_{1}(h^{\prime\prime}))\circ\mathbf{L}t^{\ast}_{1}(\eta_{E})=h\circ\mathbf{L}t^{\ast}_{1}(\eta_{E})=\beta.$
-However, by the previous work, we know that
+However, by the work above, we know that
 $\beta=\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\beta^{\prime}=\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\mathbf{L}t_{1}^{\ast}((\alpha^{A})^{-1}\circ\beta^{\prime\prime\prime})\circ\mathbf{L}t_{1}^{\ast}(\eta_{E}),$
@@
 Assume $G$ preserves small coproducts.
-Denote by $\eta$ for the unit of the adjunction.
+Denote by $\eta\colon 1\to GF$ for the unit of the adjunction.
 Then the collection $\mathcal{T}$ in $\mathcal{S}_{1}$ consisting of objects $E$ for which $\eta_{E}$ is an isomorphism is a triangulated subcategory of $\mathcal{S}_{1}$ which is closed under small coproducts.
@@
 Consider an $A\in\mathcal{T}$ and morphisms $f_{i}\colon E_{i}\to A$.
-This gives us morphisms $f_{i}\colon\eta^{-1}_{E_{i}}\colon GF(E_{i})\to A$, and hence, a unique morphism $f\colon\oplus_{i\in I}GF(E_{i})\to A$ such that $f_{i}\circ\eta^{-1}_{E_{i}}=f\circ\pi^{\prime}_{i}$ for all $i\in I$.
+This gives us morphisms $f_{i}\circ\eta^{-1}_{E_{i}}\colon GF(E_{i})\to A$, and hence, a unique morphism $f\colon\oplus_{i\in I}GF(E_{i})\to A$ such that $f_{i}\circ\eta^{-1}_{E_{i}}=f\circ\pi^{\prime}_{i}$ for all $i\in I$.
 It can be checked that $f_{i}\circ\eta^{-1}_{E_{i}}=f\circ\alpha^{-1}\circ GF(\pi_{i})$, and by uniqueness of $f$, we can see that $f\circ\alpha^{-1}$ is the only morphism from $GF(\oplus_{i\in I}E_{i})$ which can satisfy this property.
@@
 Denote by $K^{\prime}$ for the kernel of the right adjoint for $\Phi_{K}$ obtained in Proposition˜3.8.
-Set $\eta$ to be the unit of the adjunction $\Phi_{K}$ and $\Phi_{K^{\prime}}$ on $D_{\operatorname{qc}}$.
+Set $\eta\colon 1\to\Phi_{K}\Phi_{K^{\prime}}$ to be the unit of the adjunction $\Phi_{K}$ and $\Phi_{K^{\prime}}$ on $D_{\operatorname{qc}}$.
 Then the following are equivalent:
@@
 From the first claim, it suffices to show $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is essentially surjective.
-If $\Phi_{K}$ is an equivalence and is left adjoint to $\Phi_{K^{\prime}}$ on $D_{\operatorname{qc}}$, then we can arrange this data so that $\Phi_{K^{\prime}}$ is left adjoint to $\Phi_{K}$ on $D_{\operatorname{qc}}$.
+If $\Phi_{K}$ is an equivalence and is left adjoint to $\Phi_{K^{\prime}}$ on $D_{\operatorname{qc}}$, then one can show that $\Phi_{K^{\prime}}$ is also left adjoint to $\Phi_{K}$ on $D_{\operatorname{qc}}$.
 By Proposition˜3.8, $K^{\prime}\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$.
 Moreover, as $\Phi_{K^{\prime}}$ is an equivalence on $D_{\operatorname{qc}}$, it must also restrict to an equivalence on $\operatorname{Perf}$.
-Hence, $K^{\prime}$ must be relatively perfect over $Y_{2}$.
-Now, we may argue as above with the adjoint pair $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ and $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ to show the unit of this adjunction is an isomorphism.
+Hence, $K^{\prime}$ must be relatively perfect over $Y_{1}$.
+Now, by the above $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ is fully faithful and by Proposition 4.4 the unit of the adjoint pair $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ and $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is an isomorphism.
 However, this implies $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is essentially surjective, which completes the proof.
@@
 Next, we check the case of $\Phi_{K}$ being an equivalence on $D^{b}_{\operatorname{coh}}$.
-By Remark˜2.7, $\Phi_{K}$ restricts to an equivalence on $D^{b}_{\operatorname{coh}}$ implies it restricts to an equivalence on $D_{\operatorname{qc}}$.
+By Remark˜2.7, $\Phi_{K}$ restricts to an equivalence on $D^{b}_{\operatorname{coh}}$ implies $\Phi_{K}$ is an equivalence on $D_{\operatorname{qc}}$.
 From Theorem˜4.6, $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is an equivalence on $D_{\operatorname{qc}}$.
@@
 ∎
-5. Affine variations
-This section is concerned with analogs of Sections˜3 and 4 where $t$ in ˜3.1 need not be locally of finite tor-dimension.
-Instead, we consider the special case that $t$ is an affine morphism to the base with a mild restriction to one of the $Y_{i}\to S$.
-Lemma 5.1.
-Consider ˜3.1 where $t$ is an affine morphism and $f_{1}$ is locally projective.
-Suppose $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ is relatively perfect over $Y_{2}$.
-Then $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ is right adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ on $D_{\operatorname{qc}}$ where $K^{\prime}:=\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$.
-In particular, $K^{\prime}\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$.
-Proof.
-That the last claim holds follows from [Riz17, Definition 2.11 & Lemma 2.12].
-So, we check the first claim.
-To start, as $K$ is relatively perfect over $Y_{2}$, Lemma˜3.3 tells us that $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{2}\times_{S}T$.
-Hence, via Lemma˜3.11, we know that $\Phi_{\widetilde{K^{\prime}}}$ is left adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ where
-$\widetilde{K^{\prime}}=\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,(g^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}\times_{S}T}).$
-So, it is enough to show that
-$\mathbf{L}(t^{\prime})^{\ast}K^{\prime}=\mathbf{L}(t^{\prime})^{\ast}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})\cong\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,(g^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}\times_{S}T}).$
-Observe, by [Sta25, Tag 08I3] and Lemma˜3.7, there exists a natural morphism
-$\mathbf{L}(t^{\prime})^{\ast}$ $\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})\to\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,\mathbf{L}(t^{\prime})^{\ast}(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$
-$\xrightarrow{\cong}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,(g^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}\times_{S}T}).$
-However, $t$, and hence $t^{\prime}$, is an affine morphism, which ensures their derived pushforwards reflect the zero objects (see e.g.
-[Sta25, Tag 0BQT]).
-Hence, it suffices to check that the induced map
-$\mathbf{R}t^{\prime}_{\ast}\mathbf{L}(t^{\prime})^{\ast}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})\to\mathbf{R}t^{\prime}_{\ast}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,\mathbf{L}(t^{\prime})^{\ast}(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$
-is an isomorphism after applying $\mathbf{R}t^{\prime}_{\ast}$. Now, we have a string of isomorphisms,
-$\mathbf{R}t^{\prime}_{\ast}$ $\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,\mathbf{L}(t^{\prime})^{\ast}(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$
-$\cong\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,\mathbf{R}t^{\prime}_{\ast}(\mathbf{L}(t^{\prime})^{\ast}(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}}))$ ([GW23, Theorem 21.128])
-$\cong\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,\mathbf{R}t^{\prime}_{\ast}\mathcal{O}_{Y_{1}\times_{S}Y_{2}\times_{S}T}\otimes^{\mathbf{L}}(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$ (Projection formula)
-$\cong\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}}\otimes\mathbf{L}(f_{1}^{\prime})^{\ast}\mathbf{R}(t_{2})_{\ast}\mathcal{O}_{Y_{2}\times T})$ $\textrm{(Flat base change)}.$
-Then we have a commutative square
-$\mathbf{R}t^{\prime}_{\ast}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,\mathbf{L}(t^{\prime})^{\ast}(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$$\mathbf{R}t^{\prime}_{\ast}\mathbf{L}(t^{\prime})^{\ast}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$$\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})\otimes\mathbf{L}(f_{1}^{\prime})^{\ast}\mathbf{R}(t_{2})_{\ast}\mathcal{O}_{Y_{2}\times T}$$\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}}\otimes\mathbf{L}(f_{1}^{\prime})^{\ast}\mathbf{R}(t_{2})_{\ast}\mathcal{O}_{Y_{2}\times T})$$\scriptstyle{\cong}\scriptstyle{\cong}$
-where the bottom vertical morphism is an isomorphism (see [Riz17, Lemma 2.15]). This gives us the required isomorphism and completes the proof. ∎
-Lemma 5.2.
-Consider ˜3.1 where $t$ is affine, $f_{1}$ is locally projective, and $S$ is semiseparated.
-Suppose $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ is relatively perfect over $Y_{2}$.
-Then the right adjoint of $\Phi_{K}$ restricts to an exact functor $D^{b}_{\operatorname{coh}}(Y_{2})\to D^{b}_{\operatorname{coh}}(Y_{1})$.
-A similar statement holds for the right adjoint of $\Phi_{\mathbf{L}(t^{\prime})K}$.
-Proof.
-The first claim is exactly Corollary˜3.9.
-By Lemma˜3.3, our hypothesis implies $\mathbf{L}(t^{\prime})K$ is relatively perfect over each $Y_{i}\times_{S}T$.
-So, once again, the second claim follows from Corollary˜3.9.
-∎
-Proposition 5.3.
-Consider ˜3.1 where $t$ is affine, $f_{1}$ is locally projective, and $S$ is semiseparated.
-Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{2}$.
-Denote by $K^{\prime}$ the kernel of the integral transform associated to the right adjoint of $\Phi_{K}\colon D_{\operatorname{qc}}(Y_{1})\to D_{\operatorname{qc}}(Y_{2})$ obtained in Proposition˜3.8.
-If $\Phi_{K}$ is fully faithful on $D_{\operatorname{qc}}$, then $\mathbf{L}t^{\ast}_{1}$ preserves the unit $\eta$ of the adjoint pair $\Phi_{K}$ and $\Phi_{K^{\prime}}$ on $D_{\operatorname{qc}}$.
-That is, $\mathbf{L}t_{1}^{\ast}\eta_{E}$ is the unit of $\mathbf{L}t_{1}^{\ast}E$ for the adjoint pair $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ and $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$.
-Proof.
-The claims follow essentially verbatim to the proof of Proposition˜4.2 if coupled with Lemmas˜5.1 and 5.2. ∎
-Theorem 5.4.
-Consider ˜3.1 where $t$ is affine, $f_{1}$ is locally projective, and $S$ is semiseparated.
-Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ is relatively perfect over $Y_{2}$.
-If $\Phi_{K}$ is fully faithful (resp.
-an equivalence) on $D_{\operatorname{qc}}$, then so is $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$.
-Proof.
-We can argue as in Theorem˜4.6 by replacing Proposition˜3.8 with Lemma˜5.1 and Proposition˜4.2 with Proposition˜5.3. ∎
-Corollary 5.5.
-Consider ˜3.1 where $t$ is affine, $f_{1}$ is locally projective, and $S$ is semiseparated.
-Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over each $Y_{i}$.
-If $\Phi_{K}$ restricts to a fully faithful functor (resp.
-an equivalence) on $D^{b}_{\operatorname{coh}}$, then so does $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$.
-Proof.
-This is argued similar to Corollary˜4.7 if coupled with Theorem˜5.4. ∎
 5. Applications
-In this section, we study the consequences of utilizing Theorems˜4.6 and 5.4 together.
+In this section, we study the consequences of Theorem˜4.6. By ‘genus’ of a curve, we mean ‘arithmetic’(see e.g. [Sta25, Tag 0BY6]).
 Proposition 5.1.
@@
 Let $\overline{k}$ be the algebraic closure of $k$.
-Suppose $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{k}Y_{2})$ has an associated integral transform $\Phi_{K}$ restricting to an equivalence $D^{b}_{\operatorname{coh}}(C)\to D^{b}_{\operatorname{coh}}(C^{\prime})$.
-Denote by $\pi\colon(Y_{1}\times_{k}Y_{2})\times_{k}\operatorname{Spec}(\overline{k})\to Y_{1}\times_{k}Y_{2}$ the natural morphism.
+Suppose $K\in D^{b}_{\operatorname{coh}}(C\times_{k}C^{\prime})$ has an associated integral transform $\Phi_{K}$ restricting to an equivalence $D^{b}_{\operatorname{coh}}(C)\to D^{b}_{\operatorname{coh}}(C^{\prime})$.
+Denote by $\pi\colon(C\times_{k}C^{\prime})\times_{k}\operatorname{Spec}(\overline{k})\to C\times_{k}C^{\prime}$ the natural morphism.
 From Theorem˜4.6, we see that $\Phi_{\mathbf{L}\pi^{\ast}K}$ restricts to an equivalence $D^{b}_{\operatorname{coh}}(C\times_{k}\operatorname{Spec}(\overline{k}))\to D^{b}_{\operatorname{coh}}(C^{\prime}\times_{k}\operatorname{Spec}(\overline{k}))$.
@@
 By [CNS25, Corollary 5.4], we have that $\Phi_{\mathbf{L}\pi^{\ast}K}$ restricts to an equivalence $\operatorname{Perf}(C\times_{k}\operatorname{Spec}(\overline{k}))\to\operatorname{Perf}(C^{\prime}\times_{k}\operatorname{Spec}(\overline{k}))$.
-From the following facts, we know that $C\times_{k}\operatorname{Spec}(\overline{k})\cong C^{\prime}\times_{k}\operatorname{Spec}(\overline{k})$:
+Consequently, we know that $C\times_{k}\operatorname{Spec}(\overline{k})\cong C^{\prime}\times_{k}\operatorname{Spec}(\overline{k})$ in the following cases:
   • $C\times_{k}\operatorname{Spec}(\overline{k})$ is Gorenstein with ample or anti-ample canonical bundle [Bal11, Theorem 2]
   • $C\times_{k}\operatorname{Spec}(\overline{k})$ is not Gorenstein [Spe23, Theorem A]
-  • $C\times_{k}\operatorname{Spec}(\overline{k})$ Gorenstein of arithmetic genus one and trivial canonical line bundle [LM14, Theorem 1.1] ($C\times_{k}\operatorname{Spec}(\overline{k})$ is integral with arithmetic genus one such that $\operatorname{deg}(\omega_{X})=0$ implies $\omega_{X}\cong\mathcal{O}_{X}$ by [Cat82, Proposition 1.11] where $\omega_{X}$ is the canonical sheaf).
-Observe these are the only possible cases for projective curves over an algebraically closed field in characteristic zero.
+  • $C\times_{k}\operatorname{Spec}(\overline{k})$ Gorenstein of genus one and trivial canonical line bundle [LM14, Theorem 1.1] ($C\times_{k}\operatorname{Spec}(\overline{k})$ is integral with genus one such that $\operatorname{deg}(\omega_{X})=0$ implies $\omega_{X}\cong\mathcal{O}_{X}$ by [Cat82, Proposition 1.11] where $\omega_{X}$ is the canonical sheaf).
+Observe these are the only possible cases for projective curves over an algebraically closed field in characteristic zero (see e.g.
+the two paragraphs above [Spe23, Theorem A] or use [Cat82, Theorem B, Proposition 1.8 and Proposition 1.11]).
 Hence, there is a finite field extension $L/k$ such that $C\times_{k}\operatorname{Spec}(L)\cong C^{\prime}\times_{k}\operatorname{Spec}(L)$ (see e.g.
@@
 ∎
-Example 6.2.
-There is a pair of non-isomorphic elliptic curves $E_{1}$ and $E_{2}$ over $\mathbb{Q}$ that are Fourier–Mukai partners (see [War14, Example 2.3.11]).
-From [Orl02, Theorem 2.2], it follows that any triangulated equivalence $D^{b}_{\operatorname{coh}}(E_{1})\to D^{b}_{\operatorname{coh}}(E_{2})$ is given by an integral transform.
-Note that an elliptic curve is geometrically integral and projective.
-By Proposition˜6.1, there is a finite field extension $L/\mathbb{Q}$ such that $E_{1}\times_{\mathbb{Q}}\operatorname{Spec}(L)\cong E_{2}\times_{\mathbb{Q}}\operatorname{Spec}(L)$.
 Corollary 5.2.
-Let $S$ be a separated Noetherian $\mathbb{Q}$-scheme.
+Let $S$ be a quasi-compact separated regular $\mathbb{Q}$-scheme.
 Suppose $f_{i}\colon Y_{i}\to S$ are proper flat morphisms whose special fibers are geometrically integral curves.
@@
 Our hypothesis ensures that each $Y_{i}^{s}$ is a geometrically integral proper curve over $\kappa(s)$.
-Then, by Theorem˜5.4, $Y^{s}_{1}$ and $Y_{2}^{s}$ are Fourier–Mukai partners over $\kappa(s)$.
-Furthermore, Proposition˜6.1 guarantees that there is a finite field extension $K(s)/\kappa(s)$ such that $Y^{s}_{1}\times_{\kappa(s)}\operatorname{Spec}(K(s))$ is isomorphism to $Y^{s}_{2}\times_{\kappa(s)}\operatorname{Spec}(K(s))$.
+Then, by Theorem˜4.6, $Y^{s}_{1}$ and $Y_{2}^{s}$ are Fourier–Mukai partners over $\kappa(s)$.
+Indeed, the hypothesis on $S$ ensures it is Noetherian, and so the closed immersion $\operatorname{Spec}(\kappa(s))\to S$ is locally of finite tor-dimension (see e.g.
+[GW23, Remark 25.39(2)]).
+Furthermore, Proposition˜5.1 guarantees that there is a finite field extension $K(s)/\kappa(s)$ such that $Y^{s}_{1}\times_{\kappa(s)}\operatorname{Spec}(K(s))$ is isomorphism to $Y^{s}_{2}\times_{\kappa(s)}\operatorname{Spec}(K(s))$.
 Hence, if coupled with [Sta25, Tag 0BY9 & Tag 0FD2], the genii of each $Y^{s}_{i}\times_{\kappa(s)}\operatorname{Spec}(K(s))$ coincides with $Y^{s}_{j}$ for $i\not=j$.
@@
 If $Y_{1}$ is geometrically regular, then so is $Y_{2}$.
-In other words, smoothness is a derived invariance.
+In other words, smoothness is an invariance for Fourier-Mukai partners.
 Proof.
@@
 Corollary 5.4.
-Let $f_{i}\colon Y_{i}\to S$ be locally projective flat morphisms to a separated Noetherian scheme.
+Let $f_{i}\colon Y_{i}\to S$ be locally projective flat morphisms to a quasi-compact separated regular scheme.
 Assume that $Y_{1}$ and $Y_{2}$ are Fourier–Mukai partners over $S$.
@@
 It suffices, by [Sta25, Tag 01V8], to show that all the fibers $Y_{2}\times_{S}\operatorname{Spec}(\kappa(s))$ are smooth over $\kappa(s)$ where $s\in S$ is any point.
-Fix $s\in S$ (perhaps, not closed).
+Fix a point $s\in S$ (which need not be closed).
 The natural morphism $t_{s}\colon\operatorname{Spec}(\mathcal{O}_{S,s})\to S$ is affine (see e.g.
@@
 Hence, Theorem˜4.6 tells us the base changes of $Y_{1}$ and $Y_{2}$ along $t_{s}$ are Fourier–Mukai partners.
-Then Theorem˜5.4 implies these base changes remain Fourier–Mukai partners if we base change further along the natural closed immersion $\operatorname{Spec}(\kappa(p))\to\operatorname{Spec}(\mathcal{O}_{S,s})$.
+A further application Theorem˜4.6 implies these base changes remain Fourier–Mukai partners if we base change further along the natural closed immersion $\operatorname{Spec}(\kappa(p))\to\operatorname{Spec}(\mathcal{O}_{S,s})$.
+Indeed, $S$ being regular implies $\mathcal{O}_{S,s}$ is as well, and so the closed immersion from $\operatorname{Spec(\kappa(p))}$ is of finite tor-dimension (see e.g.
+[GW23, Remark 25.39(2)]).
 However, smoothness is stable under base change, so each base change of $Y_{1}$ along such morphisms remains smooth.
-Hence, the desired claim follows from Proposition˜6.4.
+Hence, the desired claim follows from Proposition˜5.3.
 ∎
 Proposition 5.5.
-Let $f_{i}\colon Y_{i}\to S$ be projective flat morphisms to a separated Noetherian scheme with geometrically integral fibers.
+Let $f_{i}\colon Y_{i}\to S$ be projective flat morphisms with geometrically integral fibers.
+Suppose $S$ is quasi-compact, separated, and regular.
 Assume that $Y_{1}$ and $Y_{2}$ are Fourier–Mukai partners over $S$.
@@
 Gorenstein) if $f_{1}$ satisfies the same condition because the proof is analogous for the converse.
-In fact, we argue in a similar fashion to Corollary˜6.5.
+In fact, we argue in a similar fashion to Corollary˜5.4.
 Moreover, by [Sta25, Tag 045O] (resp.
@@
 Hence, Theorem˜4.6 tells us the base changes of $Y_{1}$ and $Y_{2}$ along $t_{s}$ are Fourier–Mukai partners.
-Then Theorem˜5.4 implies these base changes remain Fourier–Mukai partners if we base change further along the natural closed immersion $\operatorname{Spec}(\kappa(p))\to\operatorname{Spec}(\mathcal{O}_{S,s})$.
+Then Theorem˜4.6 implies these base changes remain Fourier–Mukai partners if we base change further along the natural closed immersion $\operatorname{Spec}(\kappa(p))\to\operatorname{Spec}(\mathcal{O}_{S,s})$.
 Furthermore, if we base change further along $\overline{\kappa(s)}/\kappa(s)$ with $\overline{\kappa(s)}$ an algebraic closure of $\kappa(s)$, these base changes remain Fourier–Mukai partners via Theorem˜4.6.
@@
 ∎
+Under additional conditions, we can improve Proposition˜5.5 in the Gorenstein case.
+Proposition 5.6.
+Let $f_{i}\colon Y_{i}\to S$ be projective flat morphisms to a separated Noetherian scheme, where $Y_{1}$ and $Y_{2}$ are of finite Krull dimension, satisfying the resolution property (see e.g.
+[Sta25, Tag 0F86]).
+Assume that $Y_{1}$ and $Y_{2}$ are derived equivalences (e.g.
+the equivalence need not arise from an integral transform).
+Then $f_{1}$ is Gorenstein if, and only if, $f_{2}$ satisfies the same condition.
+Proof.
+We prove the case where $f_{1}$ is Gorenstein implies that of $f_{2}$ because the proof for the converse holds verabtim.
+Denote by $D^{b}_{\operatorname{fid}}(Y_{2})$ the full triangulated subcategory in $D^{b}_{\operatorname{coh}}(Y_{2})$ of complexes with bounded cohomology and with a bounded injective resolution (i.e.
+can be represented by a complex with finitely many nonzero components where each is injective $\mathcal{O}_{X}$-module).
+By [Har66, Subchapter V.2], the dualizing sheaf $f_{2}^{!}\mathcal{O}_{S}$ is contained in $D^{b}_{\operatorname{fid}}(Y_{2})$.
+Furthermore, [Sta25, Tag 0A89] tells us that there is an equivalence
+(5.1) $D:=\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(-,f_{2}^{!}\mathcal{O}_{S})\colon D^{b}_{\operatorname{coh}}(Y_{2})\xrightarrow{\sim}D^{b}_{\operatorname{coh}}(Y_{2})^{\operatorname{op}}$
+satisfying $D\circ D=\operatorname{id}$. Moreover, this equivalence restricts to an equivalence [Har66, Proposition V.2.6 a)],
+(5.2) $D\colon\operatorname{Perf}(Y_{2})\xrightarrow{\sim}D^{b}_{\operatorname{fid}}(Y_{2})^{\operatorname{op}}$
+So, we have the following intrinsic characterization of $D^{b}_{\operatorname{fid}}(Y_{1})$ in $D^{b}_{\operatorname{coh}}(Y_{1})$,
+(5.3) $D^{b}_{\operatorname{fid}}(Y_{1})=\{B\in D^{b}_{\operatorname{coh}}(Y_{1})\mid\operatorname{Hom}(A,B[p])=0,\text{for}\ 0\ll|p|\text{ and all }A\in D^{b}_{\operatorname{coh}}(Y_{1})\}.$
+Indeed, by [Orl06, Proposition 1.11], we have a dual characterization for perfect complexes in $D^{b}_{\operatorname{coh}}(Y_{1})$,
+$\operatorname{Perf}(Y_{1})=\{B\in D^{b}_{\operatorname{coh}}(Y_{1})\mid\operatorname{Hom}(B,A[p])=0,\text{for}\ 0\ll|p|\text{ and all }A\in D^{b}_{\operatorname{coh}}(Y_{1})\},$
+because $Y_{1}$ is separated Noetherian of finite Krull dimension and satisfies the resolution property (Orlov calls this (ELF) scheme - note that the proof in loc.
+cit.
+does not use that the scheme is defined over a field).
+This characterization of $\operatorname{Perf}$, together with the equivalences (5.1) and (5.2), implies (5.3).
+Recall that $Y_{1}$ over $S$ is Gorenstein if, and only if, the dualizing sheaf $f_{1}^{!}\mathcal{O}_{S}$ is invertible (see e.g.
+[Har66, Exercise V.9.7]).
+In particular, $f_{1}^{!}\mathcal{O}$ is a perfect complex.
+If now $\Phi\colon D_{\operatorname{qc}}(Y_{1})\to D_{\operatorname{qc}}(Y_{2})$ is an equivalence, then $\Phi(f_{1}^{!}\mathcal{O})$ is a perfect complex in $D_{\operatorname{qc}}(Y_{2})$.
+Furthermore, as $\Phi$ restrict to an equivalence on $D^{b}_{\operatorname{coh}}$, we see by the characterization (5.3) that $\Phi(f_{1}^{!}\mathcal{O})$ is contained in $D^{b}_{\operatorname{fid}}(Y_{2})$.
+Hence, by (5.2), we see that $\Phi(f_{1}^{!}\mathcal{O})$ is of the form
+$\Phi(f_{1}^{!}\mathcal{O})\cong D(P)=\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(P,f_{2}^{!}\mathcal{O})\cong P^{\vee}\otimes^{\mathbf{L}}f_{2}^{!}\mathcal{O}\in D^{b}_{\operatorname{fid}}(Y_{2})$
+for some $P$ in $\operatorname{Perf}(Y_{2})$.
+However, $\Phi(f_{1}^{!}\mathcal{O})$, and so, $P^{\vee}\otimes^{\mathbf{L}}f_{2}^{!}\mathcal{O}$ is in $\operatorname{Perf}(Y_{2})$.
+Hence, it follows that $f_{2}^{!}\mathcal{O}$ is also in $\operatorname{Perf}(Y_{2})$.
+Moreover, since $f_{2}^{!}\mathcal{O}^{\vee}\otimes^{\mathbf{L}}f_{2}^{!}\mathcal{O}\cong\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(f_{2}^{!}\mathcal{O},f_{2}^{!}\mathcal{O})\cong\mathcal{O}_{Y_{2}}$ by the definition of a dualizing complex, it is easy to deduce that $f_{2}^{!}\mathcal{O}$ is a shift of a line bundle of $Y_{2}$ (see e.g.
+[Har66, Lemma V.3.3]).
+Again, by [Har66, Exercise V.9.7], we conclude that $Y_{2}$ over $S$ is Gorenstein, which completes the proof.
+∎

### v3

@@
 Abstract.
-We study the behavior of integral transforms under various forms of base change.
-In particular, we establish a yoga of local algebra and fibers to test for derived equivalences via integral transforms.
-This partially generalizes a result of Orlov to singular varieties and strengthens many results in the literature to arbitrary fields.
-Additionally, it provides new insight into fibrations and their singularities arising in more arithmetically flavored contexts (e.g.
+We study the behavior of integral transforms under base change.
+In particular, we establish a yoga of local algebra and fibers to test for derived equivalences or fully faithfulness via integral transforms.
+This generalizes a result of Orlov to singular varieties and strengthens several results in the literature by allowing arbitrary base fields.
+Additionally, it provides new insight into fibrations and their singularities in arithmetic settings (e.g.
 projective and flat schemes over a DVR).
 Key words and phrases:
-Fourier–Mukai partners, integral transforms, singular varieties, derived categories
+Fourier–Mukai partners, integral transforms, singular varieties, derived categories, fully faithfulness
 2020 Mathematics Subject Classification:
@@
 1.1. What is known
-The context of our work starts with Mukai [Muk81].
-Loc.
-cit.
-introduced integral transforms to study triangulated equivalences between derived categories associated with abelian varieties.
-Over time, integral transforms have played a central role in the interplay between derived categories and algebraic geometry (e.g.
-[Bri02, Kaw02]), with further applications in string theory [AHR05, ASS05].
-Much of this progress has focused on smooth varieties, where geometric results rely on categorical methods.
+The context for our work begins with Mukai [MUK81].
+In that work, integral transforms were introduced to study triangulated equivalences between derived categories of abelian varieties.
+Since then, integral transforms have played a central role in the interplay between derived categories and algebraic geometry (e.g.
+[BRI02, KAW02]).
+Much of this progress concerns smooth varieties, where geometric results can be obtained using categorical methods.
 However, extending these methods to singular varieties remains challenging, largely because bounded coherent complexes need not be perfect.
-Thus, an important task is to develop techniques that improve our understanding of integral transforms in such singular settings.
+Thus, an important task is to develop techniques for studying integral transforms in singular settings.
 Recent work has begun to address these difficulties.
-Criteria determining when an integral transform is fully faithful or an equivalence have been established for varieties over algebraically closed fields, generalizing results from the smooth case [HLS07, HLS09] and building upon foundational work of Bondal–Orlov [AB95].
+Criteria determining when an integral transform is fully faithful or an equivalence have been established for varieties over algebraically closed fields, generalizing results from the smooth case [HLS07, HLS09] and building upon foundational work of Bondal–Orlov [A.
+95].
 Likewise, Bondal–Orlov’s reconstruction theorem for smooth projective varieties has been extended to projective curves [SPE23] and (anti-)Fano Gorenstein varieties [BAL11].
 There is also growing interest in studying integral transforms in the relative setting.
-Specifically, let $Y_{1}$ and $Y_{2}$ be proper $S$-schemes where $S$ is Noetherian.
-An integral transform is an exact functor $\Phi_{K}:=\mathbf{R}(p_{2})_{\ast}(\mathbf{L}p_{1}^{\ast}(-)\otimes^{\mathbf{L}}K)\colon D_{\operatorname{qc}}(Y_{1})\to D_{\operatorname{qc}}(Y_{2})$ where $K\in D_{\operatorname{qc}}(Y_{1}\times_{S}Y_{2})$ and $p_{i}\colon Y_{1}\times_{S}Y_{2}\to Y_{i}$ are the projection morphisms.
-If $\Phi_{K}$ is an equivalence, then we say $Y_{1}$ and $Y_{2}$ are Fourier–Mukai partners (over $S$).
+Specifically, let $Y_{1}$ and $Y_{2}$ be proper $S$-schemes, where $S$ is Noetherian.
+An integral transform is an exact functor
+$\Phi_{K}:=\mathbf{R}(p_{2})_{\ast}(\mathbf{L}p_{1}^{\ast}(-)\otimes^{\mathbf{L}}K)\colon D_{\operatorname{qc}}(Y_{1})\to D_{\operatorname{qc}}(Y_{2})$
+where $K\in D_{\operatorname{qc}}(Y_{1}\times_{S}Y_{2})$ and $p_{i}\colon Y_{1}\times_{S}Y_{2}\to Y_{i}$ are the projection morphisms.
+If $\Phi_{K}$ is an equivalence, we say $Y_{1}$ and $Y_{2}$ are Fourier–Mukai partners (over $S$).
 Such transforms naturally arise in the study of fibrations; that is, in situations involving proper and flat morphisms.
@@
 cit.
-investigated when an integral transform between locally projective fibrations $Y_{i}\to S$ is an equivalence for $S$ of finite type over an algebraically closed field.
+studied when an integral transform between locally projective fibrations $Y_{i}\to S$ is an equivalence for $S$ of finite type over an algebraically closed field.
 It was shown that the induced integral transform on special fibers (i.e.
-closed points) over $S$ controls the behavior of equivalences and full faithfulness.
+closed points) over $S$ controls being an equivalence or fully faithful.
 1.2. What we do
-In this work, we deepen this perspective by systematically exploiting special fibers.
-Specifically, we study the behavior of integral transforms under base change in settings that incorporate local algebra.
-Moreover, our approach naturally extends several results in the literature to arbitrary fields, and at the same time yields new insight for singularities of fibrations.
+In this work, we deepen this perspective by systematically exploiting special fibers. Specifically, we study the behavior of integral transforms under base change using techniques from local algebra.
 Preservation of adjointness
@@
 [CNS25, NEE23]), which has resolved open conjectures in algebraic geometry [NEE21a, NEE24].
-As a consequence of this machinery, any integral transform $\Phi_{K}$ inducing an equivalence on $D_{\operatorname{qc}}$ restricts to equivalences on the subcategories of bounded complexes with coherent cohomology $D^{b}_{\operatorname{coh}}$ and perfect complexes $\operatorname{Perf}$.
-Moreover, an equivalence on these ‘smaller’ categories implies an equivalence on the ‘larger’ category.
-See Remark˜2.7 for details.
+As a consequence, any integral transform $\Phi_{K}$ inducing an equivalence on $D_{\operatorname{qc}}$ restricts to equivalences on $D^{b}_{\operatorname{coh}}$ and $\operatorname{Perf}$, see Lemma˜2.10.
 Consequently, questions about equivalences may be reduced to the behavior of integral transforms on these smaller subcategories.
-Whether an integral transform restricts to these smaller categories is now well-understood [Riz17, Bal09, DLM25].
+Whether an integral transform restricts to these smaller categories is now understood [RIZ17, BAL09, DLM25].
 These results shed light on the existence of left or right adjoints to integral transforms and the conditions under which such adjoints also restrict to these subcategories.
-However, the behavior of adjointness under base change is not yet fully understood.
-Our focus is on base changes induced by affine morphisms of finite tor-dimension.
-In particular, this leads us to the following, which summarizes Section˜3.
-Proposition 1.1.
-Suppose $t\colon T\to S$ is a morphism of Noetherian schemes.
-Let $Y_{1}$ and $Y_{2}$ be proper and flat $S$-schemes.
-Denote by $\pi\colon Y_{1}\times_{S}Y_{2}\times_{S}T\to Y_{1}\times_{S}Y_{2}$ the natural morphism.
-Consider $K\in D^{-}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$.
+However, the behavior of adjointness under base change is not yet fully understood. We focus is on base changes induced by certain morphisms. This leads us to the following.
+Theorem 1.1.
+Let $S$ be a Noetherian scheme. Suppose $Y_{1}$ and $Y_{2}$ are proper and flat $S$-schemes. Consider $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$. Then the following are equivalent:
   1. (1)
-    If $t$ is affine and $\Phi_{K}$ restricts to $D^{b}_{\operatorname{coh}}$ (resp.
-$\operatorname{Perf}$), then so does $\Phi_{\mathbf{L}\pi^{\ast}K}$ (see Lemma˜3.3).
-Additionally, the converse holds if $t$ is faithfully flat (see Corollary˜3.6).
+    $K$ is relatively perfect over $Y_{1}$ (resp. $Y_{2}$)
   2. (2)
-    Assume $t$ is affine and locally of finite tor-dimension.
-    1. (a)
-      If $\Phi_{K}$ restricts to $\operatorname{Perf}$, then $\mathbf{L}\pi^{\ast}$ preserves the adjunction of $\Phi_{K}$ with its right adjoint (see Proposition˜3.8).
-    2. (b)
-      If $\Phi_{K}$ restricts to $D^{b}_{\operatorname{coh}}$, then $\mathbf{L}\pi^{\ast}$ preserves the adjunction of $\Phi_{K}$ with its left adjoint (see Proposition˜3.12).
-In short, Proposition˜1.1 describes the behavior of integral transforms under base change when restricted to the small categories, specifically addressing whether the preservation of adjoints occurs.
-Key instances where it applies are morphisms $t$ arising from field extensions or finite surjective morphisms to a Dedekind scheme (i.e.
-smooth projective curves).
+    $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}T$ (resp. $Y_{2}\times_{S}T$) for any affine morphism $t$ from a Noetherian scheme
+  3. (3)
+    $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}\operatorname{Spec}(k)$ (resp.
+$Y_{2}\times_{S}\operatorname{Spec}(k)$) for all morphisms $t\colon\operatorname{Spec}(k)\to S$ from a field
+  4. (4)
+    $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}\operatorname{Spec}(\kappa(p))$ (resp.
+$Y_{2}\times_{S}\operatorname{Spec}(\kappa(p))$) for every closed point $p\in S$ with associated closed immersion $t\colon\operatorname{Spec}(\kappa(p))\to S$
+  5. (5)
+    $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}T$ (resp. $Y_{2}\times_{S}T$) for some affine surjective morphism $t\colon T\to S$.
+Here, $t^{\prime}\colon Y_{1}\times_{S}Y_{2}\times_{S}T\to Y_{1}\times_{S}Y_{2}$ is the natural morphism. In fact, in either case the corresponding left or right adjoints are preserved.
+This follows from Theorem˜3.6 together with Propositions˜3.8 and 3.11.
+In short, Theorem˜1.1 describes how integral transforms behave under base change when restricted to $\operatorname{Perf}$ or $D^{b}_{\operatorname{coh}}$.
+Importantly, it addresses the preservation of adjoints under base change.
+To the best of our knowledge, this result is new.
+In particular, it applies in the presence of singularities on the $Y_{i}$.
+An important instance where the result is applicable is $t$ arising from field extensions.
 Fully faithfulness & equivalences
-With this understanding of adjointness under base change in hand, we turn to questions of fully faithfulness and equivalence.
-Particularly, regarding the preservation of adjointness and the restriction to smaller categories.
-Next, we shift gears to studying fully faithfulness and equivalences under base change; specifically, the ascent and descent of these properties.
-Before doing so, we review the relevant literature a bit more.
-As mentioned earlier, the work of [HLS07, HLS09] studied these questions for fibrations whose base is of finite type over an algebraically closed field.
-A key ingredient in loc.
-cit.
-was to look at the behavior on special fibers.
+Next, we shift gears to studying fully faithfulness and equivalences under base change.
+Specifically, the ascent and descent of these properties.
+Before doing so, we briefly review the relevant literature.
+As mentioned earlier, [HLS07, HLS09] studied these questions for fibrations whose base is of finite type over an algebraically closed field.
+A key ingredient in their work was to look at the behavior on special fibers.
 More recently, such questions have also been studied for smooth proper morphisms over a Noetherian base scheme (see [KUR24, Proposition/Definition 2.3]).
@@
 cit.
-leverages ‘canonical morphisms’ between the kernels of integral transforms that describe the (co)units of adjunction.
+ises ‘canonical morphisms’ between the kernels of integral transforms that describe the (co)units of adjunction.
 However, it remains unclear whether such canonical morphisms exist in the singular setting.
 This uncertainty provides further motivation for our work: we bypass the absence of such tools in the singular setting by developing different techniques.
-This brings attention to the following, which summarizes Section˜4.
-Theorem 1.2 (see Corollary˜4.7).
-Consider affine morphism $t\colon T\to S$ locally of finite tor-dimension of Noetherian schemes whose target is semiseparated.
+This brings attention to the following.
+Theorem 1.2.
+Let $S$ be a Noetherian scheme.
 Suppose $Y_{1}$ and $Y_{2}$ are proper and flat $S$-schemes.
-Denote by $\pi\colon Y_{1}\times_{S}Y_{2}\times_{S}T\to Y_{1}\times_{S}Y_{2}$ the natural morphism.
-Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over each $Y_{i}$.
-If $\Phi_{K}$ restricts to a fully faithful functor (resp.
-an equivalence) on $D^{b}_{\operatorname{coh}}$, then so does $\Phi_{\mathbf{L}\pi^{\ast}K}$.
-In practice, many schemes we may consider are separated, and hence, semiseparated. As a pleasant consequence, this allows us to partially generalize Orlov’s result to the singular setting.
-Corollary 1.3 (cf. [Orl02, Lemma 2.12]).
+Consider $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ which is relatively perfect over each $Y_{i}$.
+Then the following are equivalent:
+  1. (1)
+    $\Phi_{K}$ is fully faithful (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$
+  2. (2)
+    $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is fully faithful (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$ for any affine morphism $t$
+  3. (3)
+    $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is fully faithful (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$ for every morphism $t\colon\operatorname{Spec}(k)\to S$ from a field
+  4. (4)
+    $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is fully faithful (resp.
+an equivalence) on $D^{b}_{\operatorname{coh}}$ for every closed point $p\in S$ with associated closed immersion $t\colon\operatorname{Spec}(\kappa(p))\to S$
+  5. (5)
+    $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is fully faithful (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$ for some affine surjection $t$.
+Here, $t^{\prime}\colon Y_{1}\times_{S}Y_{2}\times_{S}T\to Y_{1}\times_{S}Y_{2}$ is the natural morphism induced by $t\colon T\to S$.
+This appears later as Theorem˜4.13.
+Notably, it generalizes [HLS09, Proposition 2.15], which only proved the statement for closed points.
+As a consequence, this allows us to generalize Orlov’s result to the singular setting (cf.
+[ORL02, Lemma 2.12]).
+Corollary 1.3.
 Consider proper schemes $Y_{1}$ and $Y_{2}$ over a field $k$.
 Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{k}Y_{2})$ be relatively perfect over each $Y_{i}$.
-If $\Phi_{K}$ restricts to a fully faithful functor (resp.
-an equivalence) on $D^{b}_{\operatorname{coh}}$, then so does $\Phi_{K_{L}}$ for any field extension $L/k$.
-Our proofs require careful bookkeeping of how the (co)units of adjunctions for integral transforms behave under base change.
-However, additional subtleties also arise in this context.
-In particular, we need to work in the larger category $D_{\operatorname{qc}}$ in order to have a firm grasp on ascending fully faithfulness; see e.g.
-Proposition˜4.4 and Theorem˜4.6.
+Then $\Phi_{K_{L}}$ is fully faithful (resp.
+an equivalence) on $D^{b}_{\operatorname{coh}}$ for any field extension $L/k$ if, and only if, $\Phi_{K}$ is fully faithful (resp.
+an equivalence) on $D^{b}_{\operatorname{coh}}$.
+Our proofs require careful bookkeeping of how the (co)units of adjunction for integral transforms behave under base change.
+See Section˜4 for details.
 Notably, the techniques we use are independent of those appearing in [ORL02, Lemma 2.12].
-1.3. Why it matters
-To conclude our work, we discuss the ramifications of the results above.
+1.3. What is next
+We conclude by discussing several consequences of the results above.
 Our results extend the literature on schemes of finite type over algebraically closed fields to the more general setting of arbitrary fields.
 They reveal a new interplay between the intuition of [HLS07, HLS09] concerning special fibers and faithfully flat descent along affine morphisms.
-In particular, certain singularities are shown to be derived invariants for fibrations over Noetherian schemes, broadening the scope to include arithmetically flavored situations of potential interest for future research.
-For a pair of geometrically integral projective curves over a field $k$ of characteristic zero, Fourier–Mukai partnership implies they are $k$-forms (see Proposition˜5.1).
+In particular, certain singularities are shown to be derived invariants for fibrations over Noetherian schemes.
+This broadens the scope to include arithmetically flavored situations of potential interest for future research.
+For a pair of geometrically integral projective curves over a field $k$ of characteristic zero, we show Fourier–Mukai partnership implies they are $k$-forms. Particularly:
+Proposition 1.4.
+Let $k$ be a field of characteristic zero.
+Suppose $C$ and $C^{\prime}$ are geometrically integral projective curves over $k$.
+If $C$ and $C^{\prime}$ are Fourier–Mukai partners over $k$, then $\operatorname{Spec}(L)\times_{k}C\cong\operatorname{Spec}(L)\times_{k}C^{\prime}$ for some finite field extension $L/k$ (i.e.
+are $k$-forms).
 Here, a $k$-form means there exists a finite extension $L/k$ such that their base changes along $L/k$ become isomorphic.
-The result holds for singular curves and enriches algebraically closed case of [Spe23].
-For fibrations of relative dimension one over a regular Noetherian $\mathbb{Q}$-scheme, Fourier–Mukai partnership forces the genera of special fibers to coincide (see Corollary˜5.2).
-Smoothness for proper schemes over an arbitrary field is likewise invariant under Fourier–Mukai partnership (see Proposition˜5.3).
+The result holds for singular curves and extends the algebraically closed case of [SPE23].
+For fibrations of relative dimension one over a Noetherian $\mathbb{Q}$-scheme, Fourier–Mukai partnership forces the genera of special fibers to coincide (see Corollary˜5.1).
+It would be interesting to find a prime or mixed characteristic analog of these results.
+Next, we show that smoothness for proper schemes over an arbitrary field is likewise invariant under Fourier–Mukai partnership (see Lemma˜5.2).
 This subtlety arises especially over imperfect fields.
@@
 [KOL11, Remark 16]).
-In the setting of projective fibrations, smoothness of the total space is also a derived invariant if the base is regular (see Corollary˜5.4).
-Finally, generalizing [HLS09], we show that the Cohen–Macaulay and Gorenstein properties of a morphism remain invariant under Fourier–Mukai partnership for regular Noetherian base schemes (see Proposition˜5.5).
-In a related but distinct direction, we study a variation of Proposition˜5.5 in the case of derived equivalences.
-Particularly, we prove a version for the Gorenstein condition for arbitrary derived equivalences where the required functor need not be an integral transform.
-See Proposition˜5.6.
+In the setting of fibrations, smoothness of the total space is also a derived invariant:
+Proposition 1.5.
+Let $f_{i}\colon Y_{i}\to S$ be proper flat morphisms to a Noetherian scheme.
+Assume that $Y_{1}$ and $Y_{2}$ are Fourier–Mukai partners over $S$.
+Then $f_{1}$ is smooth if, and only if, $f_{2}$ is smooth.
+Finally, generalizing [HLS09], we show that the Cohen–Macaulay and Gorenstein properties of a morphism remain invariant under Fourier–Mukai partnership for regular Noetherian base schemes.
+Specifically:
+Proposition 1.6.
+Let $f_{i}\colon Y_{i}\to S$ be projective flat morphisms with geometrically integral fibers to a Noetherian scheme.
+Assume that $Y_{1}$ and $Y_{2}$ are Fourier–Mukai partners over $S$.
+Then $f_{1}$ is Cohen–Macaulay (resp.
+Gorenstein) if, and only if, $f_{2}$ satisfies the same condition.
+Under additional assumptions, Proposition˜1.6 can be strengthened in the Gorenstein case. See Proposition˜5.3 for details. Its proof relies on techniques that differ from those in Proposition˜1.6.
+Finally, we comment on potential generalizations of our work.
+The authors are currently investigating stack theoretic consequences of these results.
+Under mild assumptions, the work of [AJS23] extends to suitable algebraic stacks.
+At present, it appears that much of our work carries over to this setting with a bit of effort.
+We also plan to explore deeper applications to Fourier–Mukai theory for moduli spaces using the tools developed here.
 Notation
-Let $X$ be a Noetherian scheme. We work with the following triangulated categories:
+Let $X$ be a scheme. We work with the following triangulated categories:
   1. (1)
@@
   3. (3)
-    $D_{\operatorname{coh}}^{b}(X)$ (resp.
+    If $X$ is Noetherian, then $D_{\operatorname{coh}}^{b}(X)$ (resp.
 $D_{\operatorname{qc}}^{b}(X)$) is the (strictly full) subcategory of $D(X)$ consisting of complexes having bounded and coherent (resp.
@@
     $\operatorname{Perf}(X)$ is the (strictly full) subcategory of $D_{\operatorname{qc}}(X)$ consisting of the perfect complexes on $X$.
-At times if $X$ is affine, we abuse notation and write $D_{\operatorname{qc}}(R):=D_{\operatorname{qc}}(X)$ where $R:=H^{0}(X,\mathcal{O}_{X})$ are the global sections; similar conventions will occur for the other categories mentioned here and later.
+At times if $X$ is affine, we might abuse notation and write $D^{b}_{\operatorname{coh}}(R):=D^{b}_{\operatorname{coh}}(X)$ where $R:=H^{0}(X,\mathcal{O}_{X})$ are the global sections; similar conventions will occur for the other categories mentioned here and later.
 Acknowledgments.
@@
 Part of this work was completed while Lank visited the Basque Center for Applied Mathematics (BCAM) and thanks them for their hospitality.
-The authors thank Dylan Spence, Ana Cristina López-Martín, and Elías Guisado Villalgordo for discussions in an earlier version of the work.
+Guisado Villalgordo was supported under the grant PRE2022-104796 by the Spanish Ministry of Science and Innovation.
+Manali Rahul is grateful to Max Planck Institute for Mathematics in Bonn for its hospitality and financial support.
+The authors thank Dylan Spence, Ana Cristina López-Martín, and Fernando Sancho de Salas for discussions.
 2. Preliminaries
@@
 2.1. Perfect & pseudocoherent complexes
-We say $P\in D_{\operatorname{qc}}(X)$ is perfect if it is affine locally isomorphic to a bounded complex of locally free sheaves of finite rank.
+We say $P\in D_{\operatorname{qc}}(X)$ is perfect if it is locally isomorphic to a bounded complex of locally free sheaves of finite rank.
 Denote by $\operatorname{Perf}(X)$ for the full subcategory of perfect complexes.
@@
 We say $E\in D_{\operatorname{qc}}(X)$ is pseudocoherent if it is $N$-pseudocoherent for every $N\in\mathbb{Z}$.
-In general, since $X$ is a scheme, $E\in D_{\operatorname{qc}}(X)$ is pseudocoherent if and only if locally on $X$ there is a quasi-isomorphism $P\to E$, with $P$ a bounded above complex of finite-rank free modules [GW23, Corollary 22.47].
+In general, $E\in D_{\operatorname{qc}}(X)$ is pseudocoherent if, and only if, locally on $X$ there is a quasi-isomorphism $P\to E$, with $P$ a bounded above complex of finite-rank free modules [GW23, Corollary 22.47].
 As $X$ is Noetherian, $E\in D_{\operatorname{qc}}(X)$ is pseudocoherent if, and only if, $\mathcal{H}^{n}(E)$ is coherent for all $n$ and vanishes for $n\gg 0$ (see e.g.
@@
 [STA25, Tag 08H4]).
-We will use this characterization liberally.
-While straightforward, the following is useful.
+We record a few straightforward results.
 Lemma 2.1.
-Let $f\colon Y\to X$ be an affine faithfully flat morphism of Noetherian schemes.
+Let $f\colon Y\to X$ be a faithfully flat morphism of Noetherian schemes.
 An object $E\in D_{\operatorname{qc}}(X)$ belongs to $D^{-}_{\operatorname{coh}}(X)$ (resp.
@@
 Proof.
-This is a special case of [GW23, Proposition 22.52], but we spell it out for convenience.
-The problem is affine local.
-So, we can reduce to the affine case.
-Then the claim follows from [Sta25, Tag 068R] (resp.
-[GW23, Lemma 21.172]).
+This is a special case of [GW23, Proposition 22.52]. ∎
+Lemma 2.2.
+Let $f\colon W\to Y$ be a morphism of schemes. Choose $E,G\in D(Y)$. Assume one of the following holds:
+  1. (1)
+    $E$ is perfect
+  2. (2)
+    $E$ is pseudocoherent and both $G,\mathbf{L}f^{\ast}G$ are locally bounded below.
+Then the natural morphism
+$\mathbf{L}f^{\ast}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(E,G)\to\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}f^{\ast}E,\mathbf{L}f^{\ast}G)$
+is an isomorphism.
+Proof.
+On one hand, (1) is a special case of [GW23, Proposition 22.70].
+On the other hand, (2) is argued exactly as in the case of loc.
+cit.
+where ‘$f$ has finite tor-dimension’ can be replaced with our hypothesis that $\mathbf{L}f^{\ast}G$ is locally bounded below.
+Indeed, this is the only part in the proof of loc.
+cit.
+which requires the assumption of finite tor-dimension.
 ∎
+Next, we give a characterization of perfectness.
+To do so, we need notation.
+Given a scheme $X$ and a point $p\in X$, denote by $\sigma_{p}\colon\operatorname{Spec}(\mathcal{O}_{X,p})\to X$ and $i_{p}\colon\operatorname{Spec}(\kappa(p))\to X$ for the natural morphisms.
+Note that the set theoretic image of $\sigma_{p}$ is the intersection of all open subschemes $U$ of $X$ which contain $p$.
+Lemma 2.3.
+Let $X$ be a scheme, $E\in D_{\operatorname{qc}}^{b}(X)$ be pseudocoherent, and $p\in X$. Then the following are equivalent:
+  1. (1)
+    There is an open immersion $j\colon U\to X$ such that $p\in U$ and $\mathbf{L}j^{\ast}E$ is perfect
+  2. (2)
+    $\mathbf{L}\sigma_{p}^{\ast}E$ is perfect
+  3. (3)
+    $\mathbf{L}i_{p}^{\ast}E$ is bounded.
+Proof.
+[…]
+$\operatorname{Spec}(\kappa(p))\to\operatorname{Spec}(\mathcal{O}_{X,p})\to U\xrightarrow{j}X.$
+Indeed, one uses the fact that derived pullback preserves perfect complexes (see e.g. [STA25, Tag 09UA]).
+[…]
+The following is an extension of [AJS23, Theorem 2.3] to closed points.
+Proposition 2.4.
+Let $X$ be a quasi-separated quasi-compact scheme and $E\in D^{b}_{\operatorname{qc}}(X)$ be pseudocoherent. Then the following conditions are equivalent:
+  1. (1)
+    $E$ is perfect
+  2. (2)
+    $\mathbf{L}i_{p}^{\ast}E$ is bounded for every $p\in X$
+  3. (3)
+    $\mathbf{L}i_{p}^{\ast}E$ is bounded for every $p\in X$ which is closed
+  4. (4)
+    $\mathbf{L}\sigma_{p}^{\ast}E$ is perfect for every $p\in X$
+  5. (5)
+    $\mathbf{L}\sigma_{p}^{\ast}E$ is perfect for every $p\in X$ which is closed.
+Proof.
+[…]
 2.2. Integral transforms
@@
 When clear from context, the hyphen ‘$S$-’ is dropped.
-Now, we impose each $f_{i}$ be proper and flat. An object $K\in D^{-}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ is called relatively perfect over $Y_{1}$ if
+Now, suppose each $f_{i}$ be proper and flat. An object $K\in D^{-}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ is called relatively perfect over $Y_{1}$ if
 $\mathbf{R}(p_{1})_{\ast}(K\otimes^{\mathbf{L}}\operatorname{Perf}(Y_{1}\times_{S}Y_{2}))\subseteq\operatorname{Perf}(Y_{1}).$
@@
 So, in our setting, $K\in D^{-}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ is relatively perfect over $Y_{1}$ precisely when it is $p_{1}$-perfect.
-This notion has appeared in the literature, but we record a lemma for convenience to avoid any potential confusion.
+This notion has appeared in the literature but we record a lemma for convenience to avoid any potential confusion.
 Lemma 2.5.
@@
   1. (1)
-    $E$ is $f$-perfect.
+    $E$ is $f$-perfect
   2. (2)
-    $E\otimes^{\mathbf{L}}\mathbf{L}f^{\ast}D^{b}_{\operatorname{coh}}(X)\subseteq D^{b}_{\operatorname{coh}}(Y)$.
+    $E\otimes^{\mathbf{L}}\mathbf{L}f^{\ast}D^{b}_{\operatorname{coh}}(X)\subseteq D^{b}_{\operatorname{coh}}(Y)$
   3. (3)
@@
 Remark 2.6.
-A useful observation from Lemma˜2.2 is that any object $E\in D^{-}_{\operatorname{coh}}(Y)$ being $f$-perfect implies it belongs to $D^{b}_{\operatorname{coh}}(Y)$.
+A useful observation from Lemma˜2.5 is that any object $E\in D^{-}_{\operatorname{coh}}(Y)$ being $f$-perfect implies it belongs to $D^{b}_{\operatorname{coh}}(Y)$.
 2.3. Generation for triangulated categories
@@
 Fix a subcategory $\mathcal{S}\subseteq\mathcal{T}$.
-Denote by $\operatorname{add}(\mathcal{S})$ the strictly full (that is, closed under isomorphisms) subcategory of $\mathcal{T}$ consisting of all direct summands of objects of the form $\oplus_{n\in\mathbb{Z}}S_{n}^{\oplus r_{n}}[n]$ where $S_{n}\in\mathcal{S}$ for all $n$ and finitely many $r_{n}$ are nonzero.
+Denote by $\operatorname{add}(\mathcal{S})$ for the strictly full (that is, closed under isomorphisms) subcategory of $\mathcal{T}$ consisting of all direct summands of objects of the form $\oplus_{n\in\mathbb{Z}}S_{n}^{\oplus r_{n}}[n]$ where $S_{n}\in\mathcal{S}$ for all $n$ and finitely many $r_{n}$ are nonzero.
 Inductively, let
-$\langle\mathcal{S}\rangle_{n}:=\begin{cases}\operatorname{add}(\varnothing)&\text{if }n=0,\\ \operatorname{add}(\mathcal{S})&\text{if }n=1,\\ \operatorname{add}(\{\operatorname{cone}\phi\mid\phi\in\operatorname{Hom}(\langle\mathcal{S}\rangle_{n-1},\langle\mathcal{S}\rangle_{1})\})&\text{if }n>1.\end{cases}$
-Additionally, let $\langle\mathcal{S}\rangle:=\langle\mathcal{S}\rangle_{\infty}:=\cup_{n\geq 0}\langle\mathcal{S}\rangle_{n}$.
+$\langle\mathcal{S}\rangle_{n}:=\begin{cases}\operatorname{add}(\varnothing)&\text{if }n=0\\ \operatorname{add}(\mathcal{S})&\text{if }n=1\\ \operatorname{add}(\{\operatorname{cone}\phi\mid\phi\in\operatorname{Hom}(\langle\mathcal{S}\rangle_{n-1},\langle\mathcal{S}\rangle_{1})\})&\text{if }n>1.\end{cases}$
+Additionally, let $\langle\mathcal{S}\rangle:=\cup_{n\geq 0}\langle\mathcal{S}\rangle_{n}$.
 It may be checked that $\langle\mathcal{S}\rangle$ is the smallest thick (that is, triangulated subcategory closed under direct summands) subcategory containing $\mathcal{S}$.
@@
 An object $G\in\mathcal{T}$ is called a classical generator for $\mathcal{T}$ if $\langle G\rangle=\mathcal{T}$.
-Now, we impose $\mathcal{T}$ be a compactly generated triangulated category.
-To clarify, we require that $\mathcal{T}$ admits all small coproducts.
+Suppose that $\mathcal{T}$ is compactly generated triangulated category (to clarify, we require that $\mathcal{T}$ admits all small coproducts in the definition of being compactly generated).
 Denote the collection of compact objects of $\mathcal{T}$ by $\mathcal{T}^{c}$.
 We say $G\in\mathcal{T}^{c}$ is a compact generator for $\mathcal{T}$ if for any $E\in\mathcal{T}$, one has $\operatorname{Hom}(G,E[n])=0$ for all $n\in\mathbb{Z}$ if, and only if, $E=0$.
-A well-known fact is that classical generators for $\mathcal{T}^{c}$ coincide with being compact generators for $\mathcal{T}$ (see e.g.
+A well-known fact is that classical generators for $\mathcal{T}^{c}$ coincide with compact generators for $\mathcal{T}$ (see e.g.
 [STA25, Tag 09SR]).
-The following is an important case and is freely used without mention.
-Example 2.4 ([BvdB03, Theorem 3.1.1]).
+The following is an important case.
+Example 2.7 ([Bv03, Theorem 3.1.1]).
 Let $X$ be a quasi-compact quasi-separated scheme.
-The collection of compact objects in $D_{\operatorname{qc}}(X)$ coincides with $\operatorname{Perf}(X)$.
-Moreover, $D_{\operatorname{qc}}(X)$ is compactly generated by a single object from $\operatorname{Perf}(X)$.
-See [BvdB03, Theorem 3.1.1] for details.
-The following is known for some circles, but we add it for self-containedness (see e.g. [KL15, Lemma’s 2.10 & 2.12]).
+The subcategory of compact objects in $D_{\operatorname{qc}}(X)$ coincides with $\operatorname{Perf}(X)$.
+Moreover, $D_{\operatorname{qc}}(X)$ is compactly generated by a single object from $\operatorname{Perf}(X)$, see [Bv03, Theorem 3.1.1] for details.
+We add the following for self-containedness (see e.g. [KL15, Lemma’s 2.10 & 2.12]).
 Lemma 2.8.
@@
 ∎
-Terminology 2.6.
+Terminology 2.9.
 Let $\Phi\colon D_{\operatorname{qc}}(Y_{1})\to D_{\operatorname{qc}}(Y_{2})$ be an exact functor where each $Y_{i}$ is Noetherian.
@@
 The same goes for properties such as being fully faithful or an equivalence.
-Remark 2.7.
-Let $\Phi\colon D_{\operatorname{qc}}(Y_{1})\to D_{\operatorname{qc}}(Y_{2})$ be an exact functor where each $Y_{i}$ is Noetherian. Then the following are equivalent:
+Lemma 2.10.
+Let $Y_{1}$ and $Y_{2}$ be Noetherian schemes.
+Suppose $\Phi\colon D_{\operatorname{qc}}(Y_{1})\to D_{\operatorname{qc}}(Y_{2})$ is an exact functor which preserves small coproducts.
+Then the following are equivalent:
   1. (1)
@@
     $\Phi$ restricts to an equivalence on $\operatorname{Perf}$.
+Proof.
 […]
@@
 Lemma 3.2.
-Consider ˜3.1 with $t$ a morphism locally of finite tor-dimension.
-Then $t_{1},t_{2},t^{\prime}$ are locally of finite tor-dimension.
-In particular, each face of Diagram˜3.1 is a tor-independent square (in the sense of [Lip09, Definition 3.10.2]).
+Consider ˜3.1. Then each face of Diagram˜3.1 is a tor-independent square (in the sense of [LIP09, Definition 3.10.2]).
 Proof.
-This first claim is affine local on $S$.
-Indeed, if needed, we can find an affine open cover of $S$ such that the base changes of $t$ along the associated open immersion are of finite tor-dimension.
-Then, for each such open immersion of a component in the affine cover, we have a similar diagram to Diagram˜3.1 (i.e.
-a ‘base change cube’).
-If we can show the desired properties on the morphisms arising in each diagram associated to the components of the affine open cover of $S$, then the desired claim follows.
-So, we may impose that $t$ be of finite tor-dimension.
-By base change, we know that $g_{1},g_{2},f^{\prime}_{2},f^{\prime}_{1},g_{2}^{\prime},g_{1}^{\prime}$ are proper flat morphisms.
-Consider the fibered square in Diagram˜3.1 induced from the base change of $f_{1}$ along $t$.
-It follows by [GW23, Lemma 23.27] that $t_{1}$ is of finite tor-dimension.
-A similar argument shows $t_{2}$ is of finite tor-dimension as well.
-In fact, if we work with the fibered square in Diagram˜3.1 obtained by base change of $f_{2}^{\prime}$ along $t_{1}$, we see from similar reasoning that $t^{\prime}$ is locally of finite tor-dimension.
-This completes the proof for our first claim.
-To check the second claim, it can be checked by one face at a time.
-Note that the same argument in what follows works whether $t$ has finite tor-dimension locally or absolutely because the conditions checked are at the level of stalks.
-The fibered square obtained by base change of $f_{1}$ along $t$ is tor-independent.
-Indeed, use [GW23, Exercise 21.32(iii)] with the constraint that $f_{1}$ is flat.
-Then [GW23, Exercise 21.33(v)], coupled with [Lip09, Definition 3.10.2(i) & Theorem 3.10.3], tells us the particular face is tor-independent.
-Now, we can proceed in a similar fashion with the remaining faces to finish the proof.
-∎
+As each $f_{i}$, $f^{\prime}_{i}$, $g_{i}$, and $g^{\prime}_{i}$ are flat, the desired claim follows from [GW23, Remark 22.94]. ∎
 Lemma 3.3.
 Consider ˜3.1 with $t$ an affine morphism.
-Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{1}$.
-Then $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ restricts to an exact functor on $D^{b}_{\operatorname{coh}}$.
-Similarly, if $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ is relatively perfect over $Y_{2}$, then $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ restricts to an exact functor on $\operatorname{Perf}$.
+If $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ is relatively perfect over $Y_{1}$ (resp.
+over $Y_{2}$), then $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ restricts to an exact functor on $D^{b}_{\operatorname{coh}}$ (resp.
+on $\operatorname{Perf}$).
 Proof.
-We only prove the case of $K$ being relatively perfect over $Y_{2}$ implying $\Phi_{\mathbf{L}(t^{\prime})K}$ restricts to an exact functor on $\operatorname{Perf}$ because a similar argument applies for $D^{b}_{\operatorname{coh}}$.
+We only prove the case of $K$ being relatively perfect over $Y_{2}$ because a similar argument applies for $D^{b}_{\operatorname{coh}}$.
 As $\Phi_{K}(\operatorname{Perf}(Y_{1}))\subseteq\operatorname{Perf}(Y_{2})$, [DLM25, Proposition 3.2] implies
@@
 Lemma 3.4.
-Consider ˜3.1.
-Let $K\in D_{\operatorname{qc}}(Y_{1}\times_{S}Y_{2})$.
-Then there is a natural isomorphism $\beta^{K}\colon\mathbf{L}t_{2}^{\ast}\circ\Phi_{K}\to\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\circ\mathbf{L}t_{1}^{\ast}$ on $D_{\operatorname{qc}}$.
+Consider ˜3.1. Let $K\in D_{\operatorname{qc}}(Y_{1}\times_{S}Y_{2})$. Then on $D_{\operatorname{qc}}$ there is a natural isomorphism
+$\beta^{K}\colon\mathbf{L}t_{2}^{\ast}\circ\Phi_{K}\to\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\circ\mathbf{L}t_{1}^{\ast}.$
 Proof.
-This follows from the string of functorial isomorphisms for each $E\in D_{\operatorname{qc}}(Y_{1}\times_{S}T)$:
-$\mathbf{L}t_{2}^{\ast}\Phi_{K}(E)$ $\cong\mathbf{L}t_{2}^{\ast}\mathbf{R}(f^{\prime}_{1})_{\ast}(\mathbf{L}(f^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}K)$
-$\cong\mathbf{R}(g^{\prime}_{1})_{\ast}\mathbf{L}(t^{\prime})^{\ast}(\mathbf{L}(f^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}K)$
-$\cong\mathbf{R}(g^{\prime}_{1})_{\ast}(\mathbf{L}(t^{\prime})^{\ast}\mathbf{L}(f^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K)$
-$\cong\mathbf{R}(g^{\prime}_{1})_{\ast}(\mathbf{L}(g^{\prime}_{2})^{\ast}\mathbf{L}(t_{1}^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K))$
+This follows from the string of isomorphisms for each $E\in D_{\operatorname{qc}}(Y_{1}\times_{S}T)$:
+$\mathbf{L}t_{2}^{\ast}\Phi_{K}(E)$ $=\mathbf{L}t_{2}^{\ast}\mathbf{R}(f^{\prime}_{1})_{\ast}(\mathbf{L}(f^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}K)$
+$\cong\mathbf{R}(g^{\prime}_{1})_{\ast}\mathbf{L}(t^{\prime})^{\ast}(\mathbf{L}(f^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}K)$ (flat base change)
+$\cong\mathbf{R}(g^{\prime}_{1})_{\ast}(\mathbf{L}(t^{\prime})^{\ast}\mathbf{L}(f^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K)$ (monoidality of $\mathbf{L}(-)^{\ast}$)
+$\cong\mathbf{R}(g^{\prime}_{1})_{\ast}(\mathbf{L}(g^{\prime}_{2})^{\ast}\mathbf{L}(t_{1})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K)$ (pseudofunctoriality of $\mathbf{L}(-)^{\ast}$)
 $=\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}(\mathbf{L}t_{1}^{\ast}E).$
-Proposition 3.5.
-Consider ˜3.1 where $t$ is an affine faithfully flat morphism.
-Let $K\in D^{-}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$.
+Lemma 3.5.
+Consider ˜3.1. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$.
+  1. (1)
     If $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}T$ (resp.
-$Y_{2}\times_{S}T$), then $K$ is relatively perfect over $Y_{1}$ (resp.
+$Y_{2}\times_{S}T$) where $t$ is affine and faithfully flat, then $K$ is relatively perfect over $Y_{1}$ (resp.
+$Y_{2}$).
+  2. (2)
+    If $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}\operatorname{Spec}(k)$ (resp.
+$Y_{2}\times_{S}\operatorname{Spec}(k)$) for all morphisms $\operatorname{Spec}(k)\to S$ from a field, then $K$ is relatively perfect over $Y_{1}$ (resp.
+$Y_{2}$).
+  3. (3)
+    If $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}\operatorname{Spec}(\kappa(p))$ (resp.
+$Y_{2}\times_{S}\operatorname{Spec}(\kappa(p))$) for all closed immersions $\operatorname{Spec}(\kappa(p))\to S$ associated to a closed point $p\in S$, then $K$ is relatively perfect over $Y_{1}$ (resp.
 $Y_{2}$).
 Proof.
-First, we show the case for $\mathbf{L}(t^{\prime})^{\ast}K$ being relatively perfect over $Y_{1}\times_{S}T$ implying $K$ is relatively perfect over $Y_{1}$.
-Choose $P_{i}\in\operatorname{Perf}(Y_{i})$ such that $\langle P_{i}\rangle=\operatorname{Perf}(Y_{i})$ for each $i$.
-From [BvdB03, Lemma 3.4.1], we know that $\langle\mathbf{L}f_{2}^{\ast}P_{1}\otimes^{\mathbf{L}}\mathbf{L}f_{1}^{\ast}P_{2}\rangle=\operatorname{Perf}(Y_{1}\times_{S}Y_{2})$.
-Then, by [Sta25, Tag 0BQT], it follows that
-$\langle\mathbf{L}(t^{\prime})^{\ast}(\mathbf{L}f_{2}^{\ast}P_{1}\otimes^{\mathbf{L}}\mathbf{L}f_{1}^{\ast}P_{2})\rangle=\operatorname{Perf}(Y_{1}\times_{S}Y_{2}\times_{S}T).$
-However, we have an isomorphism in Diagram˜3.1,
-$\mathbf{L}(t^{\prime})^{\ast}(\mathbf{L}f_{2}^{\ast}P_{1}\otimes^{\mathbf{L}}\mathbf{L}f_{1}^{\ast}P_{2})$ $\cong\mathbf{L}(t^{\prime})^{\ast}\mathbf{L}f_{2}^{\ast}P_{1}\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}\mathbf{L}f_{1}^{\ast}P_{2}$
-$\cong\mathbf{L}(g_{2}^{\prime})^{\ast}\mathbf{L}t_{1}^{\ast}P_{1}\otimes^{\mathbf{L}}\mathbf{L}(g_{1}^{\prime})^{\ast}\mathbf{L}t_{2}^{\ast}P_{2}.$
-Now, we finish the proof for the claim.
+We only show the case for $\mathbf{L}(t^{\prime})^{\ast}K$ being relatively perfect over $Y_{1}\times_{S}T$. To show the other case, one can argue like below and use Lemma˜2.1.
+To start, we prove the first claim.
 Let $E\in D^{b}_{\operatorname{coh}}(Y_{1})$.
@@
 So, it suffices to show that $\mathcal{H}^{j}(\Phi_{K}(E))=0$ for all but finitely many $j\in\mathbb{Z}$.
-Our hypothesis ensures that $\Phi_{\mathbf{L}(t^{\prime})K}(\mathbf{L}t_{1}^{\ast}E)\in D^{b}_{\operatorname{coh}}(Y_{2}\times_{S}T)$. Then, as $t_{2}$ is flat, for each $j\in\mathbb{Z}$,
+Our hypothesis ensures that $\Phi_{\mathbf{L}(t^{\prime})K}(\mathbf{L}t_{1}^{\ast}E)\in D^{b}_{\operatorname{coh}}(Y_{2}\times_{S}T)$.
+Then, as $t_{2}$ is flat, for each $j\in\mathbb{Z}$,
 $\operatorname{supp}(\mathcal{H}^{j}(\Phi_{K}(E)))$ $=t_{2}(\operatorname{supp}(t_{2}^{\ast}\mathcal{H}^{j}(\Phi_{K}(E))))$ ([STA25, Tag 056J])
@@
 Consequently, we have $\Phi_{K}(D^{b}_{\operatorname{coh}}(Y_{1}))\subseteq D^{b}_{\operatorname{coh}}(Y_{2})$, which implies $K$ is relatively perfect over $Y_{1}$.
-To show the case of perfect complexes being preserved by integral transforms, we can argue as above and use Lemma˜2.1.
+Next, we check the second claim. Choose any $p\in Y_{1}$. Denote by $t\colon\operatorname{Spec}(\kappa(f_{1}(p)))\to S$ for the natural morphism. There is a commutative diagram
+[…]
+where $h_{1}$ is the natural morphism. Using the hypothesis, it follows that
+$\mathbf{R}(g^{\prime}_{2})_{\ast}$ $(\mathbf{L}(t^{\prime})^{\ast}\big(K\otimes^{\mathbf{L}}\operatorname{Perf}(Y_{1}\times_{S}Y_{2})\big))$
+$\subseteq\mathbf{R}(g^{\prime}_{2})_{\ast}(\mathbf{L}(t^{\prime})^{\ast}K\otimes^{\mathbf{L}}\operatorname{Perf}(Y_{1}\times_{S}Y_{2}\times_{S}\operatorname{Spec}(\kappa(f_{1}(p)))))$
+$\subseteq\operatorname{Perf}(Y_{1}\times_{S}\operatorname{Spec}(\kappa(f_{1}(p)))).$
+There is a fibered square
+$Y_{1}\times_{S}Y_{2}\times_{S}\operatorname{Spec}(\kappa(f_{1}(p)))$$Y_{1}\times_{S}\operatorname{Spec}(\kappa(f_{1}(p)))$$Y_{1}\times_{S}Y_{2}$$Y_{1}.$$\scriptstyle{g_{2}^{\prime}}$$\scriptstyle{t^{\prime}}$$\scriptstyle{t_{1}}$$\scriptstyle{f^{\prime}_{2}}$
+So, from flat base change, we have for all $P\in\operatorname{Perf}(Y_{1}\times_{S}Y_{2})$,
+$\mathbf{L}t_{1}^{\ast}\mathbf{R}(f^{\prime}_{2})_{\ast}(K\otimes^{\mathbf{L}}P)$ $\cong\mathbf{R}(g^{\prime}_{2})_{\ast}\mathbf{L}(t^{\prime})^{\ast}(K\otimes^{\mathbf{L}}P)$
+$\in\operatorname{Perf}(Y_{1}\times_{S}\operatorname{Spec}(\kappa(f_{1}(p)))).$
+This implies that
+$\mathbf{L}h_{1}^{\ast}\mathbf{R}(f^{\prime}_{2})_{\ast}(K\otimes^{\mathbf{L}}P)\cong\mathbf{L}h^{\ast}\mathbf{L}t_{1}^{\ast}\mathbf{R}(f^{\prime}_{2})_{\ast}(K\otimes^{\mathbf{L}}P)\in D^{b}_{\operatorname{coh}}(\kappa(p)),$
+because the derived pullback of perfect complexes remain perfect (see e.g.
+[STA25, Tag 09UA]).
+Using [AJS23, Theorem 2.3(iv)], we deduce that $\mathbf{R}(f^{\prime}_{2})_{\ast}(K\otimes^{\mathbf{L}}P)\in\operatorname{Perf}(Y_{1})$; note that $\mathbf{R}(f^{\prime}_{2})_{\ast}(K\otimes^{\mathbf{L}}P)$ is bounded because $K\otimes^{\mathbf{L}}P$ is a bounded pseudocoherent complex and $f^{\prime}_{2}$ is proper.
+Indeed, we have shown that $\mathbf{L}b^{\ast}\mathbf{R}(f^{\prime}_{2})_{\ast}(K\otimes^{\mathbf{L}}P)$ is bounded for all $q\in Y_{1}$ with natural morphism $b\colon\operatorname{Spec}(\kappa(q))\to Y_{1}$.
+Thus, $K$ is relatively perfect over $Y_{1}$.
+Lastly, we check the third claim.
+However, we can argue essentially like above.
+In particular, the argument above shows that the stalks of $\mathbf{R}(f^{\prime}_{2})_{\ast}(K\otimes^{\mathbf{L}}P)$ at all closed points of $Y_{1}$ are perfect.
+Consequently, the desired claim follows from Proposition˜2.4.
 ∎
-Corollary 3.6.
-Consider ˜3.1 where $t$ is an affine faithfully flat morphism.
-Let $K\in D^{-}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$.
-Then $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}T$ (resp.
-$Y_{2}\times_{S}T$) if, and only if, $K$ is relatively perfect over $Y_{1}$ (resp.
-$Y_{2}$).
+Now, for the main result of the section.
+Theorem 3.6.
+Consider ˜3.1. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$. Then the following are equivalent:
+  1. (1)
+    $K$ is relatively perfect over $Y_{1}$ (resp. $Y_{2}$)
+  2. (2)
+    $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}T$ (resp. $Y_{2}\times_{S}T$) for every affine morphism $t$ from a Noetherian scheme
+  3. (3)
+    $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}\operatorname{Spec}(k)$ (resp.
+$Y_{2}\times_{S}\operatorname{Spec}(k)$) for all morphisms $t\colon\operatorname{Spec}(k)\to S$ from a field
+  4. (4)
+    $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}\operatorname{Spec}(\kappa(p))$ (resp.
+$Y_{2}\times_{S}\operatorname{Spec}(\kappa(p))$) for all closed immersions $t\colon\operatorname{Spec}(k(p))\to S$ associated to a closed point $p\in S$
+  5. (5)
+    $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}T$ (resp. $Y_{2}\times_{S}T$) for some affine surjective morphism $t\colon T\to S$.
 Proof.
-This follows from Lemma˜3.3 and Proposition˜3.5
+We only prove the case of being relatively perfect over $Y_{1}$ because the other can be shown analogously.
+By Lemma˜3.3, we see that (1) implies (2).
+Moreover, we know that (2) implies (3) by [STA25, Tag 01SI].
+Clearly, (3) implies (4).
+Furthermore, to see (4) implies (5), use that Item˜3.5(3) tells us $K$ is relatively perfect over $Y_{1}$, i.e.
+take $t=1_{S}$.
+So, we need to check (5) implies (1).
+Let $t\colon T\to S$ be an affine surjective morphism such that $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}T$.
+Choose any morphism $s\colon\operatorname{Spec}(k)\to S$ from a field.
+Consider the fiber product
+[…]
+By base change, $s^{\prime}$ is affine and $t^{\prime}$ is an affine surjection.
+To avoid complicated diagrams involving base changes, we need some notation.
+Specifically, let $K_{\#}$ be the derived pullback of $K$ along the natural morphism $Y_{1}\times_{S}Y_{2}\times_{S}\#\to Y_{1}\times_{S}Y_{2}$ for #\# any scheme appearing in the diagram above.
+Now, the hypothesis tells us that $K_{T}$ is relatively perfect over $Y_{1}\times_{S}T$.
+Hence, Lemma˜3.3 implies $K_{T\times_{S}\operatorname{Spec}(k)}$ is relatively perfect over $Y_{1}\times_{S}T\times_{S}\operatorname{Spec}(k)$ because $s^{\prime}$ is affine.
+However, $t^{\prime}$ is affine and faithfully flat, so Item˜3.5(1) ensures that $K_{\operatorname{Spec}(k)}$ is relatively perfect over $Y_{1}\times_{S}\operatorname{Spec}(k)$.
+However, $s$ was an arbitrary morphism from a field, so Item˜3.5(2) tells us $K$ is relatively perfect over $Y_{1}$ as desired.
+∎
+Next, we turn to studying adjointness under base change. First, we focus on right adjoints.
 Lemma 3.7 (Base change for relative dualizing complex).
@@
 Proof.
-This is well-known to some, but we include it for convenience.
+This is known but we include it for convenience.
 By [STA25, Tag 0B6S & Tag 0E2Z], we know that $(f^{\prime})^{!}\mathcal{O}_{S^{\prime}}$ and $f^{!}\mathcal{O}_{S}$ are respectively relative dualizing complexes for $f^{\prime}$ and $f$.
-Moreover, [Sta25, Tag 0E2Y] tells us that $\mathbf{L}(g^{\prime})^{\prime}f^{!}\mathcal{O}_{S}$ is a relative dualizing complex for $f^{\prime}$.
-However, [Sta25, Tag 0E2W] implies relative dualizing complexes are unique up to isomorphism, so $\mathbf{L}(g^{\prime})^{\prime}f^{!}\mathcal{O}_{S}\cong(f^{\prime})^{!}\mathcal{O}_{S^{\prime}}$.
+Moreover, [STA25, Tag 0E2Y] tells us that $\mathbf{L}(g^{\prime})^{\ast}f^{!}\mathcal{O}_{S}$ is a relative dualizing complex for $f^{\prime}$.
+However, [STA25, Tag 0E2W] implies relative dualizing complexes are unique up to isomorphism, and so $\mathbf{L}(g^{\prime})^{\ast}f^{!}\mathcal{O}_{S}\cong(f^{\prime})^{!}\mathcal{O}_{S^{\prime}}$.
 ∎
 Proposition 3.8.
-Consider ˜3.1 where $t$ is an affine morphism locally of finite tor-dimension (e.g.
-$t=1_{S}$).
+Consider ˜3.1 with $t$ affine.
 Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{2}$.
 Then $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ is right adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ on $D_{\operatorname{qc}}$ where $K^{\prime}:=\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$.
-In this case, $K^{\prime}\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$.
+In both cases, $K^{\prime}\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$.
 Proof.
@@
 $\cong\mathbf{R}(g^{\prime}_{2})_{\ast}\bigg(\mathbf{L}(t^{\prime})^{\ast}\big(\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})\big)\otimes^{\mathbf{L}}\mathbf{L}(g^{\prime}_{1})^{\ast}A\bigg)=:\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A).$
-Note that Lemma˜3.2, as well as affineness being stable under base change, implies each vertical morphism in Diagram˜3.1 are affine and locally of finite tor-dimension.
 Moreover, $(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}}\in D^{+}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ as upper shriek functors for proper morphisms preserve complexes with bounded below and coherent cohomology (see e.g.
@@
 $\mathbf{L}(t^{\prime})^{\ast}$ $\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$
-$\cong\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,\mathbf{L}(t^{\prime})^{\ast}(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$ ([GW23, Proposition 22.70])
+$\cong\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,\mathbf{L}(t^{\prime})^{\ast}(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$ (Lemma˜2.2)
 […]
-in $D^{+}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2}\times_{S}T)$.
 By Lemma˜3.3, we see that $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ restricts to an exact functor $\operatorname{Perf}(Y_{1}\times_{S}T)\to\operatorname{Perf}(Y_{2}\times_{S}T)$.
@@
 $\cong\mathbf{R}(g^{\prime}_{2})_{\ast}\bigg(\mathbf{L}(t^{\prime})^{\ast}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})\otimes^{\mathbf{L}}\mathbf{L}(g^{\prime}_{1})^{\ast}A\bigg).$
-The last isomorphism comes from the work above, whereas the second isomorphism is [Riz17, Lemma 2.13] together with the fact that $\mathbf{L}(t^{\prime})^{\ast}K$ is $g_{1}^{\prime}$-perfect (see see Corollary˜3.6 and the discussion in Section˜2.2), which completes the proof.
+The last isomorphism comes from the work above, whereas the second isomorphism is [RIZ17, Lemma 2.13] together with the fact that $\mathbf{L}(t^{\prime})^{\ast}K$ is $g_{1}^{\prime}$-perfect via Theorem˜3.6, which completes the proof.
 ∎
 Corollary 3.9.
-Consider ˜3.1 where $t=1_{S}$ and $S$ is semiseparated (i.e.
-that intersection of two affine open subschemes is affine, see e.g.
-[AJPV08, §2]).
+Consider ˜3.1.
 Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{2}$.
@@
 Proof.
-Let $U_{i}$ be an affine open cover of $S$.
-After base changing along each open immersion $U_{i}\to S$, we can form open covers of each $Y_{i}$.
-To check $\Phi_{K^{\prime}}(E)\in D^{b}_{\operatorname{coh}}(Y_{2})$, it suffices to show the restriction of $\Phi_{K^{\prime}}(E)$ to each component of the induced open cover on $Y_{2}$ has bounded and coherent cohomology.
-Hence, using Proposition˜3.8 and Lemma˜3.3, we can reduce to the case $S$ is affine.
+By Theorem˜3.6, it suffices to check that $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ restricts to $D^{b}_{\operatorname{coh}}$ for every $t\colon\operatorname{Spec}(\kappa(p))\to S$ for $p\in S$ a closed point.
+So, we can reduce to the case $S=\operatorname{Spec}(k)$ for $k$ a field.
 In this setting, the desired claim follows from [DLM25, Lemma 3.1].
@@
 ∎
-Remark 3.10.
-In Corollary˜3.9, we require $S$ to be semiseparated solely so that the open immersion from an affine scheme to $S$ is an affine morphism.
-This allows us to leverage Proposition˜3.8.
-Note that semiseparated is a mild constraint because any separated scheme is such.
+Next, we focus on left adjoints.
 Lemma 3.10.
@@
 Proof.
-This is baked into the proof of [DLM25, Proposition 4.3], but we spell it out for convenience.
+This is baked into the proof of [DLM25, Proposition 4.3] but we spell it out for convenience.
 In short, the desired claim follows from the string of natural isomorphisms for all $E\in D_{\operatorname{qc}}(Y_{2})$ and $G\in D_{\operatorname{qc}}(Y_{1})$:
@@
 Proposition 3.11.
-Consider ˜3.1 where $t$ is an affine morphism locally of finite tor-dimension.
+Consider ˜3.1 with $t$ affine.
 Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{1}$.
@@
 By Lemma˜3.3, $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}T$.
-Moreover, from Lemma˜3.11, the kernel of the integral transform which is left adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is given $\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,(g^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}\times_{S}T})$.
+Moreover, from Lemma˜3.10, the kernel of the integral transform which is left adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is given by the object $\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,(g^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}\times_{S}T})$.
 So, it suffices to show there is an isomorphism
@@
 $\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,(g^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}\times_{S}T})\to\mathbf{L}(t^{\prime})^{\ast}(\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}})).$
-As Lemma˜3.2 and affineness is stable under base change, it follows that each vertical morphism in Diagram˜3.1 are affine and locally of finite tor-dimension.
-Also, $(g^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}}\in D^{+}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ as upper shriek functors for proper morphisms preserve complexes with bounded below and coherent cohomology (see e.g.
+As in Lemma˜3.2, $(g^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}\times_{S}T}\in D^{+}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2}\times_{S}T)$ because upper shriek functors for proper morphisms preserve complexes with bounded below and coherent cohomology (see e.g.
 [STA25, Tag 0AU1]).
@@
 $\mathbf{L}(t^{\prime})^{\ast}$ $\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}})$
-$\cong\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,\mathbf{L}(t^{\prime})^{\ast}(f^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}})$ ([GW23, Proposition 22.70])
+$\cong\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,\mathbf{L}(t^{\prime})^{\ast}(f^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}})$ (Lemma˜2.2)
 […]
-in $D_{\operatorname{coh}}^{+}$. This completes the proof. ∎
+This completes the proof. ∎
 4. Fully faithfulness & equivalences
 This section is concerned with the behavior of fully faithfulness or an equivalence for an integral transform under a change of base scheme.
+Lemma 4.1 ([MAC78, IV.7, Exercise 4]).
+Consider a diagram of functors and adjoint functors
+(4.1) $\mathcal{C}$$\mathcal{C}^{\prime}$$\mathcal{D}$$\mathcal{D}^{\prime}.$$\scriptstyle{H}$$\scriptstyle{F}$$\scriptstyle{F^{\prime}}$$\scriptstyle{G}$$\scriptstyle{K}$$\scriptstyle{G^{\prime}}\scriptstyle{\dashv}\scriptstyle{\dashv}$
+Let $\beta\colon F^{\prime}H\Rightarrow KF$ and $\alpha\colon HG\Rightarrow G^{\prime}K$ be a pair of natural transformations. Then the following conditions are equivalent:
+  1. (1)
+    The following diagram commutes:
+[…]
+  2. (2)
+    The following diagram commutes:
+[…]
+  3. (3)
+    For all objects $c\in\mathcal{C}$, $d\in\mathcal{D}$, the following diagram commutes:
+[…]
+Proof.
+[…]
+$f\mapsto Gf\cdot\eta_{c}\mapsto H(Gf\cdot\eta_{c})\mapsto\alpha_{d}\cdot H(Gf\cdot\eta_{c})=\alpha_{d}\cdot HGf\cdot H\eta_{c}.$
+Moreover, $f$ is assigned along the left bottom composite as
+$f\mapsto Kf\mapsto Kf\cdot\beta_{c}\mapsto G^{\prime}(Kf\cdot\beta_{c})\cdot\eta^{\prime}_{Hc}=G^{\prime}Kf\cdot G^{\prime}\beta_{c}\cdot\eta^{\prime}_{Hc}.$
+So, by the naturality of $\alpha$ and (4.1(2)), we have
+$G^{\prime}Kf\cdot G^{\prime}\beta_{c}\cdot\eta^{\prime}_{Hc}=G^{\prime}Kf\cdot\alpha_{Fc}\cdot H\eta_{c}=\alpha_{d}\cdot HGf\cdot H\eta_{c}.$
+[…]
+[…]
+Definition 4.2.
+A (lax) morphism of adjunctions from $F\dashv G$ to $F^{\prime}\dashv G^{\prime}$ is comprised of a pair of functors
+$\mathcal{C}$$\mathcal{C}^{\prime}$$\mathcal{D}$$\mathcal{D}^{\prime}$$\scriptstyle{H}$$\scriptstyle{F}$$\scriptstyle{F^{\prime}}$$\scriptstyle{G}$$\scriptstyle{K}$$\scriptstyle{G^{\prime}}\scriptstyle{\dashv}\scriptstyle{\dashv}$
+plus a pair of natural transformations $\beta\colon F^{\prime}H\Rightarrow KF$, $\alpha\colon HG\Rightarrow G^{\prime}K$ such that the equivalent conditions in Lemma 4.1 hold.
+We call $\beta$ (resp.
+$\alpha$) the left (resp.
+right) comparison transformation.
+Furthermore, a strict morphism of adjunctions is just a lax morphism of adjunctions such that $F^{\prime}H=KF$, $HG=G^{\prime}K$, and $\beta$, $\alpha$ are the identity transformation.
+Remark 4.3.
+In the category theory literature, a ‘morphism of adjunctions’ is usually referred in its strict sense [MAC78, IV.7] and [RIE17, Exercise 4.2.v], but a ‘lax’ one is instead described by saying that $\beta$ and $\alpha$ are mates under Diagram˜4.1.
+See [CGR14, Sect.
+1] for the precise definition of mates.
+Indeed, the mate $\overline{\beta}$ of $\beta$ is by definition the natural transformation
+$\overline{\beta}\colon HG\xrightarrow{\eta^{\prime}_{HG}}G^{\prime}F^{\prime}HG\xrightarrow{G^{\prime}\beta_{G}}G^{\prime}KFG\xrightarrow{G^{\prime}K\varepsilon}G^{\prime}K.$
+Therefore, $\overline{\beta}$ will equal $\alpha$ if, and only if, the corresponding adjunct morphisms $F^{\prime}HG\to K$ (with respect to $F^{\prime}\dashv G^{\prime}$) are equal.
+This is exactly Item˜4.1(1).
+Dually, one also sees that Item˜4.1(2) amounts to the mate $\overline{\alpha}$ of $\alpha$ being equal to $\beta$.
 Lemma 4.4.
-Consider ˜3.1.
-Let $K\in D_{\operatorname{qc}}(Y_{1}\times_{S}Y_{2})$.
-Then there is a natural isomorphism $\alpha^{K}\colon\mathbf{R}(t_{2})_{\ast}\circ\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\to\Phi_{K}\circ\mathbf{R}(t_{1})_{\ast}$ on $D_{\operatorname{qc}}$.
+Consider ˜3.1. Let $K\in D_{\operatorname{qc}}(Y_{1}\times_{S}Y_{2})$. Then on $D_{\operatorname{qc}}$ there is a natural isomorphism
+$\alpha^{K}\colon\Phi_{K}\circ\mathbf{R}(t_{1})_{\ast}\to\mathbf{R}(t_{2})_{\ast}\circ\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}.$
 Proof.
 This follows from the string of natural isomorphisms for each $E\in D_{\operatorname{qc}}(Y_{1}\times_{S}T)$:
-$\mathbf{R}(t_{2})_{\ast}$ $\circ\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}(E)$
-$=\mathbf{R}(t_{2})_{\ast}\mathbf{R}(g^{\prime}_{1})_{\ast}\big(\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K\big)$
-$\cong\mathbf{R}(t_{2}\circ g^{\prime}_{1})_{\ast}\big(\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K\big)$
-$\cong\mathbf{R}(f_{1}^{\prime}\circ t^{\prime})_{\ast}\big(\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K\big)$
-$\cong\mathbf{R}(f_{1}^{\prime})_{\ast}\bigg(\mathbf{R}(t^{\prime})_{\ast}\big(\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K\big)\bigg)$ ([GW23, Proposition 22.84])
-$\cong\mathbf{R}(f_{1}^{\prime})_{\ast}\bigg(\mathbf{R}(t^{\prime})_{\ast}\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}K\bigg)$
-$\cong\mathbf{R}(f_{1}^{\prime})_{\ast}\bigg(\mathbf{L}(f^{\prime}_{2})^{\ast}\mathbf{R}(t_{1})_{\ast}E\otimes^{\mathbf{L}}K\bigg)$
-$=\Phi_{K}\circ\mathbf{R}(t_{1})_{\ast}(E).$
+$\Phi_{K}\circ\mathbf{R}(t_{1})_{\ast}(E)$ $=\mathbf{R}(f_{1}^{\prime})_{\ast}\big(\mathbf{L}(f^{\prime}_{2})^{\ast}\mathbf{R}(t_{1})_{\ast}E\otimes^{\mathbf{L}}K\big)$
+$\cong\mathbf{R}(f_{1}^{\prime})_{\ast}\big(\mathbf{R}t^{\prime}_{\ast}\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}K\big)$ (flat base change)
+$\cong\mathbf{R}(f_{1}^{\prime})_{\ast}\mathbf{R}t^{\prime}_{\ast}\big(\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K\big)$ (projection formula)
+$\cong\mathbf{R}(t_{2})_{\ast}\mathbf{R}(g^{\prime}_{1})_{\ast}\big(\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K\big)$ (pseudofunctoriality of $\mathbf{R}(-)_{\ast}$)
+$=\mathbf{R}(t_{2})_{\ast}\circ\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}(E).$
+Lemma 4.5.
+Consider ˜3.1. Let $K\in D_{\operatorname{qc}}(Y_{1}\times_{S}Y_{2})$. Then we have a morphism of adjoint pairs as depicted in the diagram
+$D_{\operatorname{qc}}(Y_{1})$$D_{\operatorname{qc}}(Y_{2})$$D_{\operatorname{qc}}(Y_{1}\times_{S}T)$$D_{\operatorname{qc}}(Y_{2}\times_{S}T)$$\scriptstyle{\Phi_{K}}$$\scriptstyle{\mathbf{L}t_{1}^{\ast}}$$\scriptstyle{\mathbf{L}t_{2}^{\ast}}$$\scriptstyle{\mathbf{R}(t_{1})_{\ast}}$$\scriptstyle{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}}$$\scriptstyle{\mathbf{R}(t_{2})_{\ast}}\scriptstyle{\dashv}\scriptstyle{\dashv}$
+with left and right comparison transformations given respectively by the isomorphisms $\beta^{K}$, $\alpha^{K}$ of Lemmas˜4.4 and 3.4.
+Proof.
+Denote by $\xi^{i}\colon\mathbf{L}t_{i}^{\ast}\mathbf{R}(t_{i})_{\ast}\to\operatorname{id}$ for the counit of the adjunction $\mathbf{L}t_{i}^{\ast}\dashv\mathbf{R}(t_{i})_{\ast}$ with $i=1,2$.
+To prove the desired claim, it suffices to verify that Item˜4.1(1) holds; that is, that the following diagram commutes:
+$\mathbf{L}t_{2}^{\ast}\Phi_{K}\mathbf{R}(t_{1})_{\ast}$$\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\mathbf{L}t_{1}^{\ast}\mathbf{R}(t_{1})_{\ast}$$\mathbf{L}t_{2}^{\ast}\mathbf{R}(t_{2})_{\ast}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$$\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}.$$\scriptstyle{\beta^{K}_{\mathbf{R}(t_{1})_{\ast}}}$$\scriptstyle{\mathbf{L}t_{2}^{\ast}(\alpha^{K})}$$\scriptstyle{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}(\xi^{1})}$$\scriptstyle{\xi^{2}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}}}$
+For brevity sake, in the remainder of the proof, we drop the $\mathbf{L}$’s and $\mathbf{R}$’s in the notation for the derived functors (i.e.
+all functors now are understood to be derived).
+Hence, the diagram above is now
+(4.2) $t_{2}^{\ast}(t_{2})_{\ast}\Phi_{(t^{\prime})^{\ast}K}$$t_{2}^{\ast}\Phi_{K}(t_{1})_{\ast}$$\Phi_{(t^{\prime})^{\ast}K}.$$\Phi_{(t^{\prime})^{\ast}K}t_{1}^{\ast}(t_{1})^{\ast}$$\scriptstyle{\xi^{2}_{\Phi_{(t^{\prime})^{\ast}K}}}$$\scriptstyle{t_{2}^{\ast}(\alpha^{K})}$$\scriptstyle{\beta^{K}_{(t_{1})_{\ast}}}$$\scriptstyle{\Phi_{(t^{\prime})^{\ast}K}(\xi^{1})}$
+Choose $E\in D_{\operatorname{qc}}(Y_{1}\times_{S}T)$.
+If we expand the definition of $\alpha^{K}$ and $\beta^{K}$ in Diagram˜4.2 (see the proofs of Lemmas Lemmas˜4.4 and 3.4), we obtain a large diagram consisting of various faces.
+In what follows, we make this a bit more explicit.
+Particularly, we spell out each face needed to understand Diagram˜4.2.
+To describe said diagram more explicitly, we use the following abbreviations for the (natural) isomorphisms:
+  • $\operatorname{BC}_{i}$, $i\in\{1,2\}$ for the flat base change isomorphisms
+  • $\operatorname{PF}$ for the pojection formula
+  • $F_{(-)_{\ast}}$ for functoriality of $(-)_{\ast}$
+  • $F_{(-)^{\ast}}$ for functoriality of $(-)^{\ast}$
+  • $M_{(-)^{\ast}}$ for monoidality of $(-)^{\ast}$
+  • $\xi^{\prime}$ for the counit of $(t^{\prime})^{\ast}\dashv(t^{\prime})_{\ast}$.
+Now, consider the following diagrams:
+(4.3) $t_{2}^{\ast}(f_{1}^{\prime})_{\ast}((f_{2}^{\prime})_{\ast}(t_{1})_{\ast}E\otimes K)$$t_{2}^{\ast}(f_{1}^{\prime})_{\ast}((t^{\prime})_{\ast}(g_{2}^{\prime})^{\ast}E\otimes K)$$(g_{1}^{\prime})_{\ast}(t^{\prime})^{\ast}((f_{2}^{\prime})^{\ast}(t_{1})_{\ast}E\otimes K)$$(g_{1}^{\prime})_{\ast}(t^{\prime})^{\ast}(t_{\ast}^{\prime}(g_{2}^{\prime})^{\ast}E\otimes K)$$\scriptstyle{\operatorname{BC}_{1}}$$\scriptstyle{\operatorname{BC}_{2}}$$\scriptstyle{\operatorname{BC}_{2}}$$\scriptstyle{\operatorname{BC}_{1}}$
+(4.4) $t_{2}^{\ast}(f_{1}^{\prime})_{\ast}((t^{\prime})_{\ast}(g_{2}^{\prime})^{\ast}E\otimes K)$$t_{2}^{\ast}(f_{1}^{\prime})_{\ast}(t^{\prime})_{\ast}((g_{2}^{\prime})^{\ast}E\otimes(t^{\prime})^{\ast}K)$$(g_{1}^{\prime})_{\ast}(t^{\prime})^{\ast}(t_{\ast}^{\prime}(g_{2}^{\prime})^{\ast}E\otimes K)$$(g_{1}^{\prime})_{\ast}(t^{\prime})^{\ast}t_{\ast}^{\prime}((g_{2}^{\prime})^{\ast}E\otimes(t^{\prime})^{\ast}K)$$\scriptstyle{\operatorname{PF}}$$\scriptstyle{\operatorname{BC}_{2}}$$\scriptstyle{\operatorname{BC}_{2}}$$\scriptstyle{\operatorname{PF}}$
+(4.5) $(g_{1}^{\prime})_{\ast}(t^{\prime})^{\ast}((f_{2}^{\prime})^{\ast}(t_{1})_{\ast}E\otimes K)$$(g_{1}^{\prime})_{\ast}(t^{\prime})^{\ast}(t_{\ast}^{\prime}(g_{2}^{\prime})^{\ast}E\otimes K)$$(g_{1}^{\prime})_{\ast}((t^{\prime})^{\ast}(f_{2}^{\prime})^{\ast}(t_{1})_{\ast}E\otimes(t^{\prime})^{\ast}K)$$(g_{1}^{\prime})_{\ast}((t^{\prime})^{\ast}t_{\ast}^{\prime}(g_{2}^{\prime})^{\ast}E\otimes(t^{\prime})^{\ast}K)$$\scriptstyle{\operatorname{BC}_{1}}$$\scriptstyle{M_{(-)^{\ast}}}$$\scriptstyle{M_{(-)}^{\ast}}$$\scriptstyle{\operatorname{BC}_{1}}$
+(4.6) $(g_{1}^{\prime})_{\ast}((t^{\prime})^{\ast}t_{\ast}^{\prime}(g_{2}^{\prime})^{\ast}E\otimes(t^{\prime})^{\ast}K)$$(g_{1}^{\prime})_{\ast}(t^{\prime})^{\ast}(t_{\ast}^{\prime}(g_{2}^{\prime})^{\ast}E\otimes K)$$(g_{1}^{\prime})_{\ast}((g_{2}^{\prime})^{\ast}E\otimes(t^{\prime})^{\ast}K)$$(g_{1}^{\prime})_{\ast}(t^{\prime})^{\ast}t_{\ast}^{\prime}((g_{2}^{\prime})^{\ast}E\otimes(t^{\prime})^{\ast}K)$$\scriptstyle{\xi^{\prime}}$$\scriptstyle{M_{(-)}^{\ast}}$$\scriptstyle{\operatorname{PF}}$$\scriptstyle{\xi^{\prime}}$
+(4.7) $t_{2}^{\ast}(f_{1}^{\prime})_{\ast}(t^{\prime})_{\ast}((g_{2}^{\prime})^{\ast}E\otimes(t^{\prime})^{\ast}K)$$t_{2}^{\ast}(t_{2})_{\ast}(g_{1}^{\prime})_{\ast}((g_{2}^{\prime})^{\ast}E\otimes(t^{\prime})^{\ast}K)$$(g_{1}^{\prime})_{\ast}(t^{\prime})^{\ast}t_{\ast}^{\prime}((g_{2}^{\prime})^{\ast}E\otimes(t^{\prime})^{\ast}K)$$(g_{1}^{\prime})_{\ast}((g_{2}^{\prime})^{\ast}E\otimes(t^{\prime})^{\ast}K)$$\scriptstyle{F_{(-)^{\ast}}}$$\scriptstyle{\operatorname{BC}_{2}}$$\scriptstyle{\xi^{2}}$$\scriptstyle{\xi^{\prime}}$
+(4.8) $(g_{1}^{\prime})_{\ast}((t^{\prime})^{\ast}(f_{2}^{\prime})^{\ast}(t_{1})_{\ast}E\otimes(t^{\prime})^{\ast}K)$$(g_{1}^{\prime})_{\ast}((t^{\prime})^{\ast}t_{\ast}^{\prime}(g_{2}^{\prime})^{\ast}E\otimes(t^{\prime})^{\ast}K)$$(g^{\prime}_{1})_{\ast}((g_{2}^{\prime})^{\ast}t_{1}^{\ast}(t_{1})_{\ast}E\otimes(t^{\prime})^{\ast}K)$$(g_{1}^{\prime})_{\ast}((g_{2}^{\prime})^{\ast}E\otimes(t^{\prime})^{\ast}K)$$\scriptstyle{\operatorname{BC}_{1}}$$\scriptstyle{F_{(-)^{\ast}}}$$\scriptstyle{\xi^{\prime}}$$\scriptstyle{\xi^{1}}$
+Assume that we have shown Diagrams˜4.3, 4.4, 4.5, 4.6, 4.7 and 4.8 are commutative. Then Diagram˜4.2 (in the case of $E$) is the pasting of these diagrams, i.e. gives the desired morphism
+$t_{2}^{\ast}(f_{1}^{\prime})_{\ast}((f_{2}^{\prime})_{\ast}(t_{1})_{\ast}E\otimes K)$$(g_{1}^{\prime})_{\ast}((g_{2}^{\prime})^{\ast}E\otimes(t^{\prime})^{\ast}K).$}}
+So, let us explain why Diagrams˜4.3, 4.4, 4.5, 4.6, 4.7 and 4.8 are commutative.
+As for Diagram˜4.3, one may use naturality of $\operatorname{BC}_{2}$, whereas Diagram˜4.4 is due to the naturality of $\operatorname{BC}_{2}$.
+Also, for Diagram˜4.5, one can use the naturality of $M_{(-)^{\ast}}$.
+Next, we explain Diagram˜4.6. Observe that it is the functor $(g_{1}^{\prime})_{\ast}$ applied to the following diagram evaluated at $((g_{2}^{\prime})^{\ast}E,K)$:
+$(t^{\prime})^{\ast}(t_{\ast}^{\prime}(-)\otimes-)$$(t^{\prime})^{\ast}t_{\ast}^{\prime}(-\otimes(t^{\prime})^{\ast}(-))$$(t^{\prime})^{\ast}t_{\ast}^{\prime}(-)\otimes(t^{\prime})^{\ast}(-)$$(-)\otimes(t^{\prime})^{\ast}(-)$$\scriptstyle{M_{(-)^{\ast}}}$$\scriptstyle{\xi^{\prime}}$$\scriptstyle{\xi^{\prime}}$
+which commutes by definition of the projection formula, see e.g. proof of [GW23, Proposition 22.81].
+Now, for Diagram˜4.7. Note that it is the following diagram evaluated at $(g_{2}^{\prime})^{\ast}E\otimes(t^{\prime})^{\ast}K$:
+[…]
+which commutes by definition of the base change morphism. See [GW23, Definition and Remark 21.129] or [LIP09, (3.7.2)].
+Lastly, we check Diagram˜4.8. However, it is the functor $(g_{1}^{\prime})_{\ast}(-\otimes(t^{\prime})^{\ast}K)$ applied to the following diagram evaluated at $E$:
+[…]
+which commutes by definition of the base change morphism. ∎
+Remark 4.6.
+Assume the hypothesis of Lemma˜4.5.
+Denote by $\zeta^{i}$ for the unit of $t_{i}^{\ast}\dashv(t_{i})_{\ast}$ if $i\in\{1,2\}$.
+From Lemmas˜4.1 and 4.5, we have that the following diagram is commutative:
+(4.9) $\Phi_{K}$$(t_{2})_{\ast}t_{2}^{\ast}\Phi_{K}$$\Phi_{K}(t_{1})_{\ast}t_{1}^{\ast}$$(t_{2})_{\ast}\Phi_{(t^{\prime})^{\ast}K}t_{1}^{\ast}.$$\scriptstyle{\zeta^{2}_{\Phi_{K}}}$$\scriptstyle{\Phi_{K}(\zeta^{1})}$$\scriptstyle{(t_{2})_{\ast}(\beta^{K})}$$\scriptstyle{\alpha^{K}_{t_{1}^{\ast}}}$
 Proposition 4.7.
-Consider ˜3.1 where $t$ is an affine morphism locally of finite tor-dimension.
+Consider ˜3.1 where $t$ is affine.
 Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{2}$.
 Denote by $K^{\prime}$ for the kernel of the integral transform obtained in Proposition˜3.8 which is right adjoint to $\Phi_{K}$ on $D_{\operatorname{qc}}$.
-If $\Phi_{K}$ is fully faithful on $D_{\operatorname{qc}}$, then $\mathbf{L}t^{\ast}_{1}$ preserves the unit $\eta$ of the adjoint pair $\Phi_{K}$ and $\Phi_{K^{\prime}}$ on $D_{\operatorname{qc}}$.
-That is, $\mathbf{L}t_{1}^{\ast}\eta_{E}$ is the unit of $\mathbf{L}t_{1}^{\ast}E$ for the adjoint pair $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ and $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$.
+Then, after possibly replacing $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ by an isomorphic functor, we have a morphism of adjoint pairs:
+$D_{\operatorname{qc}}(Y_{1})$$D_{\operatorname{qc}}(Y_{1}\times_{S}T)$$D_{\operatorname{qc}}(Y_{2})$$D_{\operatorname{qc}}(Y_{2}\times_{S}T)$$\scriptstyle{\mathbf{L}t_{1}^{\ast}}$$\scriptstyle{\Phi_{K}}$$\scriptstyle{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}}$$\scriptstyle{\Phi_{K^{\prime}}}$$\scriptstyle{\mathbf{L}t_{2}^{\ast}}$$\scriptstyle{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}}\scriptstyle{\dashv}\scriptstyle{\dashv}$
+with left and right comparison transformations given respectively by the isomorphisms
+$(\beta^{K})^{-1}\colon\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\circ\mathbf{L}t_{1}^{\ast}\to\mathbf{L}t_{2}^{\ast}\circ\Phi_{K},$
+$\beta^{K^{\prime}}\colon\mathbf{L}t_{1}^{\ast}\circ\Phi_{K^{\prime}}\to\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}\circ\mathbf{L}t_{2}^{\ast}.$
 Proof.
-By Proposition˜3.8, $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ and $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ are adjoint functors on $D_{\operatorname{qc}}$.
-Choose $E\in D_{\operatorname{qc}}(Y_{1})$.
-We want to show that $\mathbf{L}t_{1}^{\ast}(\eta_{E})$ is the unit of $\mathbf{L}t_{1}^{\ast}E$ for this adjunction.
-Let $\beta\colon\mathbf{L}t^{\ast}_{1}E\to\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$ be a morphism in $D_{\operatorname{qc}}(Y_{1}\times_{S}T)$ for any $A\in D_{\operatorname{qc}}(Y_{2}\times_{S}T)$.
-Using that $\mathbf{L}t_{1}^{\ast}$ is left adjoint to $\mathbf{R}(t_{1})_{\ast}$ on $D_{\operatorname{qc}}$, adjunction tells us there exists a unique morphism $\beta^{\prime}$ of the form $\mathbf{L}t^{\ast}_{1}(\beta^{\prime\prime})$ where $\beta^{\prime\prime}\colon E\to\mathbf{R}(t_{1})_{\ast}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$, making the following diagram commute,
-$\mathbf{L}t^{\ast}_{1}E$$\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$$\mathbf{L}t_{1}^{\ast}\mathbf{R}(t_{1})_{\ast}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$$\scriptstyle{\beta}$$\scriptstyle{\beta^{\prime}}$$\scriptstyle{\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}}$
-where $\xi^{t_{1}}$ is the counit of derived pullback and pushforward along $t_{1}$.
-Furthermore, by the universal property of the counit, the morphism $\beta^{\prime\prime}$ is the unique morphism such that $\beta=\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\mathbf{L}t^{\ast}_{1}(\beta^{\prime\prime})$.
-Similarly, using that $\Phi_{K}$ is left adjoint to $\Phi_{K^{\prime}}$ and the natural isomorphism $\alpha$ of Lemma˜4.1, there is a $\beta^{\prime\prime\prime}$ giving a diagram
-[…]
-which commutes. Pulling back this diagram via $\mathbf{L}t_{1}^{\ast}$, we obtain
-$\beta^{\prime}=\mathbf{L}t_{1}^{\ast}((\alpha^{A})^{-1}\circ\beta^{\prime\prime\prime})\circ\mathbf{L}t_{1}^{\ast}(\eta_{E}),$
-which tells us $\beta$ factors through $\mathbf{L}t_{1}^{\ast}(\eta_{E})$.
-It remains to show that the above factorization for $\beta^{\prime}$ is unique.
-So, suppose there is a morphism $h\colon\mathbf{L}t_{1}^{\ast}\Phi_{K^{\prime}}\Phi_{K}(E)\to\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$ such that $\beta=h\circ\mathbf{L}t_{1}^{\ast}(\eta_{E})$.
-Once again, using that $\mathbf{L}t_{1}^{\ast}$ is left adjoint to $\mathbf{R}(t_{1})_{\ast}$, we have a morphism
-$h^{\prime}\colon\mathbf{L}t_{1}^{\ast}\Phi_{K^{\prime}}\Phi_{K}(E)\to\mathbf{L}t_{1}^{\ast}\mathbf{R}(t_{1})_{\ast}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$
-giving a commutative diagram
-$\mathbf{L}t_{1}^{\ast}\Phi_{K^{\prime}}\Phi_{K}(E)$$\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$$\mathbf{L}t_{1}^{\ast}\mathbf{R}(t_{1})_{\ast}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$$\scriptstyle{h}$$\scriptstyle{h^{\prime}}$$\scriptstyle{\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}}$
-with $\xi^{t_{1}}$ as above and $h^{\prime}=\mathbf{L}t^{\ast}_{1}h^{\prime\prime}$ for a morphism $h^{\prime\prime}\colon\Phi_{K^{\prime}}\Phi_{K}(E)\to\mathbf{R}(t_{1})_{\ast}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$.
-Here, $h^{\prime\prime}$ is the unique morphism such that $h=\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\mathbf{L}t^{\ast}_{1}h^{\prime\prime}$.
-Tying things together, we see that
-$\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\mathbf{L}t^{\ast}_{1}(h^{\prime\prime}\circ\eta_{E})=(\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\mathbf{L}t^{\ast}_{1}(h^{\prime\prime}))\circ\mathbf{L}t^{\ast}_{1}(\eta_{E})=h\circ\mathbf{L}t^{\ast}_{1}(\eta_{E})=\beta.$
-However, by the work above, we know that
-$\beta=\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\beta^{\prime}=\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\mathbf{L}t_{1}^{\ast}((\alpha^{A})^{-1}\circ\beta^{\prime\prime\prime})\circ\mathbf{L}t_{1}^{\ast}(\eta_{E}),$
-So, from uniqueness, it follows that
-$h^{\prime\prime}\circ\eta_{E}=(\alpha^{A})^{-1}\circ\beta^{\prime\prime\prime}\circ\eta_{E}.$
-As $\Phi_{K}$ is fully faithful on $D^{b}_{\operatorname{coh}}$, $\eta_{E}$ must be an isomorphism, which implies $h^{\prime\prime}=(\alpha^{A})^{-1}\circ\beta^{\prime\prime\prime}$.
+Write $\eta$ (resp.
+$\eta^{\prime}$) for the unit of the adjunction $\Phi_{K}\dashv\Phi_{K^{\prime}}$ (resp.
+of $\Phi_{(t^{\prime})^{\ast}K}\dashv\Phi_{(t^{\prime})^{\ast}K^{\prime}}$).
+By Item˜4.1(2), proving the Proposition amounts to showing the following diagram commutes:
+(4.10) $\mathbf{L}t_{1}^{\ast}$$\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\mathbf{L}t_{1}^{\ast}$$\mathbf{L}t_{1}^{\ast}\Phi_{K^{\prime}}\Phi_{K}$$\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}\mathbf{L}t_{2}^{\ast}\Phi_{K}.$$\scriptstyle{\eta^{\prime}_{\mathbf{L}t_{1}^{\ast}}}$$\scriptstyle{\mathbf{L}t_{1}^{\ast}(\eta)}$$\scriptstyle{\beta^{K^{\prime}}_{\Phi_{K}}}$$\scriptstyle{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(\beta^{K})}$
+As in the proof of Lemma˜4.5, we omit $\mathbf{L}$’s and $\mathbf{R}$’s in the notation for the derived functors to ease notation.
+Now, the bad news is that Diagram˜4.10 needs not commute.
+Although, the good news is that it does commute if we replace $\Phi_{(t^{\prime})^{\ast}K}$ by some appropriate isomorphic functor.
+Specifically, we claim there is an adjunction $\widetilde{\Phi}_{(t^{\prime})^{\ast}K}\dashv\Phi_{(t^{\prime})^{\ast}K^{\prime}}$ with unit $\tilde{\eta}^{\prime}$ making Diagram˜4.10 commute (i.e.
+inscribing $\widetilde{(-)}$ to $\Phi_{(t^{\prime})^{\ast}K}$ and $\eta^{\prime}$) and that moreover there is an isomorphism $\theta:\Phi_{(t^{\prime})^{\ast}K}\cong\widetilde{\Phi}_{(t^{\prime})^{\ast}K}$ satisfying
+(4.11) $\Phi_{(t^{\prime})^{\ast}K^{\prime}}(\theta)\circ\eta^{\prime}=\tilde{\eta}^{\prime}$.
+Assume that for every $E\in D_{\operatorname{qc}}(Y_{1})$ we have shown:
+($\mathcal{P}(E)$) The object $\Phi_{(t^{\prime})^{\ast}K}t_{1}^{\ast}E$ with the morphism $\Phi_{(t^{\prime})^{\ast}K}(\beta^{K}_{E})\circ\beta^{K^{\prime}}_{\Phi_{K}E}\circ t_{1}^{\ast}(\eta_{E})$ is a reflection of $t_{1}^{\ast}E$ along $\Phi_{(t^{\prime})^{\ast}K^{\prime}}$ (in the sense of [BOR94, Definition 3.1.1]).
+Then we can define $\widetilde{\Phi}_{(t^{\prime})^{\ast}K}$ by means of [BOR94, Proposition 3.1.3].
+On the one hand, we have $\Phi_{(t^{\prime})^{\ast}K}\dashv\Phi_{(t^{\prime})^{\ast}K^{\prime}}$ with unit $\eta^{\prime}$; thus by [BOR94, Theorem 3.1.5] the pair $(\Phi_{(t^{\prime})^{\ast}K}A,\eta^{\prime}_{A})$ is a reflection of $A\in D_{\operatorname{qc}}(Y_{1}\times_{S}T)$ along $\Phi_{(t^{\prime})^{\ast}K^{\prime}}$.
+Now define $\tilde{\eta}^{\prime}=\{\tilde{\eta}^{\prime}_{A}\}_{A\in D_{\operatorname{qc}}(Y_{1}\times_{S}T)}$ by setting $\tilde{\eta}^{\prime}_{A}:=\eta^{\prime}_{A}$ if $A$ is not in the image of $t_{1}^{\ast}:D_{\operatorname{qc}}(Y_{1})\to D_{\operatorname{qc}}(Y_{1}\times_{S}T)$ and setting $\tilde{\eta}^{\prime}_{t_{1}^{\ast}E}$, $E\in D_{\operatorname{qc}}(Y_{1})$, to be the unique top morphism in Diagram˜4.10 making it commute.
+By [BOR94, Proposition 3.1.3], there is a unique functor $\widetilde{\Phi}_{(t^{\prime})^{\ast}K}:D_{\operatorname{qc}}(Y_{1}\times_{S}T)\to D_{\operatorname{qc}}(Y_{2}\times_{S}T)$ whose action on objects coincides with $\Phi_{(t^{\prime})^{\ast}K}$ and such that $\tilde{\eta}^{\prime}:\operatorname{id}\to\widetilde{\Phi}_{(t^{\prime})^{\ast}K^{\prime}}\Phi_{(t^{\prime})^{\ast}K}$ is a natural transformation.
+Moreover, by [BOR94, Theorem 3.1.5], we have $\widetilde{\Phi}_{(t^{\prime})^{\ast}K}\dashv\Phi_{(t^{\prime})^{\ast}K^{\prime}}$ with unit $\tilde{\eta}^{\prime}$.
+Finally, by [RIE17, Proposition 4.4.1], there is a natural isomorphism $\theta:\Phi_{(t^{\prime})^{\ast}K}\cong\widetilde{\Phi}_{(t^{\prime})^{\ast}K}$ such that (4.11) holds.
+It is left for us to show $\mathcal{P}(E)$.
+So, let $A\in D_{\operatorname{qc}}(Y_{2})$ and consider some $\gamma\colon t_{1}^{\ast}E\to\Phi_{(t^{\prime})^{\ast}K^{\prime}}A$.
+Our goal is to check that there is a unique morphism $\tilde{\gamma}\colon\Phi_{(t^{\prime})^{\ast}K}t_{1}^{\ast}E\to A$ such that the diagram
+(4.12) $t_{1}^{\ast}E$$t_{1}^{\ast}\Phi_{K^{\prime}}\Phi_{K}E$$\Phi_{(t^{\prime})^{\ast}K^{\prime}}A$$\Phi_{(t^{\prime})^{\ast}K^{\prime}}t_{2}^{\ast}\Phi_{K}E$$\Phi_{(t^{\prime})^{\ast}K^{\prime}}\Phi_{(t^{\prime})^{\ast}K}t_{1}^{\ast}E$$\scriptstyle{t_{1}^{\ast}(\eta_{E})}$$\scriptstyle{\gamma}$$\scriptstyle{\beta^{K^{\prime}}_{\Phi_{K}(E)}}$$\scriptstyle{\Phi_{(t^{\prime})^{\ast}K^{\prime}}(\beta^{K}_{E})}$$\scriptstyle{\Phi_{(t^{\prime})^{\ast}K^{\prime}}(\tilde{\gamma})}$
+commutes. Let $i\in\{1,2\}$. Denote $\zeta^{i}$ to the unit of $t_{i}^{\ast}\dashv(t_{i})_{\ast}$. Given morphisms
+$\phi\colon t^{\ast}_{i}M\to B,\quad\psi\colon C\to(t_{i})_{\ast}D,\quad\rho\colon L\to\Phi_{K^{\prime}}P,$
+we will use notations
+$\phi^{\flat_{i}}\colon M\to(t_{i})_{\ast}B,\quad\psi^{\sharp_{i}}\colon t_{i}^{\ast}C\to D,\quad\rho^{\sharp}\colon\Phi_{K}L\to P$
+for the adjunct morphisms of $\phi$, $\psi$, $\rho$ with respect to the adjunctions $t_{i}^{\ast}\dashv(t_{i})_{\ast}$ and $\Phi_{K}\dashv\Phi_{K^{\prime}}$.
+Suppose there is $\tilde{\gamma}$ turning Diagram˜4.12 into a commutative diagram.
+Then $\gamma^{\flat_{1}}\colon E\to(t_{1})_{\ast}\Phi_{(t^{\prime})^{\ast}K^{\prime}}A$ equals the outer clockwise composite in the following diagram:
+$E{{(t_{1})_{\ast}t_{1}^{\ast}E}}$$\Phi_{K^{\prime}}\Phi_{K}E$$(t_{1})_{\ast}t_{1}^{\ast}\Phi_{K^{\prime}}\Phi_{K}E$$\Phi_{K^{\prime}}(t_{2})_{\ast}t^{\ast}_{2}\Phi_{K}E$$(t_{1})_{\ast}\Phi_{(t^{\prime})^{\ast}K^{\prime}}t_{2}^{\ast}\Phi_{K}E$$\Phi_{K^{\prime}}(t_{2})_{\ast}\Phi_{(t^{\prime})^{\ast}K}t_{1}^{\ast}E$$(t_{1})_{\ast}\Phi_{(t^{\prime})^{\ast}K^{\prime}}\Phi_{(t^{\prime})^{\ast}K}t_{1}^{\ast}E$$\Phi_{K^{\prime}}(t_{2})_{\ast}A$$(t_{1})_{\ast}\Phi_{(t^{\prime})^{\ast}K^{\prime}}A.$$\scriptstyle{\zeta_{E}^{1}}$$\scriptstyle{\eta_{E}}$$\scriptstyle{\delta}$$\scriptstyle{(t_{1})_{\ast}t_{1}^{\ast}(\eta_{E})}$$\scriptstyle{\zeta^{1}_{\Phi_{K^{\prime}}\Phi_{K}E}}$$\scriptstyle{\Phi_{K^{\prime}}(\zeta^{2}_{\Phi_{K}E})}$$\scriptstyle{(t_{1})_{\ast}(\beta^{K^{\prime}}_{\Phi_{K}(E)})}$$\scriptstyle{\alpha^{K^{\prime}}_{t^{\ast}_{2}\Phi_{K}E}}$$\scriptstyle{\Phi_{K^{\prime}}(t_{2})_{\ast}(\beta_{E}^{K})}$$\scriptstyle{(t_{1})_{\ast}\Phi_{(t^{\prime})^{\ast}K^{\prime}}(\beta_{E}^{K})}$$\scriptstyle{\Phi_{K^{\prime}}(t_{2})_{\ast}(\tilde{\gamma})}$$\scriptstyle{(t_{1})_{\ast}\Phi_{(t^{\prime})^{\ast}K^{\prime}}(\tilde{\gamma})}$$\scriptstyle{\alpha_{A}^{K^{\prime}}}$
+In this diagram, we note the upper square commutes by naturality of $\zeta^{1}$, the middle square commutes by Lemma˜4.5 (it is Diagram˜4.9 with $K$ replaced by $K^{\prime}$), and the bottom rectangle commutes by naturality of $\alpha^{K^{\prime}}$.
+Set $\delta$ for the left vertical composite in the last diagram, i.e.
+$\delta=(\alpha^{K^{\prime}}_{A})^{-1}\circ\gamma^{\flat_{1}}\colon E\to\Phi_{K^{\prime}}(t_{2})_{\ast}A$.
+Then it follows that
+$\delta^{\sharp}\colon\Phi_{K}E\xrightarrow{\zeta^{2}_{\Phi_{K}E}}(t_{2})_{\ast}t_{2}^{\ast}\Phi_{K}E\xrightarrow{(t_{2})_{\ast}(\beta^{K}_{E})}(t_{2})_{\ast}\Phi_{(t^{\prime})^{\ast}K}t_{1}^{\ast}E\xrightarrow{(t_{2})_{\ast}(\tilde{\gamma})}(t_{2})_{\ast}A,$
+which tells us
+$(\delta^{\sharp})^{\sharp_{2}}\colon t_{2}^{\ast}\Phi_{K}E\xrightarrow{\beta_{E}^{K}}\Phi_{(t^{\prime})^{\ast}K}t_{1}^{\ast}E\xrightarrow{\tilde{\gamma}}A.$
 Consequently, we have
-$h=\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\mathbf{L}t^{\ast}_{1}h^{\prime\prime}=\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\mathbf{L}t_{1}^{\ast}((\alpha^{A})^{-1}\circ\beta^{\prime\prime\prime}),$
-and hence, uniqueness as desired to complete the proof. ∎
-Lemma 4.3.
-Let $F\colon\mathcal{S}_{1}\leftrightarrows\mathcal{S}_{2}\colon G$ be an adjoint pair of exact functors between compactly generated triangulated categories.
-Assume $G$ preserves small coproducts.
-Denote by $\eta\colon 1\to GF$ for the unit of the adjunction.
-Then the collection $\mathcal{T}$ in $\mathcal{S}_{1}$ consisting of objects $E$ for which $\eta_{E}$ is an isomorphism is a triangulated subcategory of $\mathcal{S}_{1}$ which is closed under small coproducts.
-Moreover, $F$ is fully faithful if, and only if, $\eta_{E}$ is an isomorphism for every $E\in\mathcal{S}_{1}^{c}$.
+$\tilde{\gamma}$ $=(\delta^{\sharp})^{\sharp_{2}}\circ(\beta_{E}^{K})^{-1}$
+$=([(\alpha^{K^{\prime}}_{A})^{-1}\circ\gamma^{\flat_{1}}]^{\sharp})^{\sharp_{2}}\circ(\beta_{E}^{K})^{-1},$
+which shows uniqueness of $\tilde{\gamma}$. Conversely, defining $\tilde{\gamma}$ by the last formula and reading the previous derivations in reverse order, it follows that Diagram˜4.12 commutes. ∎
+Remark 4.8.
+Consider the situation as in Proposition 4.7.
+Denote $\varepsilon$ and $\varepsilon^{\prime}$ respectively for the counits of the adjunctions $\Phi_{K}\dashv\Phi_{K^{\prime}}$ and $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\dashv\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$.
+From Lemma˜4.1, it follows that (after possibly replacing $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ by an isomorphic functor) the following diagram commutes:
+(4.13) $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\mathbf{L}t_{1}^{\ast}\Phi_{K^{\prime}}$$\mathbf{L}t_{2}^{\ast}\Phi_{K}\Phi_{K^{\prime}}$$\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}\mathbf{L}t_{2}^{\ast}$$\mathbf{L}t_{2}^{\ast}.$$\scriptstyle{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}(\beta^{K^{\prime}})}$$\scriptstyle{\beta^{K}_{\Phi_{K^{\prime}}}}$$\scriptstyle{\mathbf{L}t_{2}^{\ast}(\varepsilon)}$$\scriptstyle{\varepsilon^{\prime}_{\mathbf{L}t_{2}^{\ast}}}$
+Also, the proof of $\mathcal{P}(E)$ never used that $t$ is affine.
+Specifically, we showed that with ˜3.1 and given $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ relatively perfect over $Y_{2}$, it is the case that $\mathcal{P}(E)$ holds for $E\in D_{\operatorname{qc}}(Y_{1})$.
+[^1] The proof of Proposition 4.7 only needs $t$ affine to guarantee that $\Phi_{(t^{\prime})^{\ast}K}\dashv\Phi_{(t^{\prime})^{\ast}K^{\prime}}$, so that $\mathcal{P}(E)$ plus a replacement of $\Phi_{(t^{\prime})^{\ast}K}$ by a certain isomorphic functor yields a morphism of adjunctions as in the statement of Proposition 4.7.
+Thus, if it also happens that $t_{1}^{\ast}\colon D_{\operatorname{qc}}(Y_{1})\to D_{\operatorname{qc}}(Y_{1}\times_{S}T)$ is essentially surjective, then $\Phi_{(t^{\prime})^{\ast}K^{\prime}}\colon D_{\operatorname{qc}}(Y_{2}\times_{S}T)\to D_{\operatorname{qc}}(Y_{1}\times_{S}T)$ has a left adjoint $L$ (see e.g.
+[BOR94, Theorem 3.1.5]) such that there is an isomorphism
+$L(A)\cong\Phi_{(t^{\prime})^{\ast}K}(A)$
+for each $A\in D_{\operatorname{qc}}(Y_{1}\times_{S}T)$. However, one should be careful, for the isomorphism above might not be natural in $A$.
+Proposition 4.9.
+Consider ˜3.1.
+Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over each $Y_{i}$.
+If $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is fully faithful (resp.
+an equivalence) on $D^{b}_{\operatorname{coh}}$ where $t$ is one of the following:
+  1. (1)
+    $t$ is affine and faithfully flat (e.g. $t=1_{S}$)
+  2. (2)
+    $t\colon\operatorname{Spec}(k)\to S$ is any morphism from a field (so here, we want the condition to all for all such morphisms)
+  3. (3)
+    $t\colon\operatorname{Spec}(\kappa(p))\to S$ is any closed immersion associated to a closed point $p\in S$ (so here, we want the condition to all for all such morphisms),
+then so is $\Phi_{K}$.
 Proof.
-It is straightforward to see that $\mathcal{T}$ is triangulated.
-So, we start by showing $\mathcal{T}$ is closed under small coproducts.
-Let $E_{i}\in\mathcal{T}$ with $i\in I$ for $I$ a set.
-As $F$ and $G$ both commute with small coproducts, the natural morphism $\alpha\colon\oplus_{i\in I}GF(E_{i})\to GF(\oplus_{i\in I}E_{i})$ is an isomorphism.
-Also, we have a commutative diagram for each $i\in I$,
-$E_{i}$$GF(E_{i})$$\bigoplus_{i\in I}E_{i}$$GF(\bigoplus_{i}E_{i})$$\scriptstyle{\eta_{E_{i}}}$$\scriptstyle{\pi_{i}}$$\scriptstyle{GF(\pi_{i})}$$\scriptstyle{\eta_{\oplus_{i\in I}E_{i}}}$
-where $\pi_{i}$ is the natural injection. If $\pi^{\prime}_{i}\colon GF(E_{i})\to\oplus_{i\in I}GF(E_{i})$ is the natural injection, then we know that $GF(\pi_{i})=\alpha\circ\pi_{i}^{\prime}$.
-It suffices to show $\eta_{\oplus_{i\in I}E_{i}}$ is an isomorphism.
-Note that $\eta_{\oplus_{i\in I}E_{i}}$ is the unique morphism arising from the universal property of $\oplus_{i\in I}E_{i}$ being a coproduct of the $E_{i}$.
-Hence, if $GF(\oplus_{i}E_{i})$ with morphisms $GF(\pi_{i})\circ\eta_{E_{i}}$ satisfies the necessary universal property, then $\eta_{\oplus_{i\in I}E_{i}}$ is an isomorphism.
-Consider an $A\in\mathcal{T}$ and morphisms $f_{i}\colon E_{i}\to A$.
-This gives us morphisms $f_{i}\circ\eta^{-1}_{E_{i}}\colon GF(E_{i})\to A$, and hence, a unique morphism $f\colon\oplus_{i\in I}GF(E_{i})\to A$ such that $f_{i}\circ\eta^{-1}_{E_{i}}=f\circ\pi^{\prime}_{i}$ for all $i\in I$.
-It can be checked that $f_{i}\circ\eta^{-1}_{E_{i}}=f\circ\alpha^{-1}\circ GF(\pi_{i})$, and by uniqueness of $f$, we can see that $f\circ\alpha^{-1}$ is the only morphism from $GF(\oplus_{i\in I}E_{i})$ which can satisfy this property.
-Consequently, we have $GF(\oplus_{i\in I}E_{i})$ satisfying the necessary universal property to show that $\eta_{\oplus_{i\in I}E_{i}}$ is an isomorphism.
-So, $\mathcal{T}$ is closed under small coproducts.
-Now, we check the second claim.
-If $F$ is fully faithful, then $\eta_{E}$ is an isomorphism for all $E\in\mathcal{S}_{1}$.
-So, we need to check that $\eta_{E}$ being an isomorphism for every $E\in\mathcal{S}_{1}^{c}$ characterizes fully faithfulness of $F$ on $\mathcal{S}_{1}$.
-In such a case, we know that $\mathcal{S}_{1}^{c}\subseteq\mathcal{T}$.
-However, as $\mathcal{T}$ is a triangulated subcategory of $\mathcal{S}_{1}$ that is closed under small coproducts and contains $\mathcal{S}_{1}^{c}$, it follows that $\mathcal{S}_{1}\subseteq\mathcal{T}$.
-This completes the proof.
+Set $K^{\prime}$ to be the kernel of the integral transform obtained in Proposition˜3.8 which is right adjoint to $\Phi_{K}$ on $D_{\operatorname{qc}}$.
+By Theorem˜3.6, $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{2}\times_{S}T$ for each $t\colon T\to S$ in the statement.
+Hence, Proposition˜3.8 tells us that $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ is right adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ on $D_{\operatorname{qc}}$ for each $t$ in the statement.
+So, from Corollary˜3.9, we know that $\Phi_{K}\dashv\Phi_{K^{\prime}}$ and $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\dashv\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ on $D^{b}_{\operatorname{coh}}$ where $t$ is any morphism in the statement.
+We start with the first case.
+Denote $\eta$ and $\varepsilon$ (resp.
+$\eta^{\prime}$, $\varepsilon^{\prime}$) for the unit and counit of the adjunction $\Phi_{K}\dashv\Phi_{K^{\prime}}$ (resp.
+of $\Phi_{(t^{\prime})^{\ast}K}\dashv\Phi_{(t^{\prime})^{\ast}K^{\prime}}$).
+Since $t$ is faithfully flat, so are each $t_{i}$; hence each $\mathbf{L}(t_{i})^{\ast}$ reflects isomorphisms.
+Therefore, by commutativity of Diagram˜4.10 (resp.
+by Diagram˜4.13), it follows that $\eta^{\prime}$ (resp.
+$\varepsilon^{\prime}$) being an isomorphism implies $\eta$ (resp.
+$\varepsilon$) is an isomorphism.
+The result then follows from [STA25, Tag 07RB].
+Now, we check the second and third case.
+In both cases, $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is fully faithful (resp.
+an equivalence) on $D^{b}_{\operatorname{coh}}$ with $t\colon\operatorname{Spec}(\kappa(p))\to S$ is any closed immersion associated to a closed point $p\in S$.
+In this case, one can argue exactly as in [HLS09, Proposition 2.15].
+Indeed, we only need to replace Proposition 2.9 of loc.
+cit.
+with Proposition˜3.8, i.e.
+‘locally projective’ constraints of loc.
+cit.
+can be replaced with ‘proper’ in light of our work.
 ∎
 Proposition 4.10.
-Consider ˜3.1 where $t=1_{S}$ and $S$ is semiseparated.
+Consider ˜3.1 where $t$ is affine.
 Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{2}$.
-Denote by $K^{\prime}$ for the kernel of the right adjoint for $\Phi_{K}$ obtained in Proposition˜3.8.
-Set $\eta\colon 1\to\Phi_{K}\Phi_{K^{\prime}}$ to be the unit of the adjunction $\Phi_{K}$ and $\Phi_{K^{\prime}}$ on $D_{\operatorname{qc}}$.
-Then the following are equivalent:
+Denote by $K^{\prime}$ for the kernel of the integral transform obtained in Proposition˜3.8 which is right adjoint to $\Phi_{K}$ on $D_{\operatorname{qc}}$.
+Then, after possibly replacing $\Phi_{K^{\prime}}$ by an isomorphic functor, we have a morphism of adjoint pairs as depicted in the diagram
+$D_{\operatorname{qc}}(Y_{1}\times_{S}T)$$D_{\operatorname{qc}}(Y_{1})$$D_{\operatorname{qc}}(Y_{2}\times_{S}T)$$D_{\operatorname{qc}}(Y_{2})$$\scriptstyle{\mathbf{R}(t_{1})_{\ast}}$$\scriptstyle{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}}$$\scriptstyle{\Phi_{K}}$$\scriptstyle{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}}$$\scriptstyle{\mathbf{R}(t_{2})_{\ast}}$$\scriptstyle{\Phi_{K^{\prime}}}\scriptstyle{\dashv}\scriptstyle{\dashv}$
+with left and right comparison transformations given respectively by the isomorphisms
+$\alpha^{K}\colon\Phi_{K}\circ\mathbf{R}(t_{1})_{\ast}\to\mathbf{R}(t_{2})_{\ast}\circ\Phi_{\mathbf{L}(t^{\prime})^{\ast}K},$
+$(\alpha^{K^{\prime}})^{-1}\colon\mathbf{R}(t_{1})_{\ast}\circ\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}\to\Phi_{K^{\prime}}\circ\mathbf{R}(t_{2})^{\ast}.$
+Proof.
+This argument is essentially dual to that of Proposition˜4.7. ∎
+Remark 4.11.
+Consider the situation as in Proposition˜4.10.
+Denote $\varepsilon$ and $\varepsilon^{\prime}$ respectively for the counits of the adjunctions $\Phi_{K}\dashv\Phi_{K^{\prime}}$ and $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\dashv\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$.
+From Lemma˜4.1, it follows that (after possibly replacing $\Phi_{K^{\prime}}$ by an isomorphic functor) the following diagrams commute:
+(4.14) $\Phi_{K}\mathbf{L}(t_{1})_{\ast}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$$\mathbf{R}(t_{2})_{\ast}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$$\Phi_{K}\Phi_{K^{\prime}}\mathbf{R}(t_{2})_{\ast}$$\mathbf{R}(t_{2})_{\ast}$$\scriptstyle{\alpha^{K}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}}}$$\scriptstyle{\mathbf{R}(t_{2})_{\ast}(\varepsilon^{\prime})}$$\scriptstyle{\Phi_{K}(\alpha^{K^{\prime}})}$$\scriptstyle{\varepsilon_{\mathbf{R}(t_{2})_{\ast}}}$
+(4.15) $\mathbf{R}(t_{1})_{\ast}$$\Phi_{K^{\prime}}\Phi_{K}\mathbf{R}(t_{1})_{\ast}$$\mathbf{R}(t_{1})_{\ast}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$$\Phi_{K^{\prime}}\mathbf{R}(t_{2})_{\ast}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}.$$\scriptstyle{\eta_{\mathbf{R}(t_{1})_{\ast}}}$$\scriptstyle{\mathbf{R}(t_{1})_{\ast}(\eta^{\prime})}$$\scriptstyle{\Phi_{K^{\prime}}(\alpha^{K})}$$\scriptstyle{\alpha^{K^{\prime}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}}}$
+Proposition 4.12.
+Consider ˜3.1 where $t$ is affine.
+Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over each $Y_{i}$.
+If $\Phi_{K}$ is fully faithful (resp.
+an equivalence) on $D^{b}_{\operatorname{coh}}$, then so is $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$.
+Proof.
+Denote by $K^{\prime}$ for the kernel of the integral transform obtained in Proposition˜3.8 which is right adjoint to $\Phi_{K}$ on $D_{\operatorname{qc}}$.
+By Theorem˜3.6, $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{2}\times_{S}T$.
+Hence, Proposition˜3.8 tells us that $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ is right adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ on $D_{\operatorname{qc}}$.
+So, from Corollary˜3.9, we know that $\Phi_{K}\dashv\Phi_{K^{\prime}}$ and $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\dashv\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ on $D^{b}_{\operatorname{coh}}$.
+Now, let $\eta$ and $\varepsilon$ (resp.
+$\eta^{\prime}$ and $\varepsilon^{\prime}$) be the unit and counit of the adjunction $\Phi_{K}\dashv\Phi_{K^{\prime}}$ (resp.
+of $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\dashv\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$).
+Since $t$ is affine, so are $t_{1},t_{2}$.
+Therefore, by commutativity of Diagram˜4.15 (resp.
+by Diagram˜4.14), from [STA25, Tag 08I8] we get that $\eta$ (resp.
+$\varepsilon$) being an isomorphism implies $\eta^{\prime}$ (resp.
+$\varepsilon^{\prime}$) is an isomorphism.
+The result then follows from [STA25, Tag 07RB].
+∎
+Theorem 4.13.
+Consider ˜3.1. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over each $Y_{i}$. Then the following are equivalent:
   1. (1)
-    $\Phi_{K}$ is fully faithful on $D_{\operatorname{qc}}$
+    $\Phi_{K}$ is fully faithful (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$
   2. (2)
-    $\eta_{E}$ is an isomorphism for all $E\in D^{-}_{\operatorname{coh}}(Y_{1})$
+    $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is fully faithful (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$ for any affine morphism $t$
   3. (3)
-    $\eta_{E}$ is an isomorphism for all $E\in D^{b}_{\operatorname{coh}}(Y_{1})$
+    $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is fully faithful (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$ for every morphism $t\colon\operatorname{Spec}(k)\to S$ from a field
   4. (4)
-    $\eta_{E}$ is an isomorphism for all $E\in\operatorname{Perf}(Y_{1})$
-Additionally, if $K$ is relatively perfect over $Y_{1}$, then these are equivalent to $\Phi_{K}$ is fully faithful on $D^{b}_{\operatorname{coh}}$.
+    $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is fully faithful (resp.
+an equivalence) on $D^{b}_{\operatorname{coh}}$ for every closed point $p\in S$ with associated closed immersion $t\colon\operatorname{Spec}(\kappa(p))\to S$
+  5. (5)
+    $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is fully faithful (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$ for some affine surjection $t$.
 Proof.
 […]
-Remark 4.5.
-In Proposition˜4.4, a reasonable generalization of the last claim could be to replace $D^{b}_{\operatorname{coh}}$ with $\operatorname{Perf}$ or $D^{-}_{\operatorname{coh}}$.
-Theorem 4.6.
-Consider ˜3.1 where $t$ is an affine morphism locally of finite tor-dimension and $S$ is semiseparated.
-Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{2}$.
-If $\Phi_{K}$ is fully faithful (resp.
-an equivalence) on $D_{\operatorname{qc}}$, then so is $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$.
-Proof.
-First, we check the case of fully faithfulness.
-Denote by $K^{\prime}$ for the kernel of the right adjoint for $\Phi_{K}$ on $D_{\operatorname{qc}}$ that is obtained in Proposition˜3.8.
-Recall that Proposition˜3.8 ensures $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is left adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ on $D_{\operatorname{qc}}$.
-Moreover, $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ preserves small coproducts as all integral transforms do.
-Let $\mathcal{T}$ be the subcategory of $D_{\operatorname{qc}}(Y_{1}\times_{S}T)$ for which the unit of $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ and $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ is an isomorphism.
-By Lemma˜4.3, $\mathcal{T}$ is a triangulated subcategory which is closed under small coproducts.
-Using Proposition˜4.2, we see that $\mathbf{L}t_{1}^{\ast}\operatorname{Perf}(Y_{1})\subseteq\mathcal{T}$.
-However, $t_{1}$ is affine, so $\mathbf{L}t_{1}^{\ast}P$ is a compact generator of $D_{\operatorname{qc}}(Y_{1}\times_{S}T)$ if $P$ is such for $D_{\operatorname{qc}}(Y_{1})$ (see e.g.
-[Sta25, Tag 0BQT]).
-Hence, Proposition˜4.4 ensures that $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is fully faithful on $D_{\operatorname{qc}}$ as desired.
-Indeed, $\mathcal{T}$ containing $\mathbf{L}t_{1}^{\ast}P$ ensures it contains all the compact objects because
-$\operatorname{Perf}(Y_{1}\times_{S}T)=\langle\mathbf{L}t_{1}^{\ast}P\rangle\subseteq\mathcal{T}.$
-To see the the inclusion, use that any triangulated subcategory closed under small coproducts is closed under direct summands (e.g. via the ‘Eilenberg swindle’).
-Lastly, we check the case of $\Phi_{K}$ being an equivalence.
-From the first claim, it suffices to show $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is essentially surjective.
-If $\Phi_{K}$ is an equivalence and is left adjoint to $\Phi_{K^{\prime}}$ on $D_{\operatorname{qc}}$, then one can show that $\Phi_{K^{\prime}}$ is also left adjoint to $\Phi_{K}$ on $D_{\operatorname{qc}}$.
-By Proposition˜3.8, $K^{\prime}\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$.
-Moreover, as $\Phi_{K^{\prime}}$ is an equivalence on $D_{\operatorname{qc}}$, it must also restrict to an equivalence on $\operatorname{Perf}$.
-Hence, $K^{\prime}$ must be relatively perfect over $Y_{1}$.
-Now, by the above $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ is fully faithful and by Proposition 4.4 the unit of the adjoint pair $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ and $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is an isomorphism.
-However, this implies $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is essentially surjective, which completes the proof.
-∎
-Corollary 4.7.
-Consider ˜3.1 where $t$ is an affine morphism locally of finite tor-dimension and $S$ be semiseparated.
-Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over each $Y_{i}$.
-If $\Phi_{K}$ restricts to a fully faithful functor (resp.
-an equivalence) on $D^{b}_{\operatorname{coh}}$, then so does $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$.
-Proof.
-First, we show the case of being fully faithful.
-By Proposition˜4.4, $\Phi_{K}$ restricting to a fully faithful functor on $D^{b}_{\operatorname{coh}}$ is equivalent to $\Phi_{K}$ being fully faithful on $D_{\operatorname{qc}}$ because $K$ is relatively perfect over each $Y_{i}$.
-So, the desired claim follows from Theorem˜4.6 and a further application of Proposition˜4.4.
-Next, we check the case of $\Phi_{K}$ being an equivalence on $D^{b}_{\operatorname{coh}}$.
-By Remark˜2.7, $\Phi_{K}$ restricts to an equivalence on $D^{b}_{\operatorname{coh}}$ implies $\Phi_{K}$ is an equivalence on $D_{\operatorname{qc}}$.
-From Theorem˜4.6, $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is an equivalence on $D_{\operatorname{qc}}$.
-However, once again, Remark˜2.7 tells us $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ restricts to an equivalence on $D^{b}_{\operatorname{coh}}$ as desired.
-∎
 5. Applications
-In this section, we study the consequences of Theorem˜4.6. By ‘genus’ of a curve, we mean ‘arithmetic’(see e.g. [Sta25, Tag 0BY6]).
-Proposition 5.1.
-Let $k$ be a field of characteristic zero.
-Suppose $C$ and $C^{\prime}$ are geometrically integral projective curves over $k$.
-If $C$ and $C^{\prime}$ are Fourier–Mukai partners over $k$, then $\operatorname{Spec}(L)\times_{k}C\cong\operatorname{Spec}(L)\times_{k}C^{\prime}$ for some finite field extension $L/k$ (i.e.
-are $k$-forms).
-Proof.
+In this section, we study the consequences of Theorem˜4.13. By ‘genus’ of a curve, we mean ‘arithmetic’(see e.g. [STA25, Tag 0BY6]).
+Proof of Proposition˜1.4.
 Let $\overline{k}$ be the algebraic closure of $k$.
@@
 Denote by $\pi\colon(C\times_{k}C^{\prime})\times_{k}\operatorname{Spec}(\overline{k})\to C\times_{k}C^{\prime}$ the natural morphism.
-From Theorem˜4.6, we see that $\Phi_{\mathbf{L}\pi^{\ast}K}$ restricts to an equivalence $D^{b}_{\operatorname{coh}}(C\times_{k}\operatorname{Spec}(\overline{k}))\to D^{b}_{\operatorname{coh}}(C^{\prime}\times_{k}\operatorname{Spec}(\overline{k}))$.
+From Theorem˜4.13, we see that $\Phi_{\mathbf{L}\pi^{\ast}K}$ restricts to an equivalence $D^{b}_{\operatorname{coh}}(C\times_{k}\operatorname{Spec}(\overline{k}))\to D^{b}_{\operatorname{coh}}(C^{\prime}\times_{k}\operatorname{Spec}(\overline{k}))$.
 Then $C\times_{k}\operatorname{Spec}(\overline{k})$ and $C^{\prime}\times_{k}\operatorname{Spec}(\overline{k})$ are Fourier–Mukai partners over an algebraically closed field of characteristic zero.
@@
   • $C\times_{k}\operatorname{Spec}(\overline{k})$ is not Gorenstein [SPE23, Theorem A]
-  • $C\times_{k}\operatorname{Spec}(\overline{k})$ Gorenstein of genus one and trivial canonical line bundle [LM14, Theorem 1.1] ($C\times_{k}\operatorname{Spec}(\overline{k})$ is integral with genus one such that $\operatorname{deg}(\omega_{X})=0$ implies $\omega_{X}\cong\mathcal{O}_{X}$ by [Cat82, Proposition 1.11] where $\omega_{X}$ is the canonical sheaf).
+  • $C\times_{k}\operatorname{Spec}(\overline{k})$ Gorenstein of genus one and trivial canonical line bundle [LÓP14, Theorem 1.1] ($C\times_{k}\operatorname{Spec}(\overline{k})$ is integral with genus one such that $\operatorname{deg}(\omega_{X})=0$ implies $\omega_{X}\cong\mathcal{O}_{X}$ by [CAT82, Proposition 1.11] where $\omega_{X}$ is the canonical sheaf).
 Observe these are the only possible cases for projective curves over an algebraically closed field in characteristic zero (see e.g.
@@
 Corollary 5.1.
-Let $S$ be a quasi-compact separated regular $\mathbb{Q}$-scheme.
+Let $S$ be a Noetherian $\mathbb{Q}$-scheme.
 Suppose $f_{i}\colon Y_{i}\to S$ are proper flat morphisms whose special fibers are geometrically integral curves.
-Assume $f_{1}$ is locally projective.
 If $Y_{1}$ and $Y_{2}$ are Fourier–Mukai partners over $S$, then then genii of the special fibers of $Y_{1}$ coincide with that of $Y_{2}$.
@@
 $\kappa(s)$ has characteristic zero).
-Note that, as $S$ is separated, $\operatorname{Spec}(\kappa(s))\to S$ is affine (see e.g.
-[Sta25, Tag 01SG]).
+Note that $\operatorname{Spec}(\kappa(s))\to S$ is affine (see e.g.
+[STA25, Tag 01SI]).
 Denote by $Y_{i}^{s}$ for the base change of $Y_{i}$ along $\operatorname{Spec}(\kappa(s))\to S$.
-Our hypothesis ensures that each $Y_{i}^{s}$ is a geometrically integral proper curve over $\kappa(s)$.
-Then, by Theorem˜4.6, $Y^{s}_{1}$ and $Y_{2}^{s}$ are Fourier–Mukai partners over $\kappa(s)$.
-Indeed, the hypothesis on $S$ ensures it is Noetherian, and so the closed immersion $\operatorname{Spec}(\kappa(s))\to S$ is locally of finite tor-dimension (see e.g.
-[GW23, Remark 25.39(2)]).
-Furthermore, Proposition˜5.1 guarantees that there is a finite field extension $K(s)/\kappa(s)$ such that $Y^{s}_{1}\times_{\kappa(s)}\operatorname{Spec}(K(s))$ is isomorphism to $Y^{s}_{2}\times_{\kappa(s)}\operatorname{Spec}(K(s))$.
+Our hypothesis ensures that each $Y_{i}^{s}$ is a geometrically integral projective curve over $\kappa(s)$ (see e.g.
+[STA25, Tag 0A26]).
+Then, by Theorem˜4.13, $Y^{s}_{1}$ and $Y_{2}^{s}$ are Fourier–Mukai partners over $\kappa(s)$.
+Furthermore, Proposition˜1.4 guarantees that there is a finite field extension $K(s)/\kappa(s)$ such that $Y^{s}_{1}\times_{\kappa(s)}\operatorname{Spec}(K(s))$ is isomorphism to $Y^{s}_{2}\times_{\kappa(s)}\operatorname{Spec}(K(s))$.
 Hence, if coupled with [STA25, Tag 0BY9 & Tag 0FD2], the genii of each $Y^{s}_{i}\times_{\kappa(s)}\operatorname{Spec}(K(s))$ coincides with $Y^{s}_{j}$ for $i\not=j$.
@@
 ∎
-Proposition 5.3.
+Lemma 5.2.
 Let $Y_{1}$ and $Y_{2}$ be proper schemes over a field $k$ (which need not be perfect).
@@
 Proof.
-By Theorem˜4.6, $Y_{1}\times_{k}\operatorname{Spec}(L)$ and $Y_{2}\times_{k}\operatorname{Spec}(L)$ are Fourier–Mukai partners for every finitely generated field extension $L/k$.
+By Theorem˜4.13, $Y_{1}\times_{k}\operatorname{Spec}(L)$ and $Y_{2}\times_{k}\operatorname{Spec}(L)$ are Fourier–Mukai partners for every finitely generated field extension $L/k$.
 So, $Y_{1}\times_{k}\operatorname{Spec}(L)$ is regular.
@@
 ∎
-Corollary 5.4.
-Let $f_{i}\colon Y_{i}\to S$ be locally projective flat morphisms to a quasi-compact separated regular scheme.
-Assume that $Y_{1}$ and $Y_{2}$ are Fourier–Mukai partners over $S$.
-Then $f_{1}$ is smooth if, and only if, $f_{2}$ is smooth.
-Proof.
+Proof of Proposition˜1.5.
 We prove only one direction as the other follows nearly verbatim.
@@
 The natural morphism $t_{s}\colon\operatorname{Spec}(\mathcal{O}_{S,s})\to S$ is affine (see e.g.
-[Sta25, Tag 01SG]) and of finite tor-dimension (due to being flat).
-Hence, Theorem˜4.6 tells us the base changes of $Y_{1}$ and $Y_{2}$ along $t_{s}$ are Fourier–Mukai partners.
-A further application Theorem˜4.6 implies these base changes remain Fourier–Mukai partners if we base change further along the natural closed immersion $\operatorname{Spec}(\kappa(p))\to\operatorname{Spec}(\mathcal{O}_{S,s})$.
-Indeed, $S$ being regular implies $\mathcal{O}_{S,s}$ is as well, and so the closed immersion from $\operatorname{Spec(\kappa(p))}$ is of finite tor-dimension (see e.g.
-[GW23, Remark 25.39(2)]).
+[STA25, Tag 01SI]).
+Hence, Theorem˜4.13 tells us the base changes of $Y_{1}$ and $Y_{2}$ along $t_{s}$ are Fourier–Mukai partners.
+A further application Theorem˜4.13 implies these base changes remain Fourier–Mukai partners if we base change further along the natural closed immersion $\operatorname{Spec}(\kappa(p))\to\operatorname{Spec}(\mathcal{O}_{S,s})$.
 However, smoothness is stable under base change, so each base change of $Y_{1}$ along such morphisms remains smooth.
-Hence, the desired claim follows from Proposition˜5.3.
+Hence, the desired claim follows from Lemma˜5.2.
 ∎
-Proposition 5.5.
-Let $f_{i}\colon Y_{i}\to S$ be projective flat morphisms with geometrically integral fibers.
-Suppose $S$ is quasi-compact, separated, and regular.
-Assume that $Y_{1}$ and $Y_{2}$ are Fourier–Mukai partners over $S$.
-Then $f_{1}$ is Cohen–Macaulay (resp.
-Gorenstein) if, and only if, $f_{2}$ satisfies the same condition.
-Proof.
+Proof of Proposition˜1.6.
 It suffices to show $f_{2}$ is Cohen–Macaulay (resp.
 Gorenstein) if $f_{1}$ satisfies the same condition because the proof is analogous for the converse.
-In fact, we argue in a similar fashion to Corollary˜5.4.
+In fact, we argue in a similar fashion to Proposition˜1.5.
 Moreover, by [STA25, Tag 045O] (resp.
@@
 The natural morphism $t_{s}\colon\operatorname{Spec}(\mathcal{O}_{S,s})\to S$ is affine (see e.g.
-[Sta25, Tag 01SG]) and of finite tor-dimension (due to being flat).
-Hence, Theorem˜4.6 tells us the base changes of $Y_{1}$ and $Y_{2}$ along $t_{s}$ are Fourier–Mukai partners.
-Then Theorem˜4.6 implies these base changes remain Fourier–Mukai partners if we base change further along the natural closed immersion $\operatorname{Spec}(\kappa(p))\to\operatorname{Spec}(\mathcal{O}_{S,s})$.
-Furthermore, if we base change further along $\overline{\kappa(s)}/\kappa(s)$ with $\overline{\kappa(s)}$ an algebraic closure of $\kappa(s)$, these base changes remain Fourier–Mukai partners via Theorem˜4.6.
+[STA25, Tag 01SG]).
+Hence, Theorem˜4.13 tells us the base changes of $Y_{1}$ and $Y_{2}$ along $t_{s}$ are Fourier–Mukai partners.
+Then Theorem˜4.13 implies these base changes remain Fourier–Mukai partners if we base change further along the natural closed immersion $\operatorname{Spec}(\kappa(p))\to\operatorname{Spec}(\mathcal{O}_{S,s})$.
+Furthermore, if we base change further along $\overline{\kappa(s)}/\kappa(s)$ with $\overline{\kappa(s)}$ an algebraic closure of $\kappa(s)$, these base changes remain Fourier–Mukai partners via Theorem˜4.13.
 However, it follows (as fibers are geometrically integral) from [HLS09, Theorem 4.4], that these integral projective schemes over $\overline{\kappa(s)}$ enjoy the property that Cohen–Macaulayness (resp.
@@
 ∎
-Under additional conditions, we can improve Proposition˜5.5 in the Gorenstein case.
 Proposition 5.3.
@@
 [STA25, Tag 0F86]).
-Assume that $Y_{1}$ and $Y_{2}$ are derived equivalences (e.g.
+Assume that $Y_{1}$ and $Y_{2}$ are derived equivalent (e.g.
 the equivalence need not arise from an integral transform).